### v1

# SARATR-X: A Foundation Model for Synthetic Aperture Radar Images Target Recognition

Abstract
Synthetic aperture radar (SAR) is essential in actively acquiring information for Earth observation. SAR Automatic Target Recognition (ATR) focuses on detecting and classifying various target categories under different image conditions. The current deep learning-based SAR ATR methods are typically designed for specific datasets and applications. Various target characteristics, scene background information, and sensor parameters across ATR datasets challenge the generalization of those methods. This paper aims to achieve general SAR ATR based on a foundation model with Self-Supervised Learning (SSL). Our motivation is to break through the specific dataset and condition limitations and obtain universal perceptual capabilities across the target, scene, and sensor. A foundation model named SARATR-X is proposed with the following four aspects: pre-training dataset, model backbone, SSL, and evaluation task. First, we integrated 14 datasets with various target categories and imaging conditions as a pre-training dataset. Second, different model backbones were discussed to find the most suitable approaches for remote-sensing images. Third, we applied two-stage training and SAR gradient features to ensure the diversity and scalability of SARATR-X. Finally, SARATR-X has achieved competitive and superior performance on 5 datasets with 8 task settings, which shows that the foundation model can achieve universal SAR ATR. We believe it’s time to embrace fundamental models for SAR image interpretation in the era of increasing big data.
Index Terms:
Synthetic Aperture Radar (SAR), Target Recognition, Foundation Model, Self-Supervised Learning (SSL), Deep Learning, Masked Image Modeling (MIM)
1 Introduction
Based on electromagnetic scattering in microwave frequency bands, Synthetic Aperture Radar (SAR) [62, 22, 53] is essential in actively acquiring information for Earth observation with stable imaging capacity under various weather and lighting conditions. SAR Automatic Target Recognition (ATR) aims to automatically localize and classify interested objects in SAR images, e.g., detection and classification. It has been investigated extensively with various civilian and military applications [69, 36, 73, 57]. In the past decade, deep learning has reinvigorated SAR ATR with impressive breakthroughs [31, 37, 15]. However, as shown in Fig 1 and 2, due to the extensive acquisition conditions as well as the high costs of collection and annotation, there are a plethora of special datasets and algorithms designed for different applications of SAR ATR. For example, nearly a dozen new target datasets and dozens of attention modules have been developed for SAR target recognition in the last few years. However, the target characteristics, scene background information and sensor parameters across datasets challenge their generalization. In contrast to specialized algorithms for particular applications, we aim to investigate a general SAR target recognition method.
A foundation model [5] pre-trained in a task-agnostic manner (generally via self-supervision learning) on extensive data can be flexibly adapted to a wide range of downstream tasks under various conditions. Self-supervised learning (SSL) [45, 4] can mitigate label inefficiency by exploring supervision directly from the data itself, thereby reducing reliance on expensive expert labeling and efficiently scaling data and models. Hence, SSL has the capability to leverage extensive SAR unlabeled data for constructing foundational models, which can achieve universal recognition with a few labeled target samples in different categories and conditions. As shown in Table I, foundation models are thriving in remote sensing with superior performance in a diverse range of modalities and tasks. In particular, multi-modality foundation models have recently appeared, indicating this research topic’s great potential. But, there lacks a foundation model that can effectively accomplish various tasks for SAR ATR, such as FG-MAE [79] and SkySense [26] mainly focus on SAR scene classification and segmentation, and our previous SAR-JEPA [38] and MSFA [43] explored the object classification and detection tasks separately. It is still necessary to explore the development of an analogous ATR foundation model for SAR image interpretation, which lies at the intersection of SAR technologies and frontier AI technologies (in particular, the big models).
Figure 1: Motivation and performance of SARATR-X. (a) SAR ATR has various imaging conditions, including targets, scenes, and sensors. However, due to high cost, its datasets are collected under specific settings and tasks. For example, the MSTAR dataset [1] is a ten-type vehicle target classification dataset in X-band and grass scenarios, while SAR-Aircraft is a seven-type aircraft target detection dataset in three airports and a C-band satellite. Many specialized algorithms have been proposed for these datasets. However, the different target characteristics, scene background information and sensor parameters between these datasets challenge the generalization of existing algorithms. This paper aims to explore a SAR ATR foundation model, a general method to achieve various tasks. (b) Our SARATR-X performs well on 5 datasets with 8 settings compared with previous methods. For target classification, SARATR-X is better than the previous SSL method (BIDFC [92]) on fine-grained vehicle MSTAR dataset [1] with the few-shot setting. Besides, it performs well under extended operating conditions (EOCs) [94], i.e., the imaging condition variation in depression angle (EOCs-Depression), target configuration (EOCs-Config), and version (EOCs-Version). For object detection, SARATR-X demonstrates competitive performance with previous carefully designed supervised methods on various categories (SARDet-100K [43] and OGSOD [71]) and specific categories on ship (SSDD [97]), and aircraft (SAR-AIRcraft [80]). Our study shows the potential of the foundation model for universal SAR target recognition.
Thanks to the development of SAR sensors and the enormous efforts of researchers, numerous SAR target datasets have emerged over the past five years in Fig. 2. Despite image annotation remaining a challenging and expensive endeavor, SAR sensors have produced vast amounts of data at a significantly accelerated pace compared to the previous ten years. Meanwhile, the investigation of SSL and foundation models for SAR ATR is advancing with increasing datasets. As shown in Table I, existing research [92, 79, 56, 38, 43] have discussed different factors of a foundation model, and these studies are progressively advancing towards a goal of achieving general SAR ATR. For example, BIDFC [92] performs much better than supervised methods in the few-shot classification of the MSTAR vehicle dataset, SAR-JEPA [38] demonstrates that SSL can effectively improve the classification accuracy for different targets across datasets, and MSFA [43] shows improved results of ship detection. Inspired by their insightful work, we formally propose a SAR ATR foundation model, SARATR-X, to achieve a breakthrough in universal SAR target recognition.
In this work, we lay the groundwork for developing SARATR-X, a foundation model specifically tailored for SAR ATR tasks. It aims to learn generalizable representations from unlabeled SAR images, offering a basis for efficient adaptation across various downstream ATR tasks. The four parts for creating SARATR-X include pre-training datasets, model backbones, self-supervised learning, and evaluation tasks.
Pre-training datasets need to contain diverse target categories and imaging conditions to accommodate various downstream tasks. However, SAR ATR lacks a large-scale dataset like ImageNet [19], and the most used MSTAR dataset in Table II only includes fine-grained vehicle categories, which is not suitable for larger-scale pre-training. Therefore, many SSL methods combine different datasets; in particular, SARDet-100K incorporates [43] 9 SAR target detection datasets. With the increasing number of open-source datasets for SAR ATR, we integrate the vast majority of them for pre-training. As shown in Fig 2 and Table III, we apply 14 classification and detection datasets with different target categories and imaging conditions to explore the foundation model’s potential.
Model backbones aim to achieve better spatial representation in remote sensing images, especially small target signatures in large imagery. As shown in Table I, researchers used two architectures (Transformer or convolutional neural network). Transformer has better spatial resolution without downsampling, and Convolutional Neural Network (CNN) has higher efficiency with its convolution kernel. MSFA discussed the effectiveness of different architectures, and the results showed that the Swin Transformer backbone performs better. Therefore, we discuss different model backbones to find the most suitable approaches for the properties of remote sensing images. HiViT [99] is our choice, which combines Swin Transformer advantages and can drop patches in MIM.
Self-supervised learning faces the challenge of SAR image quality. SAR images contain speckle noise due to coherent imaging, and their visual features are not as distinct and rich as RGB images. As shown in Table II, contrastive learning [92, 56] uses data augmentation and pre-processing to reduce noise, while MIM [79, 38, 43] has a pixel-to-feature improvement for guide signals. Therefore, we consider that the problem to be solved by SAR SSL is how to construct high-quality guide signals. For example, PGIL [29] leveraged a sub-frequency feature of SAR complex images to learn physics information, and our SAR-JEPA [38] applied multi-scale gradient ratio to solve multiplicative speckle noise and capture target shape. Furthermore, multi-stage training [43] from ImageNet to SAR diminished noise interference on model diversity in Fig. 5. Therefore, we apply two-stage training from ImageNet to SAR and use multi-scale gradient features as high-quality guide signals for SAR MIM.
Evaluation tasks need to comprehensively evaluate the performance of the foundation model under different tasks and settings. Benefiting from the open-source target datasets, we first construct a fine-grained classification dataset with various categories to evaluate the effectiveness of the proposed improvements. In the end, we perform a comprehensive comparison between the proposed SARATR-X and existing SOTA methods in public classification and detection tasks.
Our SARATR-X has achieved a competitive and superior performance on 5 datasets with 8 task settings in Fig 1. And its codes and weights can be found at https://github.com/waterdisappear/SARATR-X. We hope this work will advance the development of foundation models and general SAR target recognition. The main contributions are summarized as follows:
TABLE I: List of visual foundation models in remote sensing. There are many foundation models for various modalities and tasks, but research (FG-MAE, SkySense, and OFA-Net) focuses on scene-level tasks with the SAR modality. Based on our previous study SAR-JEPA and MSFA [38, 43], SARATR-X fills this gap in the foundation model for SAR ATR. These foundation models’ SSL methods are divided into two approaches: one is contrastive, such as contrastive learning, to obtain invariant features; another is generative, such as masked image modeling, to generate and predict features. SD: Stable Diffusion. Swin: Swin Transformer.
Foundation model Year Modality Dataset Backbone SSL Tasks
SatMAE [13] 2022 Multi-spectral, RGB fMoW RGB/Sentinel ViT Generative Scene classification and semantic segmentation
RVSA [77] 2022 RGB MillionAID ViT, ViTAE Generative Scene classification, object detection, and semantic segmentation
RingMo [65] 2022 RGB Self-built ViT, Swin Generative Scene classification, object detection, semantic segmentation, and change detectionn
RingMo-Sense [90] 2023 RGB Self-built Video Swin Generative object detection and tracking, images segmentation, and 3 prediction tasks
CMID [55] 2023 RGB MillionAID ResNet, Swin Generative & Contrastive Scene classification, object detection, and semantic segmentation
GFM [52] 2023 RGB GeoPile Swin Generative Scene classification, semantic segmentation, change detection, and super-resolution
DiffusionSat [32] 2023 RGB fMoW, Satlas, SpaceNet SD Generative Image generation, super-resolution, temporal generation, and in-painting
Scale-MAE [58] 2023 RGB fMoW RBG ViT Generative Scene classification, object detection, and semantic segmentation
FG-MAE [79] 2023 Multi-spectral, SAR SSL4EO-S12 ViT Generative Scene classification and semantic segmentation
SMLFR [17] 2024 RGB GeoSense ConvNeXt Generative Object detection, semantic segmentation, and change detection
SkySense [26] 2024 3 modalities HSROIs, TMsI, TSARI Swin, Vit Contrastive Scene classification, object detection, semantic segmentation, and change detection
OFA-Net [87] 2024 4 modalities 5 multi-modalities datasets ViT Generative Scene Classification and semantic segmentation
SAR-JEPA [38] 2024 SAR MSAR, SAR-Ship, SARSim, SAMPLE ViT Generative Target classification
MSFA [43] 2024 SAR ImageNet, DOTA, SARDet-100k CNN, ViT, Swin Generative Object detection
SARATR-X 2024 SAR ImageNet & 14 SAR datasets HiViT Generative Target classification and object detection
  1. $\bullet$
    We propose a vision of general SAR target recognition, aiming to find universal solutions for various SAR ATR tasks and application in an era of SAR images with big unlabeled data and small labeled samples.
  2. $\bullet$
    We have systematically investigated a foundation model solution for general SAR ATR and present the first SAR ATR foundation model named SARATR-X, which achieves new performance breakthroughs on a wide range of SAR target recognition datasets and settings.
  3. $\bullet$
    We hope this promising work will stimulate research interest in SAR foundation models, thereby advancing SAR image interpretation with frontier AI technologies.
The remainder of this paper is organized as follows. Sec. 2 introduces related work in remote sensing and SAR ATR. Sec. 3 introduces the proposed foundation model SARATR-X. Sec. 4 and 5 conducts extensive experiments to demonstrate the superiority of the proposed method. Sec. 6 concludes the paper and discusses future work.
2 Related Work
TABLE II: Related work about SSL and foundation models for SAR ATR. It can be found that the exploration of the foundation model in SAR ATR. Researchers have explored different dataset settings, CNN and Transfomer architectures, and various SSL algorithms. Inspired by their successful work, we formally present the first SAR ATR foundation model, SARATR-X, to achieve various recognition tasks.
Method Year Dataset Backbone SSL Tasks Description
RotANet [82] 2021 MSTAR CNN Generative Target classsification (MSTAR) Predicting rotational patterns of targets as a loss regularization term for classification task.
UACL [88] 2021 MSTAR, FUSAR-Ship CNN Contrastive Target classsification (MSTAR) Contrastive defense method to enhance model robustness to different adversarial attack.
PGIL [29] 2022 sea-ice, urban ResNet-18 Contrastive Scene classsification (sea-ice, urban) Contrastive learning between sub-frequency features of complex images and deep features of amplitude images.
BIDFC [92] 2022 MSTAR, OpenSARShip ResNet-18 Contrastive Target classsification (MSTAR, OpenSARShip) Weakly contrastive learning for pre-training in fine-grained datasets with similar sample.
TSCL [56] 2023 MSTAR ResNet-50 Contrastive Target classsification (MSTAR, OpenSARShip) Pre-processing of SAR images for noise removal to improve image quality for Contrastive learning.
FG-MAE [79] 2023 SSL4EO-S12 ViT-S Generative Scene classification (EuroSAT, BigEarthNet-MM), semantic segmentation (DFC2020) Hand-craft feature HOG as the target signal for MIM in SAR images.
SAR-JEPA [38] 2024 MSAR, SAR-Ship, SARSim, SAMPLE ViT-B Generative Target classification (MSTAR, FUSARShip, SAR-ACD) Local reconstruction and multi-scale gradient feature for MIM with small object in noise SAR images.
MSFA [43] 2024 ImageNet, DOTA, SARDet-100k Swin-B Generative Object detection (SARDet-100K, SSDD, HRSID) Multi-stage pre-training strategy for MIM from RGB images to SAR images.
SARATR-X 2024 ImageNet, 14 SAR target datasets HiViT-B Generative Target classification and Object detection (5 datasets and 8 settings) A first foundation model for target recognition in SAR Images.
As shown in Table I and II, visual foundation models are booming in remote sensing. Many models are proposed for various modalities and tasks. Our studies focus on the foundation model for SAR ATR, i.e., SAR images-based target classification and object detection. In the following, we introduce the recent development of remote sensing foundation models.
2.1 Foundation models in remote sensing
Remote sensing foundation models [30] have received widespread attention in the last three years. Researchers have made many breakthroughs, achieving effective learning on various modalities and tasks. In terms of pre-training datasets, researchers use existing large-scale datasets or collect a large number of samples from different sources. As for the model backbone, researchers have improved attention mechanisms, positional encoding, and other aspects to enhance the perception of complex spatial information. MIM has been used to learn spatial-temporal contextual information, while contrast learning is applied to multi-modal learning.
SatMAE [13] proposed a novel masking strategy and temporal and spectral positional encoding for multi-spectral and temporal images. With a new dataset fMoW Sentinel with 13 frequency bands, SatMAE achieved new performance on scene classification and semantic segmentation tasks. RVSA [77] improved the pretraining ViT backbone with a rotated varied-size window attention method for the arbitrary-oriented objects. This work shows the importance of learning targets’ complex spatial contextual relationships in remote sensing images. RingMo [65] used a patch incomplete mask strategy for dense and small objects. Based on their self-built 2 million images, RingMo has proven effective in many tasks. RingMo-Sense [90] offered a three-branch network and masking strategy to model the spatio-temporal interaction for temporal images. CMID [55] combined contrastive learning and masked image modeling to learn global semantic information and local spatial information. GFM [52] focused on the differences between natural and remote sensing images and employed a multi-objective continual pretraining approach to leverage both knowledge. DiffusionSat [32] was the first remote sensing generative that employs geographic information embedding in stable diffusion. Scale-MAE [58] reconstructed images at different frequencies with improved positional encodings of ViT. FG-MAE [79] employed different hand-designed features to replace the original pixels in the MIM and improve the feature quality. And SMLFR [17] used a low pass filter to eliminate high-frequency information from the image pixel.
Multi-modal remote sensing foundation models have been developed, such as SkySense [26] and OFA-Net [87]. SkySense [26] proposed a multi-granularity contrastive learning method to learn representations of different modalities. Besides, a GEO-context prototype was applied to embed geographical contextual information. OFA-Net [87] applied a shared Transformer backbone for multiple modalities. Besides this, there are vision-language models [41], such as EarthGPT [98], SkyEyeGPT [93] and LHRS-Bot [54], which incorporate large language models to different remote sensing images modalities. However, due to the difficulty of annotating SAR images, the collection of public datasets used by EarthGPT only has 10,554 SAR ship images, much less than 84,838 infrared images and 907,945 optical images.
Therefore, there is relatively less research on SAR foundation models due to the scarcity and fragmentation of the high-resolution SAR target datasets in Fig 2. We want to explore the visual foundation model based on SAR images with target recognition to stimulate research enthusiasm in this direction.
Figure 2: Timeline of measured and simulated datasets for SAR target recognition (SAR magnitude images were processed in pseudo-color for better visualization). SAR target datasets have increased rapidly since the 2020s, greatly enriching the diversity of targets, scenes, and sensors. Target categories are mainly fine-grained vehicles, ships, and aircraft. Mixed means that the target category includes more than one category, such as a mixture containing various vehicles, ships, and aircraft in IRIS-SAR. Scenes contain cities, harbors, airports, etc. Sensors are devices with different resolutions and bands under satellite or airborne platforms. However, the number of samples (No. samples) in most datasets is only a few thousand with several categories due to the high costs and difficult annotation. It inspired us to propose SARATR-X with an integrated dataset in Table III.
TABLE III: Description of our pre-training dataset for SARATR-X, which contains 14 open-source SAR datasets for pre-training. # Img.: Number of images. # Target: Number of target categories. # Scene: Number of scenes. Res.: Resolution. Pol.: Polarization. Large SAR imagery in detection datasets contains more target and scene types than the annotation. We cropped some large images to increase the training samples.
Dataset Year # Img. Img. size # Target # Scene Res. (m) Band Pol. Description
AIR-SARShip [66] 2019 801 $512\sim 1000$ $\geq 1$ $\geq 3$ $1\sim 3$ C Single Ship detection dataset in complex scenes
HRSID [81] 2020 5,604 800 $\geq 1$ $\geq 2$ $0.5\sim 3$ C/X Quad Ship detection and instance segmentation dataset
Sandia MiniSAR [60] 2006 3,927 224 $\geq 1$ $\geq 7$ 0.1 Ku Single Terrestrial targets in urban areas, airports, deserts, and others
MSAR [84, 10] 2022 28,499 $256\sim 2048$ $\geq 4$ $\geq 6$ 1 C Quad Terrestrial and maritime targets detection dataset
MSTAR [1] 1995 14,577 $128\sim 193$ 10 1 0.3 X Single Fine-grained vehicle classification dataset
OGSOD [71] 2023 18,331 256 $\geq 3$ $\geq 2$ 3 C Double Detection dataset for bridges, oil tanks, and harbours
OpenSARShip [35] 2017 26,679 $9\sim 445$ 14 10 $2.3\sim 17.4$ C Double Fine-grained maritime target slices
SADD [96] 2022 883 224 $\geq 1$ $\geq 2$ $0.5\sim 3$ X Single Aircraft detection dataset
SAMPLE [34] 2019 5,380 128 10 2 0.3 X Single Simulation and measured vehicle dataset
SAR-AIRcraft [80] 2023 18,818 512 $\geq 7$ $\geq 3$ 1 C Single Aircraft detection dataset
SARSim [51, 33] 2017 21,168 139 14 3 0.3 X Single Simulation vehicle dataset
SAR-Ship [78] 2019 39,729 256 $\geq 1$ $\geq 4$ $3\sim 25$ C Quad Ship detection dataset in complex scenes
SIVED [44] 2023 1,044 512 $\geq 1$ $\geq 4$ $0.1\sim 0.3$ X/Ku/Ka Single Synthetic vehicle detection dataset
SSDD [97] 2021 1,160 $214\sim 668$ $\geq 1$ $\geq 2$ $1\sim 15$ C/X Quad Ship detection dataset
2.2 Foundation models in SAR
Researchers have explored SAR foundation models in different aspects, as shown in Table II. Inspired by previous studies, we systematically investigated how to construct a SAR ATR foundation model.
Early SSL was often used as a regularization loss for classification tasks. RotANet [82] predicted the rotational pattern of MSTAR vehicle targets to capture azimuthal features for the classification task. UACL [88] combined data augmentation and adversarial samples in contrastive learning to improve the model’s robustness to various adversarial attacks. PGIL [29] used contrastive learning between the SAR complex images’ sub-frequency feature and the amplitude images’ deep feature to inject physical knowledge into the classification task. Recently, SSL has been used in model pre-training and fine-tuning frameworks. BIDFC [92] proposed weakly contrastive learning for pre-training in fine-grained vehicle datasets MSTAR and used Gaussian noise data augmentation to simulate SAR image noise. TSCL [56] applied the pre-processing of SAR images before data augmentation in contrastive learning. FG-MAE [79] discussed different hand-craft features for multi-spectral and SAR images and used the HOG feature for SAR. Our previous study, SAR-JEPA [38] and MSFA [43], focused on target classification and object detection. SAR-JEPA [38] applied Local reconstruction and multi-scale gradient feature to capture target spital signatures better. MSFA [43] proposed a multi-stage with filter augmentation pre-training framework to use large-scale RBG and SAR data for detection.
Our insight. These studies have demonstrated that SSL can achieve performance improvements on various categories [38] and tasks [38, 43], and can even be comparable to the performance of specially designed supervised methods [92, 43]. These inspired us to conduct systematic research for foundation models to achieve general SAR target recognition, especially in this era of big data. Firstly, we must extend the pre-dataset with different tasks and scenarios based on existing research. Secondly, a suitable model backbone needs to be discussed for the small target characteristics of remote-sensing images. Thirdly, SSL needs high-quality guide signals from SAR images under noise interference. Finally, we need to evaluate the performance of foundation models comprehensively.
3 Approach
We aim to construct a foundation model for general ATR from large-scale SAR images via the SSL method. As described above, the increasing SAR datasets and SSL studies inspired us to build a foundation model for SAR ATR. Our approach revolves around pre-training datasets, model backbones, SSL methods, and evaluation tasks to provide a systematic benchmark for foundation models in SAR ATR.
Figure 3: Two steps of our pre-training. The first is to perform MIM on ImageNet to obtain better initialization weights, as shown in Fig 5 (c). Besides, we can easily obtain the open-source pre-training weights in step 1, thus reducing the pre-training time. The second is to perform MIM on SAR images with high-quality guide signals. We use multi-scale gradient features to suppress speckle noise and extract target edges.
3.1 Creating a Diverse Pre-training Dataset
Existing work mainly used MSTAR [1] as a pre-training dataset. While MSTAR is a high-quality vehicle target slice dataset, its commonly used samples are only a few thousand. Besides, this dataset suffers from background bias due to a single imaging scene [39]. By comparison, the pre-training set ImageNet-1K in computer vision contains 1.4 million images with different categories and scenes. Using an insufficient SAR target pre-training dataset would underestimate the potential of this research area. Since diverse targets, scenes, and sensor conditions constitute a huge data sampling space in real-world situations, constructing a large pre-training dataset for the foundation model is central to unifying various SAR ATR tasks.
The increasing SAR target datasets in Fig. 2 is a primary motivation for achieving this goal. Although SAR images are expensive and no single dataset contains all popular target categories and various imaging conditions, collecting target samples from various open-source datasets can still construct a pre-training dataset with different categories, scenes, and sensors. Therefore, we build a new pre-training dataset containing 186,600 SAR target samples from 14 open-source SAR target datasets in Table III. We construct the dataset to include, as much as possible, target categories (terrestrial and maritime targets such as vehicles, ships, aircraft, oil tanks, bridges, etc.), scenes (typical scenes such as cities, harbors, airports, oceans, etc.), and sensors (satellite, airborne, and simulation platforms of different resolutions and bands).
3.2 Choosing a Scalable Model
We consider two model backbones for SAR target recognition. For the first, we apply Vision Transformer (ViT) [18], commonly used in SSL and has good scalability in model parameters. For the second architecture, we experiment with ConvNeXt-V2 [83], which has the same scalability as ViT and maintains the efficiency of convolutional neural networks. In addition to the scalability of the model parameters, the image properties need to be considered for remote sensing. The SAR target usually has a small foreground and a dynamic context range. Our previous study, MSFA [43], finds that the Swin Transformer is better than ViT with a hierarchical structure, but it is unsuitable for drop patches in MIM to save computing resources. Therefore, we also consider a variant of ViT, Hierarchical Vision Transformer (HiVit) [99], which improves the input spatial resolution and retains the property of ViT with MIM.
3.3 Pre-Training with two steps
We use MIM as the pretext task and masked autoencoders (MAE) [27] to save computational resources with dropping patches. MIM can help the foundation model achieve SAR image interpretation by recognizing the contextual relationship around objects. Here is a key point when applying MIM. SAR images belong to coherent imaging, and values have speckle noise that can interfere with the pretext task. Therefore, SARATR-X uses two pre-training steps to build a foundation model in Fig 3.
The first step is to perform MIM on ImageNet to obtain better initialization weights. We simplify the multi-stage pre-training of MSFA, which performs SSL on ImageNet with backbone, detection task pre-training on DOTA with the whole framework, and detection task fine-tuning on SAR images. SARATR-X uses the pre-training weight of ImageNet as initialization weights for the SAR pre-training step. This way enhances the diversity of attention during the SAR pre-training step in Fig. 5. In contrast, random initialization leads to a convergence of attention towards the same pattern in SAR pre-training with MAE. Besides, ImageNet pre-training backbone weights can be obtained from open source, thus greatly reducing pre-training time. We refer to using ImageNet pre-trained weights as SSL-ImageNet & SAR.
The second step is to perform MIM with SAR images. As mentioned before, SAR image noise is a very tricky problem, and FG-MAR, SAR-JEPA, and MSFA have discussed many features, such as CannyEdge [8], HOG [14], CannyEdge [8], Haar-like [70], SAR-HOG [61] and SAR-SIFT [16]. We can use different feature combinations to get the best results, but here is the simplest gradient feature that follows our previous SAR-JEPA to avoid getting bogged down on complex feature selection. We use the Multi-scale Gradient Feature (MGF) [38] to suppress the speckle noise and extract the target shape.
TABLE IV: The results on the classification dataset with 25 categories. We perform the linear probing setting on the few-shot SAR target classification. For the pre-training datasets, we recommend using ImageNet’s pre-trained weights as initialization before pre-training on SAR datasets. For the model backbones, we prefer the HiVit to recognize the small target in SAR images. Model backbones are their base version. The classification metric is average accuracy. Bold indicates the best result, underline is the next best result. SL. Supervised Learning. SSL. Self-Supervised Learning.
Backbones Params Pre-training Classification (N-shot)
setting dataset method 5 10 20
ConvNeXt-V2 89M SL-ImageNet ImageNet-1K Supervised 52.5 61.7 70.5
ConvNeXt-V2 89M SSL-ImageNet ImageNet-1K FCMAE 47.2 54.5 64.0
ConvNeXt-V2 89M SSL-SAR SAR images FCMAE 52.7 60.9 67.7
ConvNeXt-V2 89M SSL-ImageNet &\&& SAR ImageNet &\&& SAR FCMAE 54.7 61.5 69.5
ViT 86M SL-ImageNet ImageNet-1K Supervised 58.6 65.7 74.2
ViT 86M SSL-ImageNet ImageNet-1K MAE 50.7 58.0 65.5
ViT 86M SSL-SAR SAR images MAE 54.1 61.5 68.2
ViT 86M SSL-ImageNet &\&& SAR ImageNet &\&& SAR MAE 65.8 76.4 83.6
HiViT 66M SL-ImageNet ImageNet-1K Supervised 49.0 55.8 63.3
HiViT 66M SSL-ImageNet ImageNet-1K MAE 53.0 60.3 69.3
HiViT 66M SSL-SAR SAR images MAE 64.9 72.7 79.9
HiViT 66M SSL-ImageNet &\&& SAR ImageNet &\&& SAR MAE 71.5 78.5 84.0
HiViT 66M SSL-ImageNet &\&& SAR ImageNet &\&& SAR Ours 76.5 80.8 85.1
Multi-scale gradient feature. The classical differential gradient is not a constant false alarm rate operator due to multiplicative speckle noise in the SAR image. It means that speckle noise would cause the gradient computation to have false points in strong target regions. Previous studies [68, 6] have shown that the computing ratio is suitable for multiplicative noise. Here, MGF uses the gradient by ratio [61, 16] to obtain gradient features $G_{\text{m}}$.
$R_{i}$ $=\frac{M_{1}(i)}{M_{2}(i)},$ (1)
$G_{\text{H}}$ $=log(R_{1}),$ (2)
$G_{\text{V}}$ $=log(R_{3}),$ (3)
$G_{\text{m}}$ $=\sqrt{G_{\text{H}}^{2}+G_{\text{V}}^{2}},$ (4)
where $R_{i}$ denotes the average ratio at different directions. $M_{1}(i)$ and $M_{2}(i)$ are the area averages on opposite sides of the current pixel along direction $i$. $i=1$ is the horizontal direction and $i=3$ means the vertical direction. The area averages can be computed from the input image and four fixed convolution kernels. Then, Eq. 2 and 3 use logarithms to solve the vertical gradient calculation [16]. $G_{\text{H}}$ is the horizontal gradient and $G_{\text{V}}$ is the vertical gradient.
$\text{MGF}=\text{concat}(G_{\text{m1}},G_{\text{m2}},G_{\text{m3}})$ (5)
Due to the dynamic range required for various targets in remote sensing [42], MGF is constructed with convolutional kernels of different sizes. We set the kernel scale $r$ equal to 9, 13, and 17 to obtain $G_{\text{m1}}$, $G_{\text{m2}}$, and $G_{\text{m3}}$, and the whole convolutional kernel size is odd square $12r+$ 1.
3.4 Evaluating with Recognition Tasks
We merged fine-grained classification datasets of Vehicles, Ships, and Aircraft as a new SAR classification dataset named (SAR-VSA). This dataset is used for comparing the SSL model’s performance with a few-shot setting in Sec. 4. We then report SARATR-X results on existing classification and detection dataset settings with other methods in Sec. 5.
4 Experiments of SARATR-X
First, we perform the SSL on the pre-training dataset without label information. Then, we fine-tune the pre-trained model on SAR-VSA with few-shot classification task and linear probing setting to analyze the improvement of SARATR-X. Finally, we discuss the scalability of the proposed approach.
We perform pre-training on 8 NVIDIA RTX3090 GPUs. The SAR pre-training dataset consists of fourteen SAR datasets. The few-shot SAR classification dataset, SAR-VSA, contains 25 fine-grained targets from the three SAR datasets. It is difficult to ensure training convergence by fine-tuning the model’s whole parameters under small sample cases. Therefore, we use linear probing [27], which includes a batch normalization layer, to adjust the differences in the data statistical properties and reduce the number of fine-tuning parameters. Detailed settings are in the Appendix  A.
4.1 Comparison of Model Backbones
Table IV compares different model backbones for SAR ATR: ConvNeXt-V2 [83], ViT [18], and HiViT [99]. ConvNeXt-V2 and ViT represent the two main architectures: CNN and Transformer. HiViT combines Vit and Swin Transformer [47].
From the result, we can see that ViT outperforms ConvNeXt-V2. On the one hand, the ViT is more flexible than ConvNeXt-V2 in learning the contextual information in SAR images. On the other hand, multiple downsamplings in ConvNeXt-V2 result in the loss of small targets, while ViT maintains the same spatial resolution in different layers. HiVit, a visual Transformer with ViT’s flexibility and hierarchical representation, performs better than ViT and ConvNeXt-V2 in our self-supervised experiments. In particular, HiViT uses small (4 × 4) input patches, capturing small target features well. Fig. 5 demonstrates that HiViT has a better various attention distance than ViT due to the small target information in remote sensing.
4.2 Strategy of two-step pre-training
Here, we discuss the two-step pre-training strategy to make full use of the available model weights and SAR datasets. Table IV has four pre-training settings: SL-ImageNet, SSL-ImageNet, SSL-SAR, and SSL-ImageNet & SAR. SL-ImageNet is pre-training on ImageNet with supervised learning [^1] We used open-source weights from GitHub to conduct the experiments. The supervised weights of ConvNeXt-V2 and HiViT were obtained by supervised fine-tuning after SSL.; SSL-ImageNet is pre-training on ImageNet with SSL from scratch; SSL-SAR is pre-training on our SAR pre-training dataset from scratch; SSL-ImageNet & SAR pre-training the model on SAR dataset based on the initialized weight from SSL-ImageNet.
It is noticed that the additional supervised information introduced by SL-ImageNet does not necessarily improve the SAR ATR performance, e.g., the linear probing performance of SL-ImageNet for HiVit is lower than that of SSL-ImageNet. SSL-SAR can achieve better results than SSL-ImageNet using less data (12%), reflecting the huge differences in the target features between these two images. However, ImageNet pre-training weights can provide a good initialization for low features such as shape and texture in SSL with visible spectral remote sensing [67] and medical image [103]. Our experiments also confirmed this conclusion: using SSL-ImageNet as initialization weights improves the pre-training performance of SAR images in Table IV and attention diversity in Fig. 5. Therefore, our SARATR-X uses the SSL-ImageNet & SAR setting to complement the richness of the pre-training.
4.3 Design of target signal for SAR images
TABLE V: Comparisons of different target features for MIM. We found many target features are unsuitable for the multiplicative speckle noise in SAR images. This result inspired us to discuss different gradient features for SAR SSL. The pre-training setting SSL-SAR and the model backbones are the base version. The classification metric is average accuracy. Bold indicates the best result, underline is the next best result.
Model Target feature Classification (N-shot)
5 10 20
ViT pixel value [27] 54.1 61.5 68.2
ViT low pass filter [46] 53.8 60.5 66.7
ViT HOG feature [79] 39.7 48.7 56.6
ViT deep feature [2] 27.8 35.6 41.7
HiViT pixel value [27] 64.9 72.7 79.9
HiViT HOG feature [79] 58.2 64.8 71.7
HiViT SAR-HOG [61] 75.1 80.2 83.9
HiViT MGF 76.0 81.1 84.5
After discussing the backbone and strategy, we focus on the target features for SSL methods with SAR images. Due to the unique multiplicative speckle noise in SAR images, a key point for MIM is designing high-quality guide signals. As shown in Table V, we consider five target features (pixel value [27], low pass filter [46], HOG feature [79], deep feature [2], SAR-HOG [61], and gradient by ratio [16, 61]). All SSL methods use the SSL-SAR setting and base version.
First, we consider whether the existing methods based on the ViT are suitable for SAR images. PixMIM [46] applies the low pass filter to remove the high-frequency components and drives the model to focus on shape information. However, PixMIM does not outperform MAE because the noise type of SAR is multiplicative, and the filter parameters require a trade-off between target and noise. Then, FG-MAE [79] uses HOG to capture the feature for the SSL with SAR scene-level tasks, but we find that HOG does not ensure high-quality SAR target features. Target regions usually have strong scattering values, and the speckle noise causes the gradient computation to have strong false points in these regions. Besides, I-JEPA [2] propose deep networks as target feature encoders to capture the deep semantic features but lead to training overfitting noise and failure to learn effective feature representations.
Therefore, we choose the SAR features as target features to enhance HiViT. SAR-HOG changes the gradient calculation of HOG features and uses the gradient by ratio to solve speckle noise, and its result performs better than pixel value and HOG. Inspired by PixMIM, we prefer to use the target shape (i.e., gradient features) directly as the target feature  [^2] SAR-HOG uses the same multi-scale setting to illustrate that simple gradient features can effectively represent target shape as an SSL guide signal. Besides, multi-scale can improve the feature representation for various small targets in remote sensing. We discuss kernel settings for computing gradient in Fig. 4. Scale can affect feature quality: a smaller scale is finer for small target edge extraction, while a larger scale is more suitable for large target and noise suppression. Therefore, combining features of different scales has an improvement over a single scale for various target sizes in images.
Figure 4: Kernel settings for computing gradient by ratio. Here, the scale 1/2/3 is with $r$ equal to 9/13/17, and the multi-scale contacts all scales. Multi-scale is more suitable than single scale for various targets in remote sensing images.
4.4 Analysis
As stated above, SARATR-X’s key points are summarized as follows: HiViT architecture avoids the loss of small target information. SSL-ImageNet & SAR use ImageNet pre-training weights to provide a good initialization for diversity perceptual capability. MGF ensures high-quality target features and suppresses speckle noise under SSL with SAR images. Benefiting the above insights, our SARATR-X can learn high-quality target features from noisy SAR remote sensing images in Table IV. Next, we analyze the diversity and scalability of SARATR-X.
Figure 5: Averaged attention distance in different attention heads (x-axis is the attention head w.r.t layer number) on SSL models. Attention distance represents the range of the receptive field. We can focus on model architectures (Fig. (a) v.s. Fig. (b)), initialization weights (Fig. (a) v.s. Fig. (c)), and SSL signals (Fig. (d) v.s. Fig. (e)) to ensure diverse attention ranges in SAR target recognition, such as HiViT architecture, ImageNet weights, and SAR target features.
Figure 6: Scaling ability of SARATR-X in dataset size, parameters, and training epochs with linear probing performance. Our method can benefit from these three aspects. However, it is important to note that excessive training epochs can lead to overfitting due to the dataset size and speckle noise.
Visualization. Research [85] has shown that supervised pre-training and contrastive learning only model global information in the high layer, and MIM can model both local and global information. However, we observe that this phenomenon is not only related to the methods but also to the data properties. Fig. 5 (a) shows that the ViT with MAE focuses on global information due to large SAR image scenes, which differs from MIM’s modeling properties. Therefore, HiViT has various attention ranges with its high spatial resolution and hierarchical structure in Fig. 5 (b). In addition, using ImageNet’s weight as initialization can enhance this problem in Fig. 5 (c). As for the target feature, Fig. 5 (d) displays the HOG feature enhances the noise interference, which harms feature diversity. MGF effectively extracts the shape information of the target, making the model more focused on diverse edge information in the lower layer. However, this approach removes texture and preserves edge information, which motivates the higher layers not to need to focus on texture details and diminishes the attention range in Fig. 5 (e). Therefore, we combine two-step training with MGF in Fig. 5 (f).
Scaling experiment. Although MIM is a good learner that scales with data and model resources [86], a question arises as to whether our method can ensure scalability for MIM when dealing with noisy data such as SAR. Fig. 6 presents the scaling experiment from three perspectives: dataset size, parameters, and training epochs. Despite our pre-training set comprising 186,660 images, which is smaller than ImageNet-1K, we observe a significant improvement in downstream tasks’ performance with increasing data and parameters. This phenomenon indicates that the foundational model can fully unleash its potential in SAR images by extracting high-quality features as guiding signals. However, just like the study [86], when the pre-training size contains 100,000 images, the model tends to overfit during extended training epochs. Moreover, SAR image noise and low resolution can further aggravate the overfitting. However, SARATR-X outperforms our previous study, SAR-JEPA, which has overfitted at 400 epochs with 94,776 SAR images. There is a need to continue investigating how to ensure high-quality feature representation when extending the SAR foundation models.
5 Leveraging SARATR-X for Recognition
We have discussed different aspects of SARATR-X, but there are many specific datasets and specialized models for SAR ATR. We thus compare our SARATR-X with other state-of-the-art methods, such as supervised learning ($\rm{CS}^{n}$et [9] and PD [95]), semi-supervised learning (EUAPS [100]), and self-supervised learning (MSFA [43] and BIDFC [92]). We focus on SAR recognition tasks, including image classification and detection. More detailed settings [^3] We removed downstream tasks’ test sets from the pre-training set samples. are in appendix B.
TABLE VI: SAR classification results on MSTAR’s SOCs and EOCs settings. SOCs are in a similar distribution and have ten types of vehicle targets. EOCs include variations for imaging conditions that belong to robustness classification and out-of-distribution problems. EOCs use depression angle, configuration, and version variations to test robustness. We find that self-supervised (BIDFC and Ours) and semi-supervised (EUAPS) methods significantly perform under the few-shot task with additional unlabeled data, demonstrating the data’s importance. Besides, our method is robust to variations in imaging parameters. The metric is $\uparrow$, and the detail setting follows the $N$-shot [94]. Bold indicates the best result, underline is the next best result. We provide detailed results in Appendix B.
SOCs-Standard operating conditions (10-way)
Method Year 1-shot 2-shot 5-shot
DKTS-N [94] 2021 49.3 58.5 72.3
ConvT [72] 2022 42.6 54.4 75.2
HDLM [74] 2022 - - 72.4
BIDFC [92] 2022 80.7 85.3 90.3
CRID [73] 2023 48.3 51.0 73.3
EUAPS [100] 2023 - - 88.7
PD [95] 2024 46.7 58.9 70.2
SARATR-X 2024 85.2 (+4.5) 91.4 (+6.1) 95.9 (+5.6)
EOCs-Depression angle variations (4-way)
Method Year 1-shot 2-shot 5-shot
DKTS-N [94] 2021 61.9 63.9 67.4
ConvT [72] 2022 59.6 64.1 68.2
CRID [73] 2023 62.1 62.3 74.5
SARATR-X 2024 93.4 (+31.3) 97.3 (+33.2) 98.9 (+24.8)
EOCs-Target configuration variations (4-way)
Method Year 1-shot 2-shot 5-shot
DKTS-N [94] 2021 47.3 53.6 62.2
ConvT [72] 2022 44.3 51.9 64.1
CRID [73] 2023 62.8 65.7 74.1
SARATR-X 2024 65.0 (+2.2) 74.0 (+8.3) 78.3 (+4.2)
EOCs-Target version variations (4-way)
Method Year 1-shot 2-shot 5-shot
DKTS-N [94] 2021 48.9 55.1 65.6
ConvT [72] 2022 42.3 58.3 68.1
CRID [73] 2023 53.5 56.2 67.2
SARATR-X 2024 65.3 (+11.8) 76.5 (+20.3) 82.8 (+15.6)
TABLE VII: SAR detection results on SARDet-100K, OGSOD, SSDD, and SAR-Aircraft. Our proposed SARATR-X has competitive performance on various detection datasets. Bold indicates the best result, underline is the next best result. More detailed results are in Appendix B.
SARDet-100K (Object detection)
Method Year $\rm{mAP}\uparrow$ $\rm{mAP_{50}}\uparrow$ $\rm{mAP_{75}}\uparrow$
Deformable DETR [104] 2020 50.0 85.1 51.7
Swin Transformer [47] 2021 53.8 87.8 59.0
VAN [25] 2022 53.5 86.8 58.0
ConvNext [48] 2022 55.1 87.8 59.5
MSFA [43] 2024 56.4 88.2 61.5
SARATR-X 2024 57.3 (+0.9) 88.7 (+0.5) 62.8 (+1.3)
OGSOD (Object detection)
Method Year $\rm{mAP}\uparrow$ $\rm{mAP_{50}}\uparrow$ $\rm{mAP_{75}}\uparrow$
Generalized Focal [28] 2019 41.8 67.6 -
Sparse R-CNN [63] 2021 38.7 65.6 -
Object Box [91] 2022 40.1 76.6 -
YOLOv7 [75] 2022 45.1 79.2 -
SARATR-X 2024 52.0 (+6.9) 85.9 (+6.7) 51.3
SSDD (Ship detection)
Method Year $\rm{AP}\uparrow$ $\rm{AP_{50}}\uparrow$ $\rm{AP_{75}}\uparrow$
FBR-Net [20] 2021 - 94.1 59.1
CenterNet++ [24] 2021 - 95.1 -
CRTransSar [84] 2022 - 97.0 76.2
YOLO-Lite [59] 2023 - 94.4 -
FEPS-Net [3] 2023 59.9 96.0 67.5
$\rm{CS}^{n}$et [9] 2023 64.9 97.1 -
SARATR-X 2024 67.5 (+2.6) 97.3 (+0.2) 83.5 (+7.3)
SAR-Aircraft (Aircraft detection)
Method Year $\rm{mAP}\uparrow$ $\rm{mAP_{50}}\uparrow$ $\rm{mAP_{75}}\uparrow$
Cascade R-CNN [7] 2018 - 75.7 58.9
RepPoints [89] 2019 - 72.6 53.3
SKG-Net [21] 2021 - 70.7 46.4
SA-Net [80] 2023 - 77.7 62.8
SARATR-X 2024 58.7 86.1 (+5.7) 64.7 (+3.3)
Classification task. We study in Table VI the performance of SARATR-X on the MSTAR [1] dataset with Standard Operating Conditions (SOCs) and Extended Operating Conditions (EOCs). We first notice that SSL (BIDFC and ours) and semi-supervised (EUAPS) significantly outperform other methods under small samples with additional unlabeled samples. Our results surpass the previous best by large margins, demonstrating the value of foundation models in an era of rapidly growing SAR data. In particular, our SARATR-X shows robustness to the EOC setting of imaging condition variations. This result indicates that the foundation model learns stable features and relationships from diverse imaging conditions in a large number of target samples.
Detection task. As illustrated in Table VII, we report box AP for SAR target detection with a horizontal bounding box on multi-category (SARDet-100K and OGSOD), ship (SSDD), and aircraft detection (SAR-Aircraft). SARATR-X performs better than our previous MSFA by 0.8 points on SARDet-100K. MSFA has more complex training processes and target features, which use multi-stage training between RGB and SAR images with three different target features and have a detection pre-training step. SARATR-X is more simple but more effective for SAR images. More significantly, SARATR-X outperforms or has comparable performance on various datasets compared to many specifically designed detection methods in Table VII. Of course, our study is only a preliminary exploration of SSL for SAR. More effective target features can be realized in a data-knowledge dual-driven manner by further mining the knowledge of SAR imaging mechanisms and properties. Moreover, with the larger dataset and parameters, the path of foundation models will hopefully lead to general SAR target recognition, but this goal requires researchers together.
6 Conclusion and Future Perspectives
In this paper, we proposed the SARATR-X for SAR ATR and systematically investigated a foundation model framework. First, a pre-training dataset was built from 14 open-source datasets, including various targets, scenes, and sensors. Then, the foundation model’s pre-training backbone, SSL methods, and downsteam tasks were discussed in detail. Importantly, SARATR-X demonstrated superior performance on different target recognition datasets, demonstrating the foundation model’s potential in this field.
We believe that research on SAR foundation models, such as SARATR-X, has the potential for generalized feature representations in SAR images and accelerates towards all-day, all-weather target recognition in Earth observation. However, the research of the foundation model requires large data, which is a real problem for SAR images. SAR images are expensive and require specific imaging equipment and algorithms, and privacy and security prevent the data from opening. Therefore, we are particularly grateful to the publishers of open-source SAR target datasets. By making SARATR-X publicly available, we aim to accelerate the progress of the foundation model in SAR target recognition by enabling researchers to use our dataset and code to design better methods or explore downstream applications.
Although this work systematically investigated a foundation model framework, several limitations and challenges require exploration in future work. The SAR images are derived from open-source SAR datasets, and the targets are mainly vehicles, ships, aircraft, oil tanks, etc. Collecting target samples from increasingly unlabelled SAR imagery could further expand the amount of data and the range of downstream applications. In addition, investigating the expert knowledge with text for multimodal interactions to describe the relationship between targets and scenes can further enhance the representation capability of the foundation model. Due to the visual variability of SAR images, textual descriptions are an effective tool for improving interpretability and comprehensibility.
In conclusion, we have verified the SARATR-X’s ability to adapt to diverse SAR target datasets, showing high performance and generalizability in classification and detection. By taking full advantage of the rapid growth of SAR images, the SSL-based foundation model opens the door to a generalized feature representation and SAR target recognition. We believe it’s time to embrace fundamental models for SAR image interpretation.
Appendix A Implementation Details of Section 4 Experiments of SARATR-X
Here are the details of the dataset and training settings.
A.1 Pre-traing dataset setting
Figure 7: Visualization of the fourteen datasets included in our pre-training. In this way, a pre-trained dataset for SAR ATR can be built with multiple targets, scenes, and sensors.
As shown in Fig. 7, we collect data from open-source datasets based on our previous research [38, 43]. Now, our pre-training dataset contains 14 open-source SAR target datasets. Here are brief descriptions of each dataset’s targets, scenes, and sensors.
AIR-SARShip [66] is a ship detection dataset based on the Chinese C-band Gaofen-3 satellite. AIR-SARShip-1.0 and AIR-SARShip-2.0 include 318 VV-polarised images with 1 and 3 m resolutions. This dataset includes harbors, islands, and different conditions of sea surfaces and covers thousands of ships.
HRSID [81] is a high-resolution dataset for ship detection and instance segmentation based on the European C-band Sentinel-1B, German X-band TerraSAR-X and TanDEM-X satellites. HRSID consists of 5,604 cropped SAR images with 0.5 to 3 m resolutions. The scene is a busy area of maritime transport, such as harbors and estuarine cities, and the annotation targets are ships of different sizes.
Sandia MiniSAR [60] is a 0.1 m resolution dataset based on a Ku-band airborne platform released by Sandia National Laboratories. The dataset contains scenes and targets such as aircraft on tarmacs, buildings in urban areas, and vehicles in desert areas but lacks official annotations.
MSAR [84, 10] is a multi-class target detection dataset based on the Chinese C-band HISEA-1 satellite in large-scale scenes. MSAR comprises 28,449 image slices with quad polarization and 1 m resolution. Scenes covered include airports, harbors, nearshore, islands, distant seas, and urban areas. The labeled target categories include aircraft, oil tanks, bridges, and ships.
MSTAR [1] is the most commonly used target classification dataset released by the Defense Advanced Research Projects Agency, USA. Its sensor is an X-band radar with HH polarization mode and 0.3 m resolution. It contains ten categories of military vehicles with various imaging angles, target variants, and other conditions but with single grass scenes.
OGSOD [71] is a city object detection dataset collected from the Chinese C-band Gaofen-3 satellite with VV and VH polarization modes, and its resolution is 3 m. This dataset also contains optical images from Google Earth with 10 m resolution. It is annotated with static objects, including bridges, harbors, and oil tanks in urban areas.
OpenSARShip [35] is a ship slices dataset based on the European C-band Sentinel-1 satellite. Its resolution is 2.3 m to 17.4 m with VV and VH polarization. The dataset contains many ship slices from 10 busy ports. It has a diverse range of ship types but a significant category imbalance.
SADD [96] is an aircraft detection dataset collected from the German X-band TerraSAR-X satellite. Its resolution is 0.5 m to 3 m with HH polarization. The dataset contains densely parked aircraft of different sizes on airport tarmacs and runways. It has a large number of small-sized planes as well as the airport perimeter area.
SAMPLE [34] is a synthetic and measured paired fine-grained vehicle dataset released by the Air Force Research Laboratory, USA. This dataset is simulated in X-band and 0.3 m resolution. The public version provides 5,380 images of ten categories of vehicle targets at partial imaging angles.
SAR-AIRcraft [80] is a aircraft detection dataset based on the Chinese C-band Gaofen-3 satellite with 1 m resolution and single polarization. The dataset collects seven types of aircraft of different sizes from three civil airports. It can support fine-grained aircraft detection and classification studies.
SARSim [51, 33] is a fine-grained vehicle dataset created by Terma A/S, Denmark. The simulation system used for this dataset can generate X-band SAR images with resolutions ranging from 0.1m to 0.3m from CAD models. SARSim provides 21,168 vehicle samples in 7 categories (truck, car, motorbike, bus, tank, bulldozer, and pickup) and 3 scenes (grass, roads, and a mean of the two) with 7 imaging depression angles.
SAR-Ship [78] is a ship target detection dataset in complex scenes based on Chinese Gaofen-3 and European Sentinel-1 satellites. The public version of this dataset contains 39,729 images from two satellites in different imaging modes and resolutions. The dataset provides ship targets of various sizes in complex ocean scenes such as nearshore, distant seas, harbors, and islands.
SIVED [44] is a vehicle detection dataset with rotatable bounding box. It consists of vehicle slices from the MSTAR dataset [1] and vehicles in urban areas from the Sandia MiniSAR and FARAD datasets [60], and scenes include car parks, buildings, trees, roads, and others.
SSDD [97] is a commonly used SAR ship detection dataset. It is constructed based on Canadian RadarSat-2, German TerraSAR-X, and European Sentinel-1 satellites and contains different scenarios for the inshore and offshore of China and India. The dataset covers various ship sizes in different oceanic conditions with diverse clutter and noise interference.
A.2 Classification dataset for performance Test
TABLE VIII: Description of our SAR classification dataset, named SAR-VSA, which contains 25 fine-grained targets. # Train: Number of train samples. # Test: Number of test sample images.
Fine-grained category # Train # Test
anti-aircraft (ZSU234) 299 274
bulldozer (D7) 299 274
howitzer (2S1) 299 274
infantry vehicle (BMP2) 698 587
main battle tank (T62) 299 273
main battle tank (T72) 691 582
patrol car (BRDM2) 298 274
personnel carrier (BTR60) 256 195
personnel carrier (BTR70) 233 196
truck (ZIL131) 299 274
bridge 1,023 438
coastal land 707 303
land patch 1,137 487
sea clutter wave 1,378 590
sea patch 1,250 535
ship (cargo) 366 156
ship (fishing) 248 106
ship (tanker) 150 64
ship (others) 312 133
strong false alarms 299 128
aircraft (Airbus A220) 91 373
aircraft (Airbus A330) 97 415
aircraft (Comac ARJ21) 103 411
aircraft (Boeing 737) 100 428
aircraft (Boeing 787) 113 391
We select three target classification datasets, including 25 fine-grained targets from vehicles, ships, aircraft, and others, to evaluate the comprehensive performance of SSL and the foundation model for SAR target recognition. The new SAR classification dataset named SAR-Target in Table VIII.
MSTAR [1] is the most commonly used SAR vehicle datasetw. It has many experimental setting variants, while we refer to the  [40] to adopt the most commonly used ten-class classification settings, such as infantry vehicle, patrol car, personnel carrier, main battle tank, and truck.
FUSAR-Ship [64] contains 15 primary ship categories and many non-ship targets based on the Gaofen-3 satellite in scenes such as sea, land, coast, river, and island. Based on the experimental setting of [76], we have ten ocean target types, such as four fine-grained ships, bridges, and ocean scene slices.
SAR-ACD [64] contains five types of aircraft based on the Gaofen-3 satellite in three civil airports. Since the released dataset does not separate the training and test data, we randomly select partial samples as the training set and others as the test set. Fine-grained recognition of aircraft targets is a more challenging task due to the smooth surface of the aircraft, resulting in insignificant SAR image features.
A.3 Hyperparameter settings
Here are the detailed settings of our pre-training and downstream tasks as shown in Table IX and X.
Pre-training. Our default pre-training setting is in Table IX, and other hyperparameters of each method use the default settings from their papers and codes. Our pre-training is applied on 8 NVIDIA RTX3090 GPUs with 200 epochs and 800 batch sizes. Compared to the training settings of MAE, we add ColorJitter ($\rm{contrast}=0.5$) to increase data richness. Moreover, we modify the batch size and epoch according to 8 GPUs. It is worth noting that although MAE uses the normalized pixel value to enhance the feature representation in the visible spectral images, we find that the normalized pixel value cannot be used due to the SAR image noise and prevents the training loss from decreasing properly.
TABLE IX: Pre-training setting.
Config Value
optimizer AdamW [50]
base learning rate 1.5e-4
weight decay 0.05
optimizer momentum $\beta_{1},\beta_{2}=0.9,0.95$ [12]
batch size 800
epoch 200
learning rate schedule cosine decay [49]
warmup epoch [23] 5
augmentation ResizedCrop, HFlip, ColorJitter
Classification setting. All models use the same training settings in downstream classification tasks. Table X gives the default setting. Our few-shot learning setting is based on the Dassl toolbox [101, 102] and is averaged over 10 random experiments. Since we focus on the small sample case of the downstream classification task, we use the linear probing method in MAE to finetune models and avoid overfitting.
TABLE X: Classification settings (fine-tuning and linear probing).
Config Value
optimizer AdamW
base learning rate 1e-3
weight decay 1e-4
optimizer momentum $\beta_{1},\beta_{2}=0.9,0.95$
batch size 25
epoch 30
learning rate schedule cosine decay
warmup epoch 1
warmup type constant
warmup learning rate 1e-5
Figure 8: Partial fine-tuning results of HiViT-B and SSL-ImageNet & SAR. Tuning 0 blocks is linear probing. With the increasing number of fine-tuned transformer blocks, our method are consistently better than MAE and experience overfitting later.
Partial fine-tuning. Fig. 8 shows why we chose linear probing for few-shot evaluation, as HiViT overfits when many blocks are fine-tuned. Experimental results show that our SAR-KGPA consistently obtains better representations than MAE, and overfitting occurs later.
Appendix B Implementation Details of Section 5 Leveraging SARATR-X for Recognition
B.1 Dataset description
We choose MSTAR, the most commonly used dataset in SAR target classification. SOCs are in similar image conditions, and the training set’s depression angle under SOC is 17°, while the test set is 15°. Ten categories of targets include BMP2, BRDM2, BTR60, BTR70, T62, T72, 2S1, D7, ZIL131, and ZSU234 as shown in Table VIII. The EOCs setting is the imaging condition variations to test robustness. Existing methods have been saturated with different experimental settings of MSTAR [40], but few samples and depression angle variations remain challenging. We use SARDet-100K and OGSOD datasets, which have many test samples and categories, to evaluate the detection performance fully. The OGSOD comparison results are derived from the original article’s [71] single-modal approach using only SAR images. SSDD and SAR-Aircraft are ship and aircraft categories.
B.2 Hyperparameter settings
Based on our scaling experiment, we use HiViT-B with 600 epochs pre-trained on SSL-ImageNet & SAR as the foundation model for classification and detection tasks.
Classification setting is follow Table X. The only difference is that we use partial fine-tuning for better performance, and the last 6 blocks are added to the fine-tuned.
Detection setting is follows the default setting in HiViT, and we adjust the learning rate to 5e-4. We used the same settings for each dataset fine-tuning, see our GitHub configuration based on the mmdetection [11] framework for details.
B.3 Detailed results
We provide detail classification and detection result on Table XI and XII. Although the proposed method outperforms existing methods in mAP, there is still scope for improvement in some refined metrics.
TABLE XI: Detailed classification results on MSTAR’s SOCs and EOCs. Bold indicates the best result, underline is the next best result.
SOCs: Standard operating conditions (10 way)
Method Year 1-shot 2-shot 3-shot 4-shot 5-shot 10-shot 20-shot 25-shot
DKTS-N [94] 2021 49.3 58.5 - - 72.3 84.6 - 96.2
ConvT [72] 2022 42.6 54.4 - - 75.2 88.6 - 96.5
HDLM [74] 2022 - - - - 72.4 88.2 95.2 -
BIDFC [92] 2022 80.7 85.3 87.3 88.4 90.3 - - -
CRID [73] 2023 48.3 51.0 - - 73.3 87.4 96.9 97.1
EUAPS [100] 2023 - - - - 88.7 98.6 99.8 -
PD [95] 2024 46.7 58.9 62.0 67.5 70.2 83.7 - -
SARATR-X 2024 85.2 91.4 93.9 95.1 95.9 97.7 98.1 98.5
EOCs: Depression angle variations (4 way)
Method Year 1-shot 2-shot 3-shot 4-shot 5-shot 10-shot 20-shot 25-shot
DKTS-N [94] 2021 61.9 63.9 - - 67.4 71.1 - 78.9
ConvT [72] 2022 59.6 64.1 - - 68.2 74.8 79.1
CRID [73] 2023 62.1 62.3 - - 74.5 85.9 - 87.0
SARATR-X 2024 93.4 97.3 98.5 98.0 98.9 99.5 99.6 99.4
EOCs: Target configuration variations (4-way)
Method Year 1-shot 2-shot 3-shot 4-shot 5-shot 10-shot 20-shot 25-shot
DKTS-N [94] 2021 47.3 53.6 - - 62.2 68.4 - 74.5
ConvT [72] 2022 44.3 51.9 - - 64.1 89.7 - 91.0
CRID [73] 2023 62.8 65.7 - - 74.1 78.7 - 84.1
SARATR-X 2024 65.0 74.0 74.0 84.2 78.3 87.0 88.9 93.5
EOCs: Target version configuration variations (4-way)
Method Year 1-shot 2-shot 3-shot 4-shot 5-shot 10-shot 20-shot 25-shot
DKTS-N [94] 2021 48.9 55.1 - - 65.6 70.2 - 77.0
ConvT [72] 2022 42.3 58.3 - - 68.1 83.6 - 92.0
CRID [73] 2023 53.5 56.2 - - 67.2 79.9 - 87.8
SARATR-X 2024 65.3 76.5 76.3 83.4 82.8 85.0 92.1 97.0
TABLE XII: Detailed detection results. The metrics are mAP, $\rm{mAP_{50}}$ (@@@@50), $\rm{mAP_{75}}$ (@@@@75), $\rm{mAP_{small}}$ (@@@@s), $\rm{mAP_{medium}}$ (@@@@m), and $\rm{mAP_{large}}$ (@@@@l). Bold indicates the best result, underline is the next best result.
SARDet-100K (Object detection)
Method Year mAP @@@@50 @@@@75 @@@@s @@@@m @@@@l
Deformable DETR [104] 2020 50.0 85.1 51.7 44.0 65.1 61.2
Swin Transformer [47] 2021 53.8 87.8 59.0 49.1 64.6 60.0
VAN [25] 2022 53.5 86.8 58.0 47.3 65.5 60.6
ConvNext [48] 2022 55.1 87.8 59.5 48.9 66.9 61.1
MSFA [43] 2024 56.4 88.2 61.5 50.5 66.5 62.5
SARATR-X 2024 57.2 (+0.8) 88.5 (+0.3) 62.6 (+1.1) 51.2 (+0.7) 69.6 (+3.1) 64.7 (+2.2)
OGSOD* (Object detection)
Method Year mAP @@@@50 @@@@75 @@@@s @@@@m @@@@l
Generalized Focal [28] 2019 41.8 67.6 - - - -
Sparse R-CNN [63] 2021 38.7 65.6 - - - -
Object Box [91] 2022 40.1 76.6 - - - -
YOLOv7 [75] 2022 45.1 79.2 - - - -
SARATR-X 2024 51.4 (+6.3) 85.2 (+6.0) 51.7 46.8 75.3 77.1
SSDD (Ship detection)
Method Year AP @@@@50 @@@@75 @@@@s @@@@m @@@@l
FBR-Net [20] 2021 - 94.1 59.1 - - -
CenterNet++ [24] 2021 - 95.1 - - - -
CRTransSar [84] 2022 - 97.0 76.2 - - -
YOLO-Lite [59] 2023 - 94.4 - - - -
FEPS-Net [3] 2023 59.9 96.0 67.5 55.1 68.2 70.6
$\rm{CS}^{n}$et [9] 2023 64.9 97.1 - - - -
SARATR-X 2024 67.4 (+2.5) 97.3 (+0.2) 83.1 (+6.9) 68.1 (+13.0) 66.1 (-2.1) 60.5 (-10.1)
SAR-Aircraft (Aircraft detection)
Method Year AP @@@@50 @@@@75 @@@@s @@@@m @@@@l
Cascade R-CNN [7] 2018 - 75.7 58.9 - - -
RepPoints [89] 2019 - 72.6 53.3 - - -
SKG-Net [21] 2021 - 70.7 46.4 - - -
SA-Net [80] 2023 - 77.7 62.8 - - -
SARATR-X 2024 58.7 86.1 (+5.7) 64.7 (+3.3) 20.0 57.2 49.7

### v2

@@
-SARATR-X: A Foundation Model for Synthetic Aperture Radar Images Target Recognition
+SARatrX: Towards Building A Foundation Model for SAR Target Recognition
 Abstract
-Synthetic aperture radar (SAR) is essential in actively acquiring information for Earth observation.
-SAR Automatic Target Recognition (ATR) focuses on detecting and classifying various target categories under different image conditions.
-The current deep learning-based SAR ATR methods are typically designed for specific datasets and applications.
-Various target characteristics, scene background information, and sensor parameters across ATR datasets challenge the generalization of those methods.
-This paper aims to achieve general SAR ATR based on a foundation model with Self-Supervised Learning (SSL).
-Our motivation is to break through the specific dataset and condition limitations and obtain universal perceptual capabilities across the target, scene, and sensor.
-A foundation model named SARATR-X is proposed with the following four aspects: pre-training dataset, model backbone, SSL, and evaluation task.
-First, we integrated 14 datasets with various target categories and imaging conditions as a pre-training dataset.
-Second, different model backbones were discussed to find the most suitable approaches for remote-sensing images.
-Third, we applied two-stage training and SAR gradient features to ensure the diversity and scalability of SARATR-X.
-Finally, SARATR-X has achieved competitive and superior performance on 5 datasets with 8 task settings, which shows that the foundation model can achieve universal SAR ATR.
-We believe it’s time to embrace fundamental models for SAR image interpretation in the era of increasing big data.
+Despite the remarkable progress in synthetic aperture radar automatic target recognition (SAR ATR), recent efforts have concentrated on the detection or classification of a specific and coarse category, e.g., vehicles, ships, airplanes, or buildings.
+One of the fundamental limitations of the top-performing SAR ATR methods is that the learning paradigm is supervised, task-specific, limited-category, closed-world learning, which depends on massive amounts of accurately annotated samples that are expensively labeled by expert SAR analysts and has limited generalization capability and scalability.
+In this work, we make the first attempt towards building a foundation model for SAR ATR, termed SARatrX.
+SARatrX learns generalizable representations via self-supervised learning (SSL) and provides a basis for label-efficient model adaptation to generic SAR target detection and classification tasks.
+Specifically, SARatrX is trained on 0.18 M unlabelled SAR target samples, which are curated by combining contemporary benchmarks and constitute the largest publicly available dataset till now.
+Considering the characteristics of SAR images, a backbone tailored for SAR ATR is carefully designed, and a two-step SSL method endowed with multi-scale gradient features was applied to ensure the feature diversity and model scalability of SARatrX.
+The capabilities of SARatrX are evaluated on classification under few-shot and robustness settings and detection across various categories and scenes, and impressive performance is achieved, often competitive with or even superior to prior fully supervised, semi-supervised, or self-supervised algorithms.
+Our SARatrX and the curated dataset are released at https://github.com/waterdisappear/SARatrX to foster research into foundation models for SAR ATR and SAR image interpretation.
 Index Terms:
-Synthetic Aperture Radar (SAR), Target Recognition, Foundation Model, Self-Supervised Learning (SSL), Deep Learning, Masked Image Modeling (MIM)
-1 Introduction
-Based on electromagnetic scattering in microwave frequency bands, Synthetic Aperture Radar (SAR) [62, 22, 53] is essential in actively acquiring information for Earth observation with stable imaging capacity under various weather and lighting conditions.
-SAR Automatic Target Recognition (ATR) aims to automatically localize and classify interested objects in SAR images, e.g., detection and classification.
-It has been investigated extensively with various civilian and military applications [69, 36, 73, 57].
-In the past decade, deep learning has reinvigorated SAR ATR with impressive breakthroughs [31, 37, 15].
-However, as shown in Fig 1 and 2, due to the extensive acquisition conditions as well as the high costs of collection and annotation, there are a plethora of special datasets and algorithms designed for different applications of SAR ATR.
-For example, nearly a dozen new target datasets and dozens of attention modules have been developed for SAR target recognition in the last few years.
-However, the target characteristics, scene background information and sensor parameters across datasets challenge their generalization.
-In contrast to specialized algorithms for particular applications, we aim to investigate a general SAR target recognition method.
-A foundation model [5] pre-trained in a task-agnostic manner (generally via self-supervision learning) on extensive data can be flexibly adapted to a wide range of downstream tasks under various conditions.
-Self-supervised learning (SSL) [45, 4] can mitigate label inefficiency by exploring supervision directly from the data itself, thereby reducing reliance on expensive expert labeling and efficiently scaling data and models.
-Hence, SSL has the capability to leverage extensive SAR unlabeled data for constructing foundational models, which can achieve universal recognition with a few labeled target samples in different categories and conditions.
-As shown in Table I, foundation models are thriving in remote sensing with superior performance in a diverse range of modalities and tasks.
-In particular, multi-modality foundation models have recently appeared, indicating this research topic’s great potential.
-But, there lacks a foundation model that can effectively accomplish various tasks for SAR ATR, such as FG-MAE [79] and SkySense [26] mainly focus on SAR scene classification and segmentation, and our previous SAR-JEPA [38] and MSFA [43] explored the object classification and detection tasks separately.
-It is still necessary to explore the development of an analogous ATR foundation model for SAR image interpretation, which lies at the intersection of SAR technologies and frontier AI technologies (in particular, the big models).
-Figure 1: Motivation and performance of SARATR-X.
-(a) SAR ATR has various imaging conditions, including targets, scenes, and sensors.
-However, due to high cost, its datasets are collected under specific settings and tasks.
-For example, the MSTAR dataset [1] is a ten-type vehicle target classification dataset in X-band and grass scenarios, while SAR-Aircraft is a seven-type aircraft target detection dataset in three airports and a C-band satellite.
-Many specialized algorithms have been proposed for these datasets.
-However, the different target characteristics, scene background information and sensor parameters between these datasets challenge the generalization of existing algorithms.
-This paper aims to explore a SAR ATR foundation model, a general method to achieve various tasks.
-(b) Our SARATR-X performs well on 5 datasets with 8 settings compared with previous methods.
-For target classification, SARATR-X is better than the previous SSL method (BIDFC [92]) on fine-grained vehicle MSTAR dataset [1] with the few-shot setting.
-Besides, it performs well under extended operating conditions (EOCs) [94], i.e., the imaging condition variation in depression angle (EOCs-Depression), target configuration (EOCs-Config), and version (EOCs-Version).
-For object detection, SARATR-X demonstrates competitive performance with previous carefully designed supervised methods on various categories (SARDet-100K [43] and OGSOD [71]) and specific categories on ship (SSDD [97]), and aircraft (SAR-AIRcraft [80]).
-Our study shows the potential of the foundation model for universal SAR target recognition.
-Thanks to the development of SAR sensors and the enormous efforts of researchers, numerous SAR target datasets have emerged over the past five years in Fig.
-2.
-Despite image annotation remaining a challenging and expensive endeavor, SAR sensors have produced vast amounts of data at a significantly accelerated pace compared to the previous ten years.
-Meanwhile, the investigation of SSL and foundation models for SAR ATR is advancing with increasing datasets.
-As shown in Table I, existing research [92, 79, 56, 38, 43] have discussed different factors of a foundation model, and these studies are progressively advancing towards a goal of achieving general SAR ATR.
-For example, BIDFC [92] performs much better than supervised methods in the few-shot classification of the MSTAR vehicle dataset, SAR-JEPA [38] demonstrates that SSL can effectively improve the classification accuracy for different targets across datasets, and MSFA [43] shows improved results of ship detection.
-Inspired by their insightful work, we formally propose a SAR ATR foundation model, SARATR-X, to achieve a breakthrough in universal SAR target recognition.
-In this work, we lay the groundwork for developing SARATR-X, a foundation model specifically tailored for SAR ATR tasks.
-It aims to learn generalizable representations from unlabeled SAR images, offering a basis for efficient adaptation across various downstream ATR tasks.
-The four parts for creating SARATR-X include pre-training datasets, model backbones, self-supervised learning, and evaluation tasks.
-Pre-training datasets need to contain diverse target categories and imaging conditions to accommodate various downstream tasks.
-However, SAR ATR lacks a large-scale dataset like ImageNet [19], and the most used MSTAR dataset in Table II only includes fine-grained vehicle categories, which is not suitable for larger-scale pre-training.
-Therefore, many SSL methods combine different datasets; in particular, SARDet-100K incorporates [43] 9 SAR target detection datasets.
-With the increasing number of open-source datasets for SAR ATR, we integrate the vast majority of them for pre-training.
-As shown in Fig 2 and Table III, we apply 14 classification and detection datasets with different target categories and imaging conditions to explore the foundation model’s potential.
-Model backbones aim to achieve better spatial representation in remote sensing images, especially small target signatures in large imagery.
-As shown in Table I, researchers used two architectures (Transformer or convolutional neural network).
-Transformer has better spatial resolution without downsampling, and Convolutional Neural Network (CNN) has higher efficiency with its convolution kernel.
-MSFA discussed the effectiveness of different architectures, and the results showed that the Swin Transformer backbone performs better.
-Therefore, we discuss different model backbones to find the most suitable approaches for the properties of remote sensing images.
-HiViT [99] is our choice, which combines Swin Transformer advantages and can drop patches in MIM.
-Self-supervised learning faces the challenge of SAR image quality.
-SAR images contain speckle noise due to coherent imaging, and their visual features are not as distinct and rich as RGB images.
-As shown in Table II, contrastive learning [92, 56] uses data augmentation and pre-processing to reduce noise, while MIM [79, 38, 43] has a pixel-to-feature improvement for guide signals.
-Therefore, we consider that the problem to be solved by SAR SSL is how to construct high-quality guide signals.
-For example, PGIL [29] leveraged a sub-frequency feature of SAR complex images to learn physics information, and our SAR-JEPA [38] applied multi-scale gradient ratio to solve multiplicative speckle noise and capture target shape.
-Furthermore, multi-stage training [43] from ImageNet to SAR diminished noise interference on model diversity in Fig.
-5.
-Therefore, we apply two-stage training from ImageNet to SAR and use multi-scale gradient features as high-quality guide signals for SAR MIM.
-Evaluation tasks need to comprehensively evaluate the performance of the foundation model under different tasks and settings.
-Benefiting from the open-source target datasets, we first construct a fine-grained classification dataset with various categories to evaluate the effectiveness of the proposed improvements.
-In the end, we perform a comprehensive comparison between the proposed SARATR-X and existing SOTA methods in public classification and detection tasks.
-Our SARATR-X has achieved a competitive and superior performance on 5 datasets with 8 task settings in Fig 1.
-And its codes and weights can be found at https://github.com/waterdisappear/SARATR-X.
-We hope this work will advance the development of foundation models and general SAR target recognition.
-The main contributions are summarized as follows:
-TABLE I: List of visual foundation models in remote sensing.
-There are many foundation models for various modalities and tasks, but research (FG-MAE, SkySense, and OFA-Net) focuses on scene-level tasks with the SAR modality.
-Based on our previous study SAR-JEPA and MSFA [38, 43], SARATR-X fills this gap in the foundation model for SAR ATR.
-These foundation models’ SSL methods are divided into two approaches: one is contrastive, such as contrastive learning, to obtain invariant features; another is generative, such as masked image modeling, to generate and predict features.
-SD: Stable Diffusion.
-Swin: Swin Transformer.
-Foundation model Year Modality Dataset Backbone SSL Tasks
-SatMAE [13] 2022 Multi-spectral, RGB fMoW RGB/Sentinel ViT Generative Scene classification and semantic segmentation
-RVSA [77] 2022 RGB MillionAID ViT, ViTAE Generative Scene classification, object detection, and semantic segmentation
-RingMo [65] 2022 RGB Self-built ViT, Swin Generative Scene classification, object detection, semantic segmentation, and change detectionn
-RingMo-Sense [90] 2023 RGB Self-built Video Swin Generative object detection and tracking, images segmentation, and 3 prediction tasks
-CMID [55] 2023 RGB MillionAID ResNet, Swin Generative & Contrastive Scene classification, object detection, and semantic segmentation
-GFM [52] 2023 RGB GeoPile Swin Generative Scene classification, semantic segmentation, change detection, and super-resolution
-DiffusionSat [32] 2023 RGB fMoW, Satlas, SpaceNet SD Generative Image generation, super-resolution, temporal generation, and in-painting
-Scale-MAE [58] 2023 RGB fMoW RBG ViT Generative Scene classification, object detection, and semantic segmentation
-FG-MAE [79] 2023 Multi-spectral, SAR SSL4EO-S12 ViT Generative Scene classification and semantic segmentation
-SMLFR [17] 2024 RGB GeoSense ConvNeXt Generative Object detection, semantic segmentation, and change detection
-SkySense [26] 2024 3 modalities HSROIs, TMsI, TSARI Swin, Vit Contrastive Scene classification, object detection, semantic segmentation, and change detection
-OFA-Net [87] 2024 4 modalities 5 multi-modalities datasets ViT Generative Scene Classification and semantic segmentation
-SAR-JEPA [38] 2024 SAR MSAR, SAR-Ship, SARSim, SAMPLE ViT Generative Target classification
-MSFA [43] 2024 SAR ImageNet, DOTA, SARDet-100k CNN, ViT, Swin Generative Object detection
-SARATR-X 2024 SAR ImageNet & 14 SAR datasets HiViT Generative Target classification and object detection
+Synthetic Aperture Radar (SAR), Object Recognition, Object Detection, Foundation Model, Self-supervised learning, Deep Learning, Masked Image Modeling
+https://github.com/waterdisappear/SARatrX
+I Introduction
+Synthetic aperture radar (SAR) [1, 2, 3, 4], developed from electromagnetic scattering in microwave frequency bands, plays a crucial role in active Earth observations, functioning effectively across diverse weather conditions and lighting environments.
+With the rapid development of SAR imaging techniques, high-resolution SAR images can be accessed more easily than before, enabling even more research opportunities for SAR image intelligent interpretation.
+SAR automatic target recognition (ATR), aiming at automatically localizing and classifying objects of interest (e.g., vehicles, ships, airplanes, or buildings) in SAR images, is a longstanding, important yet challenging problem in SAR image intelligent interpretation [5, 6, 7, 8].
+SAR ATR plays an essential role in civil and national defense applications such as modern airport management, disaster management, urban planning and infrastructure monitoring, military reconnaissance, and maritime surveillance.
+Therefore, it has become an active research area for several decades [9, 10, 11, 12, 13, 14].
+In the past decade, deep learning has brought tremendous success for SAR ATR [15, 5, 16, 17, 8].
+Despite the significant progress, there are at least the following fundamental challenges that will need to be addressed to further advance the field of SAR ATR.
+Figure 1: The motivation and performance of SARatrX.
+(a) SAR ATR includes various imaging conditions, targets, scenes, and sensors.
+However, due to high costs, the datasets are often collected in specific settings for certain tasks.
+For example, MSTAR [18] is a ten-type vehicle target classification dataset in the X-band and grass scenarios, while SAR-Aircraft is a seven-type aircraft target detection dataset collected from three airports and a C-band satellite.
+Several specialized algorithms have been proposed for application to these datasets.
+However, the differing target characteristics, scene background information, and sensor parameters have complicated the generalization of existing algorithms.
+As such, this paper aims to develop a SAR ATR foundation model, a generalized method for conducting various tasks.
+(b) SARatrX performed well across 5 datasets with 8 settings.
+It was superior to existing SSL methods (BIDFC [19]) for target classification in the fine-grained vehicle MSTAR dataset [18] with a few-shot setting.
+In addition, it performed well under extended operating conditions (EOCs) [20] (i.e., imaging conditions with variable depression angle (EOCs-Depression), target configuration (EOCs-Config), and version (EOCs-Version)).
+SARatrX also demonstrated competitive object detection performance with existing supervised methods applied to various categories (SARDet-100K [21] and OGSOD [22]), as well as specific categories for ships (SSDD [23]) and aircraft (SAR-AIRcraft [24]).
+Our study shows the potential of a foundation model for generalized SAR ATR.
+Firstly, task-specific property.
+One of the fundamental limitations of the current SAR ATR methods [25, 26, 27, 28, 29, 30] is that one model is trained and evaluated on one specific task.
+The detection and classification of a specific coarse category (e.g., vehicles, ships, airplanes, or buildings in Fig.
+1.) all require their own deep models.
+As a result, the task-specific properties of these deep models pose significant challenges for training new tasks or developing a comprehensive SAR ATR system since each task must be learned independently from the ground up, requiring vast amounts of labeled data.
+This results in computational inefficiency, lower accuracy, and inconsistent results between the different models.
+Secondly, heavy reliance on supervised learning.
+Recent progress in SAR ATR, while substantial, has been limited to supervised learning, which heavily depends on massive amounts of accurately annotated target samples that are expensively labeled by expert SAR analysts and have limited generalization capability and scalability.
+However, the scarcity of expert SAR analysts cannot meet such an exhaustive requirement, leaving vast amounts of SAR images unlabelled and unexploited.
+Thirdly, the ignorance of SAR image characteristics in model designs.
+The imaging characteristics of SAR imagery differ significantly from those of optical imagery, leading to a significant domain gap between natural and SAR images.
+This raises significant challenges when one intends to transfer prior knowledge from the natural image domain.
+Different strong prior knowledge of SAR imagery, including speckle noise, discrete target appearances, and the lack of geometry, texture, and contour cues, needs special consideration when designing backbone architectures.
+Most of the current mainstream backbones and methods designed on natural images are not suitable for the aforementioned information.
+Finally, underdeveloped open-source ecosystem.
+Due to data sensitivity, the open-source ecosystem across the entire field is underdeveloped, making it challenging to share code and data publicly.
+Currently, there are no large and representative benchmark datasets for SAR ATR.
+As a result, this locks the potential of recent deep learning techniques for SAR ATR and significantly slows down the development of this field.
+Recently, the remarkable success of foundation models (FMs) [31, 32, 33, 34] has led to a learning paradigm shift in artificial intelligence.
+Foundation models [35], pretrained on extensive data in a task-agnostic manner (generally via Self-supervised learning), can be flexibly adapted to a wide range of downstream tasks.
+Self-supervised learning (SSL) [36, 37, 7, 38] can be used to mitigate label inefficiencies by exploring supervision in the data directly, thereby reducing the reliance on expensive expert labeling while efficiently scaling the data and models.
+FMs shine in a broad range of areas, including natural language processing, computer vision, speech recognition, and medical image analysis.
+FMs have also been explored in remote sensing image understanding, but it is mostly limited to the evaluation of optical data.
+To our knowledge, the huge potential of FMs for SAR image interpretation remains completely locked.
+In our preliminary work, a novel SSL method for SAR imagery named SAR Joint-Embedding Predictive Architecture (SAR-JEPA) [39] was proposed and demonstrated promising results.
+Given the aforementioned discussion, to fully unlock the potential of LMs for SAR image interpretation, we present the first attempt toward building a foundation model for SAR ATR and propose SARatrX that learns generalizable representations via SSL and provides a basis for label-efficient model adaptation to generic SAR target detection and classification tasks.
+While conceptually simple, existing methods for building FMs in other domains cannot be just directly applied to SAR ATR.
+We are still facing significant challenges, which we will discuss below and address with SARatrX.
+Pre-training datasets must include diverse target categories and imaging conditions to accommodate various downstream tasks.
+However, SAR ATR lacks a large-scale dataset, such as ImageNet [40], and the most common MSTAR dataset only includes fine-grained vehicle categories that are not suitable for larger-scale pre-training.
+As such, several SSL methods combine different datasets.
+Specifically, SARDet-100K [21] incorporates 9 SAR target detection sets.
+With the number of open-source datasets used for SAR ATR increasing, most of the open-source datasets were integrated into this study as part of the pre-training.
+A total of 14 classification and detection datasets with different target categories and imaging conditions were included in this study as a new pre-training dataset, SARDet-180K, to explore the potential of FMs, as seen in Table I.
+Model backbones aim to achieve better spatial representations in remote sensing images, especially for small target signatures in large imagery.
+Transformers and convolutional neural networks are among the most common architectures used for these tasks.
+Transformers offer better spatial resolution without downsampling, while convolutional neural networks (CNNs) provide higher efficiency through convolution kernels.
+The performance of different architectures was tested in MSFA [21] with SAR detection tasks, the results of which have shown the Swin transformer backbone to be the most effective for remote sensing images.
+As such, HiViT [41] was selected for use in this study, as it has the advantages of a Swin transformer and can drop patches in Masked Image Modeling (MIM).
+Self-supervised learning is complicated by SAR image quality, which is negatively affected by speckle noise in coherent imaging.
+The coherent imaging resulting visual features are also not as distinct or rich as in natural RGB images.
+Contrastive learning [19, 42] uses data augmentation and preprocessing to reduce noise, while MIM [43, 39, 21] applies various target feature for guided signals to suppress noise.
+As such, the primary task of SAR SSL is the enhance the quality of feature learning and guide signals.
+For example, PGIL [6] leveraged a sub-frequency feature of complex SAR images to learn physics information, while our SAR-JEPA [39] applied multi-scale gradient ratios to solve for multiplicative speckle noise and capture target shapes.
+Furthermore, multi-stage training [21] from ImageNet to SAR diminished the effects of noise on model diversity, as seen in Fig.
+4.
+Thus, we applied two-step pretraining from ImageNet to SAR to increase model diversity during pretraining with SAR images.
+Besides, multiscale gradient features were used as high-quality guide signals for MIM with SAR images.
+Evaluation tasks need to comprehensively evaluate the performance of a foundation model for different tasks and settings.
+The availability of 3 open-source target datasets was utilized by first constructing a fine-grained classification dataset, SAR-VSA, with 25 categories to evaluate the effectiveness of the proposed improvements.
+A comprehensive comparison was then performed between the proposed SARatrX and existing methods for public classification and detection tasks.
+SARatrX achieved superior performance in 5 datasets across 8 task settings, as shown in Fig.
+1, which is competitive with prior methods on various SAR ATR tasks (few-shot vehicle classification, robust classification, detection with specific categories or various categories).
+We hope that this work could advance the development of the intersection of general SAR target recognition and FMs.
+The primary contributions of this study can be summarized as follows:
   1. $\bullet$
-    We propose a vision of general SAR target recognition, aiming to find universal solutions for various SAR ATR tasks and application in an era of SAR images with big unlabeled data and small labeled samples.
+    We present the first foundation model called SARatrX, which learns generalizable representations via SSL from large-scale unlabelled data and provides a basis for generic SAR target detection and classification tasks.
   2. $\bullet$
-    We have systematically investigated a foundation model solution for general SAR ATR and present the first SAR ATR foundation model named SARATR-X, which achieves new performance breakthroughs on a wide range of SAR target recognition datasets and settings.
+    We systematically investigate a foundation model framework for SAR ATR.
+We build a currently the largest publicly available pre-training dataset SARdet-180k and fully discussed the model architecture and proposed SSL method with many comparisons.
   3. $\bullet$
-    We hope this promising work will stimulate research interest in SAR foundation models, thereby advancing SAR image interpretation with frontier AI technologies.
+    The ability of SARatrX is comprehensively evaluated with various SAR ATR tasks, such as few-shot classification, robust classification, ship detection, aircraft detection, and detection with various categories tasks.
 The remainder of this paper is organized as follows.
 Sec.
-2 introduces related work in remote sensing and SAR ATR.
+II introduces related work in remote sensing and SAR ATR.
 Sec.
-3 introduces the proposed foundation model SARATR-X.
+III discusses the proposed foundation model (SARatrX).
+Secs.
+IV and V conduct extensive experiments to demonstrate the superiority of the proposed method.
 Sec.
-4 and 5 conducts extensive experiments to demonstrate the superiority of the proposed method.
-Sec.
-6 concludes the paper and discusses future work.
-2 Related Work
-TABLE II: Related work about SSL and foundation models for SAR ATR.
-It can be found that the exploration of the foundation model in SAR ATR.
-Researchers have explored different dataset settings, CNN and Transfomer architectures, and various SSL algorithms.
-Inspired by their successful work, we formally present the first SAR ATR foundation model, SARATR-X, to achieve various recognition tasks.
-Method Year Dataset Backbone SSL Tasks Description
-RotANet [82] 2021 MSTAR CNN Generative Target classsification (MSTAR) Predicting rotational patterns of targets as a loss regularization term for classification task.
-UACL [88] 2021 MSTAR, FUSAR-Ship CNN Contrastive Target classsification (MSTAR) Contrastive defense method to enhance model robustness to different adversarial attack.
-PGIL [29] 2022 sea-ice, urban ResNet-18 Contrastive Scene classsification (sea-ice, urban) Contrastive learning between sub-frequency features of complex images and deep features of amplitude images.
-BIDFC [92] 2022 MSTAR, OpenSARShip ResNet-18 Contrastive Target classsification (MSTAR, OpenSARShip) Weakly contrastive learning for pre-training in fine-grained datasets with similar sample.
-TSCL [56] 2023 MSTAR ResNet-50 Contrastive Target classsification (MSTAR, OpenSARShip) Pre-processing of SAR images for noise removal to improve image quality for Contrastive learning.
-FG-MAE [79] 2023 SSL4EO-S12 ViT-S Generative Scene classification (EuroSAT, BigEarthNet-MM), semantic segmentation (DFC2020) Hand-craft feature HOG as the target signal for MIM in SAR images.
-SAR-JEPA [38] 2024 MSAR, SAR-Ship, SARSim, SAMPLE ViT-B Generative Target classification (MSTAR, FUSARShip, SAR-ACD) Local reconstruction and multi-scale gradient feature for MIM with small object in noise SAR images.
-MSFA [43] 2024 ImageNet, DOTA, SARDet-100k Swin-B Generative Object detection (SARDet-100K, SSDD, HRSID) Multi-stage pre-training strategy for MIM from RGB images to SAR images.
-SARATR-X 2024 ImageNet, 14 SAR target datasets HiViT-B Generative Target classification and Object detection (5 datasets and 8 settings) A first foundation model for target recognition in SAR Images.
-As shown in Table I and II, visual foundation models are booming in remote sensing.
-Many models are proposed for various modalities and tasks.
-Our studies focus on the foundation model for SAR ATR, i.e., SAR images-based target classification and object detection.
-In the following, we introduce the recent development of remote sensing foundation models.
-2.1 Foundation models in remote sensing
-Remote sensing foundation models [30] have received widespread attention in the last three years.
-Researchers have made many breakthroughs, achieving effective learning on various modalities and tasks.
-In terms of pre-training datasets, researchers use existing large-scale datasets or collect a large number of samples from different sources.
-As for the model backbone, researchers have improved attention mechanisms, positional encoding, and other aspects to enhance the perception of complex spatial information.
-MIM has been used to learn spatial-temporal contextual information, while contrast learning is applied to multi-modal learning.
-SatMAE [13] proposed a novel masking strategy and temporal and spectral positional encoding for multi-spectral and temporal images.
-With a new dataset fMoW Sentinel with 13 frequency bands, SatMAE achieved new performance on scene classification and semantic segmentation tasks.
-RVSA [77] improved the pretraining ViT backbone with a rotated varied-size window attention method for the arbitrary-oriented objects.
-This work shows the importance of learning targets’ complex spatial contextual relationships in remote sensing images.
-RingMo [65] used a patch incomplete mask strategy for dense and small objects.
-Based on their self-built 2 million images, RingMo has proven effective in many tasks.
-RingMo-Sense [90] offered a three-branch network and masking strategy to model the spatio-temporal interaction for temporal images.
-CMID [55] combined contrastive learning and masked image modeling to learn global semantic information and local spatial information.
-GFM [52] focused on the differences between natural and remote sensing images and employed a multi-objective continual pretraining approach to leverage both knowledge.
-DiffusionSat [32] was the first remote sensing generative that employs geographic information embedding in stable diffusion.
-Scale-MAE [58] reconstructed images at different frequencies with improved positional encodings of ViT.
-FG-MAE [79] employed different hand-designed features to replace the original pixels in the MIM and improve the feature quality.
-And SMLFR [17] used a low pass filter to eliminate high-frequency information from the image pixel.
-Multi-modal remote sensing foundation models have been developed, such as SkySense [26] and OFA-Net [87].
-SkySense [26] proposed a multi-granularity contrastive learning method to learn representations of different modalities.
-Besides, a GEO-context prototype was applied to embed geographical contextual information.
-OFA-Net [87] applied a shared Transformer backbone for multiple modalities.
-Besides this, there are vision-language models [41], such as EarthGPT [98], SkyEyeGPT [93] and LHRS-Bot [54], which incorporate large language models to different remote sensing images modalities.
-However, due to the difficulty of annotating SAR images, the collection of public datasets used by EarthGPT only has 10,554 SAR ship images, much less than 84,838 infrared images and 907,945 optical images.
-Therefore, there is relatively less research on SAR foundation models due to the scarcity and fragmentation of the high-resolution SAR target datasets in Fig 2.
-We want to explore the visual foundation model based on SAR images with target recognition to stimulate research enthusiasm in this direction.
-Figure 2: Timeline of measured and simulated datasets for SAR target recognition (SAR magnitude images were processed in pseudo-color for better visualization).
-SAR target datasets have increased rapidly since the 2020s, greatly enriching the diversity of targets, scenes, and sensors.
-Target categories are mainly fine-grained vehicles, ships, and aircraft.
-Mixed means that the target category includes more than one category, such as a mixture containing various vehicles, ships, and aircraft in IRIS-SAR.
-Scenes contain cities, harbors, airports, etc.
-Sensors are devices with different resolutions and bands under satellite or airborne platforms.
-However, the number of samples (No.
-samples) in most datasets is only a few thousand with several categories due to the high costs and difficult annotation.
-It inspired us to propose SARATR-X with an integrated dataset in Table III.
-TABLE III: Description of our pre-training dataset for SARATR-X, which contains 14 open-source SAR datasets for pre-training.
+VI concludes the paper and discusses future work.
+II Related Work
+Visual foundation models are actively being used for remote sensing applications and several different algorithms have recently been proposed for various modalities and tasks.
+However, this study focuses on a foundation model used for SAR ATR (i.e., SAR image-based target classification and object detection).
+In the following sections, we introduce recent developments in remote sensing foundation models.
+II-A Foundation models for remote sensing
+Remote sensing foundation models [44] have received widespread attention in recent years and have achieved effective learning across various modalities and tasks.
+Many of these studies have used existing large-scale pre-training datasets or have collected large quantities of samples from different sources.
+Model backbones have been established by improving attention mechanisms, positional encoding, and other aspects used to enhance the perception of complex spatial information.
+MIM has also been used to learn spatial-temporal contextual information, while contrast learning has been applied to multi-modal learning.
+SatMAE [45] involves a novel masking strategy with the temporal and spectral positional encoding used for multi-spectral and temporal images.
+SatMAE has also achieved excellent performance in scene classification and semantic segmentation tasks using a new dataset (fMoW Sentinel) that includes 13 different frequency bands.
+RVSA [46] improves a pre-training ViT backbone using a rotated variable-size window attention method for arbitrarily oriented objects.
+This work demonstrates the importance of learning complex spatial contextual relationships for targets in remote sensing images.
+RingMo [47] utilizes a patch incomplete mask strategy for dense or small objects and a self-constructed set of 2 million images, proving effective in a variety of tasks.
+It shows the potential of large-scale pre-training.
+RingMo-Sense [48] offers a three-branch network and a masking strategy for modeling spatio-temporal interactions in temporal images.
+CMID [49] combines contrastive learning and masked image modeling to learn global semantic information and local spatial information.
+This work also shows the importance of learning the diversity of contextual relationships in remote-sensing images.
+GFM [50] focuses on the differences between natural and remote sensing images and employs a multi-object continual pre-training approach to leverage information from both.
+It shows that detailed information in natural images can complement remote sensing images well.
+DiffusionSat [51] is the first remote sensing generative model to employ geographic information embedding in stable diffusion.
+Scale-MAE [52] reconstructs images at different frequencies with improved positional encoding for ViT.
+FG-MAE [43] employs various hand-designed features to replace original pixels in the MIM, thereby improving the feature quality.
+Similarly, SMLFR [53] uses a low-pass filter to eliminate high-frequency information from the image pixels.
+These studies show the necessity of SSL high-quality guide signals.
+As such, our work focuses on the issues of SAR image quality while considering the range and spatial resolution of remote sensing.
+A variety of multi-modal remote sensing FMs have also been developed, including SkySense [54] and OFA-Net [55].
+SkySense [54] adopts a multi-granularity contrastive learning method to learn representations for different modalities.
+A GEO-context prototype has also been applied to embed geographical contextual information.
+OFA-Net [55] employs a shared transformer backbone for multiple modalities.
+In addition, vision-language models [56], such as EarthGPT [57], SkyEyeGPT [58], and LHRS-Bot [59] incorporate large language models into various remote sensing image modalities.
+However, due to the difficulty of annotating SAR images, the collection of public datasets used by EarthGPT only contains 10,554 SAR ship images, much less than the 84,838 infrared or 907,945 optical images.
+As a result, relatively little research is available on SAR FMs, due to the scarcity of high-resolution SAR target datasets (see Table I).
+This study explores visual FMs based on SAR images, using target recognition to promote research in this field.
+II-B Related SSL in SAR
+SSL for SAR ATR has been investigated in multiple studies.
+Early SSL was often used as a regularization loss for classification tasks, as discussed below.
+RotANet [60] predicts the rotational patterns of MSTAR vehicle targets by capturing azimuthal features for classification tasks.
+UACL [61] combines data augmentation and adversarial samples for contrastive learning, to improve model robustness to various adversarial attacks.
+PGIL [6] employs contrastive learning between complex SAR image sub-frequency features and deep amplitude image features, incorporating physical knowledge into classification tasks.
+SSL has also been used recently in model pre-training and fine-tuning frameworks.
+BIDFC [19] proposes weakly contrastive learning for pre-training in fine-grained vehicle datasets (MSTAR) and applied Gaussian noise data augmentation to simulate SAR image noise.
+TSCL [42] applies SAR image pre-processing prior to data augmentation in contrastive learning.
+FG-MAE [43] discusses different hand-crafted features for use with multi-spectral and SAR images and applies HOG features to SAR.
+In our previous studies, SAR-JEPA [39] SAR-JEPA [39] applies local reconstruction and multi-scale gradient features to collect target spatial signatures better.
+MSFA [21] proposes a multi-stage process with a filter augmentation pre-training framework for use in large-scale RGB and SAR data detection.
+However, these studies have only explored classification or detection tasks on a small number of datasets.
+Inspired by these previous works, this study aims to systematically investigate the construction of a SAR ATR foundation model for various ATR datasets via SSL.
+Our Insights - These studies have demonstrated that SSL can achieve performance improvements across multiple categories [39] and tasks [39, 21], which may be comparable to the performance of specially designed supervised methods [19, 21].
+However, it still lacks a foundation model for various SAR ATR applications.
+Besides, there is a lack of a pre-training and evaluation benchmark that contains images from classification and detection scenes.
+This inspired us to conduct systematic research into foundation models for general SAR target recognition, specifically with big data.
+We first extended the pre-dataset using different classification and detection tasks and scenarios (such as globally inland, marine, harbors, cities, and airports) based on existing research.
+A suitable model backbone is then discussed for the small target characteristics of remote sensing images.
+Since SSL requires high-quality guide signals from SAR images under the influence of noise, we applied two-step pre-training.
+Finally, we comprehensively evaluated the performance of the foundation models.
+III Approach
+We aim to construct a foundation model for general ATR from large-scale SAR images via an SSL method.
+As described above, increasing SAR datasets and SSL studies have inspired us to develop a foundation model for SAR ATR.
+This approach centers around pre-training datasets, model backbones, SSL methods, and evaluation tasks, to provide a systematic benchmark for SAR ATR foundation models.
+TABLE I: A description of the SARatrX data, which included 14 open-source SAR datasets used for pre-training.
+Large SAR imagery in the detection datasets contained more targets and scene types than the annotations.
+Some large images were cropped to increase the number of available training samples.
 # Img.: Number of images.
@@
 Pol.: Polarization.
-Large SAR imagery in detection datasets contains more target and scene types than the annotation.
-We cropped some large images to increase the training samples.
-Dataset Year # Img. Img. size # Target # Scene Res. (m) Band Pol. Description
+Dataset Year # Imgs. Img. Size # Targets # Scenes Res. (m) Band Pol. Description
 AIR-SARShip [62] 2019 801 $512\sim 1000$ $\geq 1$ $\geq 3$ $1\sim 3$ C Single Ship detection dataset in complex scenes
@@
 SSDD [23] 2021 1,160 $214\sim 668$ $\geq 1$ $\geq 2$ $1\sim 15$ C/X Quad Ship detection dataset
-2.2 Foundation models in SAR
-Researchers have explored SAR foundation models in different aspects, as shown in Table II. Inspired by previous studies, we systematically investigated how to construct a SAR ATR foundation model.
-Early SSL was often used as a regularization loss for classification tasks.
-RotANet [82] predicted the rotational pattern of MSTAR vehicle targets to capture azimuthal features for the classification task.
-UACL [88] combined data augmentation and adversarial samples in contrastive learning to improve the model’s robustness to various adversarial attacks.
-PGIL [29] used contrastive learning between the SAR complex images’ sub-frequency feature and the amplitude images’ deep feature to inject physical knowledge into the classification task.
-Recently, SSL has been used in model pre-training and fine-tuning frameworks.
-BIDFC [92] proposed weakly contrastive learning for pre-training in fine-grained vehicle datasets MSTAR and used Gaussian noise data augmentation to simulate SAR image noise.
-TSCL [56] applied the pre-processing of SAR images before data augmentation in contrastive learning.
-FG-MAE [79] discussed different hand-craft features for multi-spectral and SAR images and used the HOG feature for SAR.
-Our previous study, SAR-JEPA [38] and MSFA [43], focused on target classification and object detection.
-SAR-JEPA [38] applied Local reconstruction and multi-scale gradient feature to capture target spital signatures better.
-MSFA [43] proposed a multi-stage with filter augmentation pre-training framework to use large-scale RBG and SAR data for detection.
-Our insight.
-These studies have demonstrated that SSL can achieve performance improvements on various categories [38] and tasks [38, 43], and can even be comparable to the performance of specially designed supervised methods [92, 43].
-These inspired us to conduct systematic research for foundation models to achieve general SAR target recognition, especially in this era of big data.
-Firstly, we must extend the pre-dataset with different tasks and scenarios based on existing research.
-Secondly, a suitable model backbone needs to be discussed for the small target characteristics of remote-sensing images.
-Thirdly, SSL needs high-quality guide signals from SAR images under noise interference.
-Finally, we need to evaluate the performance of foundation models comprehensively.
-3 Approach
-We aim to construct a foundation model for general ATR from large-scale SAR images via the SSL method.
-As described above, the increasing SAR datasets and SSL studies inspired us to build a foundation model for SAR ATR.
-Our approach revolves around pre-training datasets, model backbones, SSL methods, and evaluation tasks to provide a systematic benchmark for foundation models in SAR ATR.
-Figure 3: Two steps of our pre-training.
-The first is to perform MIM on ImageNet to obtain better initialization weights, as shown in Fig 5 (c).
-Besides, we can easily obtain the open-source pre-training weights in step 1, thus reducing the pre-training time.
-The second is to perform MIM on SAR images with high-quality guide signals.
-We use multi-scale gradient features to suppress speckle noise and extract target edges.
-3.1 Creating a Diverse Pre-training Dataset
-Existing work mainly used MSTAR [1] as a pre-training dataset.
-While MSTAR is a high-quality vehicle target slice dataset, its commonly used samples are only a few thousand.
-Besides, this dataset suffers from background bias due to a single imaging scene [39].
-By comparison, the pre-training set ImageNet-1K in computer vision contains 1.4 million images with different categories and scenes.
-Using an insufficient SAR target pre-training dataset would underestimate the potential of this research area.
-Since diverse targets, scenes, and sensor conditions constitute a huge data sampling space in real-world situations, constructing a large pre-training dataset for the foundation model is central to unifying various SAR ATR tasks.
-The increasing SAR target datasets in Fig.
-2 is a primary motivation for achieving this goal.
-Although SAR images are expensive and no single dataset contains all popular target categories and various imaging conditions, collecting target samples from various open-source datasets can still construct a pre-training dataset with different categories, scenes, and sensors.
-Therefore, we build a new pre-training dataset containing 186,600 SAR target samples from 14 open-source SAR target datasets in Table III.
-We construct the dataset to include, as much as possible, target categories (terrestrial and maritime targets such as vehicles, ships, aircraft, oil tanks, bridges, etc.), scenes (typical scenes such as cities, harbors, airports, oceans, etc.), and sensors (satellite, airborne, and simulation platforms of different resolutions and bands).
-3.2 Choosing a Scalable Model
-We consider two model backbones for SAR target recognition.
-For the first, we apply Vision Transformer (ViT) [18], commonly used in SSL and has good scalability in model parameters.
-For the second architecture, we experiment with ConvNeXt-V2 [83], which has the same scalability as ViT and maintains the efficiency of convolutional neural networks.
-In addition to the scalability of the model parameters, the image properties need to be considered for remote sensing.
-The SAR target usually has a small foreground and a dynamic context range.
-Our previous study, MSFA [43], finds that the Swin Transformer is better than ViT with a hierarchical structure, but it is unsuitable for drop patches in MIM to save computing resources.
-Therefore, we also consider a variant of ViT, Hierarchical Vision Transformer (HiVit) [99], which improves the input spatial resolution and retains the property of ViT with MIM.
-3.3 Pre-Training with two steps
-We use MIM as the pretext task and masked autoencoders (MAE) [27] to save computational resources with dropping patches.
-MIM can help the foundation model achieve SAR image interpretation by recognizing the contextual relationship around objects.
-Here is a key point when applying MIM.
-SAR images belong to coherent imaging, and values have speckle noise that can interfere with the pretext task.
-Therefore, SARATR-X uses two pre-training steps to build a foundation model in Fig 3.
-The first step is to perform MIM on ImageNet to obtain better initialization weights.
-We simplify the multi-stage pre-training of MSFA, which performs SSL on ImageNet with backbone, detection task pre-training on DOTA with the whole framework, and detection task fine-tuning on SAR images.
-SARATR-X uses the pre-training weight of ImageNet as initialization weights for the SAR pre-training step.
-This way enhances the diversity of attention during the SAR pre-training step in Fig.
-5.
-In contrast, random initialization leads to a convergence of attention towards the same pattern in SAR pre-training with MAE.
-Besides, ImageNet pre-training backbone weights can be obtained from open source, thus greatly reducing pre-training time.
-We refer to using ImageNet pre-trained weights as SSL-ImageNet & SAR.
-The second step is to perform MIM with SAR images.
-As mentioned before, SAR image noise is a very tricky problem, and FG-MAR, SAR-JEPA, and MSFA have discussed many features, such as CannyEdge [8], HOG [14], CannyEdge [8], Haar-like [70], SAR-HOG [61] and SAR-SIFT [16].
-We can use different feature combinations to get the best results, but here is the simplest gradient feature that follows our previous SAR-JEPA to avoid getting bogged down on complex feature selection.
-We use the Multi-scale Gradient Feature (MGF) [38] to suppress the speckle noise and extract the target shape.
-TABLE IV: The results on the classification dataset with 25 categories.
-We perform the linear probing setting on the few-shot SAR target classification.
-For the pre-training datasets, we recommend using ImageNet’s pre-trained weights as initialization before pre-training on SAR datasets.
-For the model backbones, we prefer the HiVit to recognize the small target in SAR images.
-Model backbones are their base version.
-The classification metric is average accuracy.
-Bold indicates the best result, underline is the next best result.
-SL.
-Supervised Learning.
-SSL.
-Self-Supervised Learning.
+III-A Establishing a Diverse Dataset for Pretraining
+Previous research primarily employed MSTAR [18] as a pretraining dataset.
+While MSTAR provides high-quality vehicle targets, it only contains a few thousand commonly used samples.
+The images also suffer from background bias caused by a single imaging scene [73].
+In contrast, the ImageNet-1K pre-training set contains 1.4 million images with different categories and scenes.
+Since diverse target, scene, and sensor conditions constitute a large data sampling space in real-world scenarios, constructing a large pre-training dataset for the foundation model is central.
+The increasing availability of SAR target datasets as seen in Table.
+I, is a primary motivation for achieving this goal.
+Although SAR images are expensive and no single dataset contains all popular target categories or imaging conditions, collecting target samples from various open-source datasets can still provide a pre-training set with distinct categories, scenes, and sensors.
+As such, we constructed a new pre-training dataset, SARDet-180K, consisting of 186,600 SAR target samples from 14 open-source [^1] In this paper, we did not consider target datasets that are not open source or datasets for SAR terrain classification.
+Furthermore, we chose the typical open-source ship dataset as pre-training to avoid too many marine scenarios, and other open-source ship datasets about 28,749 images were not added, including FUSAR-Ship [74], LS-SSDD [75], DSSDD [76], HRSID [63], SRSDD [77], and RSDD [78].
+SAR target datasets’ all images, as described in Table I.
+This set aims, to the extent possible, to include common target categories (terrestrial and maritime targets such as vehicles, ships, aircraft, oil tanks, bridges, etc.), scenes (typical scenes such as cities, harbors, airports, oceans, etc.), and sensors (satellite, airborne, and simulation platforms of varying resolutions and bands).
+III-B Model Architecture
+Two model backbones were considered for SAR target recognition.
+The first utilized a vision transformer (ViT) [79], commonly used in SSL, offering good scalability of model parameters.
+The second architecture employed ConvNeXt-V2 [80], which offered the same scalability as ViT but maintained the efficiency of a convolutional neural network.
+In addition to the scalability of model parameters, remote sensing tasks must also consider image properties.
+For example, SAR targets typically exhibit a small foreground and a dynamic context range.
+Our previous study, MSFA [21], demonstrated that Swin transformers could outperform ViTs with a hierarchical structure, yet are unsuitable when using drop patches in MIM to preserve computing resources.
+Therefore, we finally considered a variant of ViT, a hierarchical vision transformer (HiViT) [41], which improved the input spatial resolution and retained ViT properties for MIM.
+Figure 2: The pre-training process included two steps.
+The first involved performing MIM on ImageNet data to obtain better initialization weights for model diversity, as shown in Fig.
+4 (c).
+The open-source pre-training weights used in step 1 could easily be obtained, thereby reducing the pre-training time.
+The second involved performing MIM on SAR images with high-quality guide signals.
+Multi-scale gradient features were used to suppress speckle noise and extract target edges.
+III-C Proposed Pretraining
+MIM was used as a pretext task, and masked autoencoders (MAE) [81] were employed to drop patches and preserve computational resources.
+MIM can help foundation models achieve SAR image interpretation by recognizing contextual relationships in objects (a key point when applying MIM).
+SAR utilizes a type of coherent imaging, which involves speckle noise that can interfere with a pretext task.
+As such, SARatrX employs two pretraining steps to construct a foundation model, as shown in Fig.
+2.
+The first step provides ImageNet weight to avoid the interference of SAR speckle noise in the early stages of the second pretraining step.
+Moreover, we use multiscale gradient features to suppress speckle noise throughout the second SAR pretraining step.
+The first step involves performing MIM with ImageNet data to obtain better initialization weights because visible light images contain more and bettter signtures compared to SAR images.
+We simplified the multistage pretraining of MSFA, which performs SSL on ImageNet with a backbone, detection task pretraining on DOTA with a whole framework, and detection task finetuning on SAR images.
+SARatrX uses the pre-training weights from ImageNet as initialization weights for the SAR pre-training step.
+This approach enhances the diversity of attention during SAR pre-training, as shown in Fig.
+4.
+In contrast, random initialization leads to the convergence of attention toward the same pattern as in SAR pre-training with MAE.
+The ImageNet pre-training backbone weights were obtained from an open source, thus reducing pre-training time significantly.
+This process of using pre-trained ImageNet weights is referred to as SSL-ImageNet & SAR.
+The second step involves performing MIM with SAR images.
+As mentioned previously, SAR image noise is a challenging problem that has been investigated in FG-MAR, SAR-JEPA, and MSFA, which have discussed features such as CannyEdge [82], HOG [83], Haar-like [84], SAR-HOG [85], and SAR-SIFT [86].
+Different feature combinations can be used to achieve the best results [21], but the simplest gradient features follow our previous SAR-JEPA approach to avoid excessive runtimes during complex feature selection.
+As such, multi-scale gradient features (MGFs) [39] were used to suppress speckle noise and extract target shapes.
+Multi-scale gradient feature - The classical differential gradient is not a constant false alarm rate operator due to multiplicative speckle noise in SAR images.
+This implies that speckle can cause gradient calculations to produce false points in strong target regions.
+Previous studies [87, 88] have demonstrated that a computing ratio is suitable for multiplicative noise.
+In this study, MGF employed a gradient by ratio [85, 86] to obtain the relevant gradient features $G_{\text{m}}$:
+$R_{i}$ $=\frac{M_{1}(i)}{M_{2}(i)},$ (1)
+$G_{\text{H}}$ $=log(R_{1}),$ (2)
+$G_{\text{V}}$ $=log(R_{3}),$ (3)
+where $R_{i}$ denotes the average ratio in different directions.
+$M_{1}(i)$ and $M_{2}(i)$ denote area averages on opposite sides of a current pixel along the direction $i$, where $i=1$ represents the horizontal direction and $i=3$ indicates the vertical direction.
+These area averages can be calculated from an input image and four fixed convolution kernels (see Fig.
+2).
+Eqs.
+2 and 3 then use logarithms to perform the vertical gradient calculation [86], where $G_{\text{H}}$ is the horizontal gradient and $G_{\text{V}}$ is the vertical gradient.
+$G_{\text{m}}$ $=\sqrt{G_{\text{H}}^{2}+G_{\text{V}}^{2}},$ (4)
+MGF $=\text{concat}(G_{\text{m1}},G_{\text{m2}},G_{\text{m3}}),$ (5)
+Due to the dynamic range required for various targets in remote sensing [89], MGF is constructed with convolutional kernels of different sizes.
+We set the kernel scale $r$ equal to 9, 13, and 17 to obtain $G_{\text{m1}}$, $G_{\text{m2}}$, and $G_{\text{m3}}$, and the whole convolutional kernel size is odd square $12r+$ 1 in order to calculate the average ratio in four different directions.
+III-D Evaluation with Recognition Tasks
+Fine-grained classification datasets comprised of vehicles, ships, and aircraft were merged to form a new SAR classification dataset called SAR-VSA (Vehicles, Ships, and Aircraft).
+This set was used to compare SSL model performance with the few-shot settings described in Sec.
+IV.
+We then report the SARatrX results for existing classification and detection settings, datasets, and algorithms in Sec.
+V.
+IV SARatrX Experiments
+We first performed SSL on the pre-training dataset without label information and then fine-tuned the pre-trained model on a classification dataset using few-shot classification tasks and linear probing settings to analyze improvements made to SARatrX.
+We also discuss the scalability of the proposed technique.
+Pre-training was first performed on eight NVIDIA RTX3090 GPUs.
+The SAR pre-training SARDet-180K dataset consisted of 14 SAR datasets (see Table I).
+Specifically, a few-shot SAR classification dataset, SAR-VSA, included 25 fine-grained targets from three SAR datasets (MSTAR [25], FUSAR-Ship [29], and SAR-ACD [29]).
+It is difficult to ensure training convergence by fine-tuning whole model parameters in small sample cases.
+As such, we used linear probing [81], which included a batch normalization layer, to adjust for differences in the statistical data properties and reduce the number of fine-tuning parameters.
+Detailed settings can be found in Appendix A.
+TABLE II: The results for a classification dataset containing 25 categories.
+Linear probing was performed during few-shot SAR target classification.
+We recommend using pre-trained ImageNet weights during initialization before pre-training the SAR datasets.
+We also recommend the HiViT model backbone for recognizing small targets in SAR images.
+Average accuracy was employed as the classification metric.
+Bolded text indicates the best result, while underlined text is the next best result.
+SL: supervised learning.
+SSL: self-supervised learning.
 Backbones Params Pre-Training Classification (N-Shot)
-setting dataset method 5 10 20
+Settings Dataset Method 5 10 20
 ConvNeXt-V2 89M SL-ImageNet ImageNet-1K Supervised 52.5 61.7 70.5
@@
 HiViT 66M SSL-ImageNet &\&& SAR ImageNet &\&& SAR Ours 76.5 80.8 85.1
-Multi-scale gradient feature.
-The classical differential gradient is not a constant false alarm rate operator due to multiplicative speckle noise in the SAR image.
-It means that speckle noise would cause the gradient computation to have false points in strong target regions.
-Previous studies [68, 6] have shown that the computing ratio is suitable for multiplicative noise.
-Here, MGF uses the gradient by ratio [61, 16] to obtain gradient features $G_{\text{m}}$.
-$R_{i}$ $=\frac{M_{1}(i)}{M_{2}(i)},$ (1)
-$G_{\text{H}}$ $=log(R_{1}),$ (2)
-$G_{\text{V}}$ $=log(R_{3}),$ (3)
-$G_{\text{m}}$ $=\sqrt{G_{\text{H}}^{2}+G_{\text{V}}^{2}},$ (4)
-where $R_{i}$ denotes the average ratio at different directions.
-$M_{1}(i)$ and $M_{2}(i)$ are the area averages on opposite sides of the current pixel along direction $i$.
-$i=1$ is the horizontal direction and $i=3$ means the vertical direction.
-The area averages can be computed from the input image and four fixed convolution kernels.
-Then, Eq.
-2 and 3 use logarithms to solve the vertical gradient calculation [16].
-$G_{\text{H}}$ is the horizontal gradient and $G_{\text{V}}$ is the vertical gradient.
-$\text{MGF}=\text{concat}(G_{\text{m1}},G_{\text{m2}},G_{\text{m3}})$ (5)
-Due to the dynamic range required for various targets in remote sensing [42], MGF is constructed with convolutional kernels of different sizes.
-We set the kernel scale $r$ equal to 9, 13, and 17 to obtain $G_{\text{m1}}$, $G_{\text{m2}}$, and $G_{\text{m3}}$, and the whole convolutional kernel size is odd square $12r+$ 1.
-3.4 Evaluating with Recognition Tasks
-We merged fine-grained classification datasets of Vehicles, Ships, and Aircraft as a new SAR classification dataset named (SAR-VSA).
-This dataset is used for comparing the SSL model’s performance with a few-shot setting in Sec.
-4.
-We then report SARATR-X results on existing classification and detection dataset settings with other methods in Sec.
-5.
-4 Experiments of SARATR-X
-First, we perform the SSL on the pre-training dataset without label information.
-Then, we fine-tune the pre-trained model on SAR-VSA with few-shot classification task and linear probing setting to analyze the improvement of SARATR-X.
-Finally, we discuss the scalability of the proposed approach.
-We perform pre-training on 8 NVIDIA RTX3090 GPUs.
-The SAR pre-training dataset consists of fourteen SAR datasets.
-The few-shot SAR classification dataset, SAR-VSA, contains 25 fine-grained targets from the three SAR datasets.
-It is difficult to ensure training convergence by fine-tuning the model’s whole parameters under small sample cases.
-Therefore, we use linear probing [27], which includes a batch normalization layer, to adjust the differences in the data statistical properties and reduce the number of fine-tuning parameters.
-Detailed settings are in the Appendix  A.
-4.1 Comparison of Model Backbones
-Table IV compares different model backbones for SAR ATR: ConvNeXt-V2 [83], ViT [18], and HiViT [99].
-ConvNeXt-V2 and ViT represent the two main architectures: CNN and Transformer.
-HiViT combines Vit and Swin Transformer [47].
-From the result, we can see that ViT outperforms ConvNeXt-V2.
-On the one hand, the ViT is more flexible than ConvNeXt-V2 in learning the contextual information in SAR images.
-On the other hand, multiple downsamplings in ConvNeXt-V2 result in the loss of small targets, while ViT maintains the same spatial resolution in different layers.
-HiVit, a visual Transformer with ViT’s flexibility and hierarchical representation, performs better than ViT and ConvNeXt-V2 in our self-supervised experiments.
-In particular, HiViT uses small (4 × 4) input patches, capturing small target features well.
-Fig.
-5 demonstrates that HiViT has a better various attention distance than ViT due to the small target information in remote sensing.
-4.2 Strategy of two-step pre-training
-Here, we discuss the two-step pre-training strategy to make full use of the available model weights and SAR datasets.
-Table IV has four pre-training settings: SL-ImageNet, SSL-ImageNet, SSL-SAR, and SSL-ImageNet & SAR.
-SL-ImageNet is pre-training on ImageNet with supervised learning [^1] We used open-source weights from GitHub to conduct the experiments.
-The supervised weights of ConvNeXt-V2 and HiViT were obtained by supervised fine-tuning after SSL.; SSL-ImageNet is pre-training on ImageNet with SSL from scratch; SSL-SAR is pre-training on our SAR pre-training dataset from scratch; SSL-ImageNet & SAR pre-training the model on SAR dataset based on the initialized weight from SSL-ImageNet.
-It is noticed that the additional supervised information introduced by SL-ImageNet does not necessarily improve the SAR ATR performance, e.g., the linear probing performance of SL-ImageNet for HiVit is lower than that of SSL-ImageNet.
-SSL-SAR can achieve better results than SSL-ImageNet using less data (12%), reflecting the huge differences in the target features between these two images.
-However, ImageNet pre-training weights can provide a good initialization for low features such as shape and texture in SSL with visible spectral remote sensing [67] and medical image [103].
-Our experiments also confirmed this conclusion: using SSL-ImageNet as initialization weights improves the pre-training performance of SAR images in Table IV and attention diversity in Fig.
-5.
-Therefore, our SARATR-X uses the SSL-ImageNet & SAR setting to complement the richness of the pre-training.
-4.3 Design of target signal for SAR images
-TABLE V: Comparisons of different target features for MIM.
-We found many target features are unsuitable for the multiplicative speckle noise in SAR images.
-This result inspired us to discuss different gradient features for SAR SSL.
-The pre-training setting SSL-SAR and the model backbones are the base version.
-The classification metric is average accuracy.
-Bold indicates the best result, underline is the next best result.
+TABLE III: Comparisons of various target features for MIM.
+Many target features were found to be unsuitable for the multiplicative speckle noise in SAR images.
+This result inspired us to pursue different gradient features for SAR SSL.
+The pre-training settings in SSL-SAR and the model backbones formed the base version, while average velocity served as the classification metric.
+Bolded text indicates the best result, while underlined text is the next best result.
 Model Target Feature Classification (N-shot)
@@
 HiViT MGF 76.0 81.1 84.5
-After discussing the backbone and strategy, we focus on the target features for SSL methods with SAR images.
-Due to the unique multiplicative speckle noise in SAR images, a key point for MIM is designing high-quality guide signals.
-As shown in Table V, we consider five target features (pixel value [27], low pass filter [46], HOG feature [79], deep feature [2], SAR-HOG [61], and gradient by ratio [16, 61]).
-All SSL methods use the SSL-SAR setting and base version.
-First, we consider whether the existing methods based on the ViT are suitable for SAR images.
-PixMIM [46] applies the low pass filter to remove the high-frequency components and drives the model to focus on shape information.
-However, PixMIM does not outperform MAE because the noise type of SAR is multiplicative, and the filter parameters require a trade-off between target and noise.
-Then, FG-MAE [79] uses HOG to capture the feature for the SSL with SAR scene-level tasks, but we find that HOG does not ensure high-quality SAR target features.
-Target regions usually have strong scattering values, and the speckle noise causes the gradient computation to have strong false points in these regions.
-Besides, I-JEPA [2] propose deep networks as target feature encoders to capture the deep semantic features but lead to training overfitting noise and failure to learn effective feature representations.
-Therefore, we choose the SAR features as target features to enhance HiViT.
-SAR-HOG changes the gradient calculation of HOG features and uses the gradient by ratio to solve speckle noise, and its result performs better than pixel value and HOG.
-Inspired by PixMIM, we prefer to use the target shape (i.e., gradient features) directly as the target feature  [^2] SAR-HOG uses the same multi-scale setting to illustrate that simple gradient features can effectively represent target shape as an SSL guide signal.
-Besides, multi-scale can improve the feature representation for various small targets in remote sensing.
-We discuss kernel settings for computing gradient in Fig.
-4.
-Scale can affect feature quality: a smaller scale is finer for small target edge extraction, while a larger scale is more suitable for large target and noise suppression.
-Therefore, combining features of different scales has an improvement over a single scale for various target sizes in images.
-Figure 4: Kernel settings for computing gradient by ratio.
-Here, the scale 1/2/3 is with $r$ equal to 9/13/17, and the multi-scale contacts all scales.
-Multi-scale is more suitable than single scale for various targets in remote sensing images.
-4.4 Analysis
-As stated above, SARATR-X’s key points are summarized as follows: HiViT architecture avoids the loss of small target information.
-SSL-ImageNet & SAR use ImageNet pre-training weights to provide a good initialization for diversity perceptual capability.
+IV-A Comparison of Model Backbones
+Table II compares different model backbones used for SAR ATR, including ConvNeXt-V2 [80], ViT [79], and HiViT [41].
+ConvNeXt-V2 and ViT represent the two primary architectures: a CNN and a transformer.
+Specifically, HiViT combines ViT with a Swin transformer [92].
+The results indicated that ViT outperformed ConvNeXt-V2.
+On the one hand, ViT is more flexible than ConvNeXt-V2 for learning contextual information in SAR images.
+On the other hand, multiple downsampling steps in ConvNeXt-V2 resulted in the loss of small targets, while ViT maintained the same spatial resolution in different layers.
+HiViT, a visual transformer with ViT’s flexibility and hierarchical representation, outperformed ViT and ConvNeXt-V2 in our self-supervised experiments.
+In particular, HiViT employed small (4 × 4) input patches, capturing small target features well.
+Fig.
+4 demonstrates that HiViT also offered a superior variable attention distance compared to ViT, due to the small target information common in remote sensing.
+Therefore, we use HiViT as backbone of our SARatrX.
+IV-B Strategy of two-step pre-training
+Here, we discuss the two-step pre-training strategy included to make full use of available model weights and SAR datasets.
+Table II includes four pre-training settings: SL-ImageNet, SSL-ImageNet, SSL-SAR, and SSL-ImageNet & SAR.
+SL-ImageNet was pre-trained on ImageNet using supervised learning [^2] We used open-source weights from GitHub to conduct the experiments.
+Supervised weights in ConvNeXt-V2 and HiViT were obtained with supervised fine-tuning after SSL.; SSL-ImageNet was pre-trained on ImageNet from scratch using SSL; SSL-SAR was pre-trained from scratch on our SAR pre-training dataset; SSL-ImageNet & SAR pre-trained the model on a SAR dataset based on initialized weights from SSL-ImageNet.
+Notice the additional supervised information introduced by SL-ImageNet did not necessarily improve SAR ATR performance (e.g., the linear probing performance of SL-ImageNet for HiVit was lower than that of SSL-ImageNet).
+SSL-SAR achieved better results than SSL-ImageNet using less data (12%), reflecting large differences in target features between the two images.
+However, ImageNet pre-training weights did provide a good initialization for lower features, such as shape and texture in SSL, with visible spectral remote sensing [93] and medical images [94].
+Our experiments also confirmed that using SSL-ImageNet as initialization weights improved the pre-training performance of SAR images (see Table II) and attention diversity (see Fig.
+4).
+As such, SARatrX employed the SSL-ImageNet & SAR settings to complement the richness of pre-training.
+IV-C Design of target signals for SAR images
+After investigating the backbone and strategy, we focused on target features for SSL methods used with SAR images.
+One key point for MIM is designing high-quality guide signals due to the unique multiplicative speckle noise in SAR images.
+This means we need to suppress noise and enhance target features.
+As seen in Table III, we considered five target features (pixel values) [81], including a low pass filter [90], HOG features [43], deep features [91], SAR-HOG [85], and gradient by ratio [86, 85]).
+All SSL methods used the SSL-SAR setting and base version.
+We first considered whether existing methods based on ViT were suitable for SAR image classification.
+PixMIM [90] applies a low pass filter to remove high-frequency components, driving the model to focus on shape information.
+However, PixMIM did not outperform MAE because the noise type in SAR is multiplicative, and the filter parameters require a trade-off between the target and noise.
+FG-MAE [43] uses HOG to capture SSL features in SAR scene-level tasks, though we found that HOG did not ensure accurate SAR target features.
+Target regions typically exhibited strong scattering values, and the included speckle noise often caused the gradient computations to exhibit strong false points in these regions.
+In addition, I-JEPA [91] proposed deep networks for use as target feature encoders to capture deep semantic features.
+However, this can lead to training overfitting noise and a failure to learn effective feature representations.
+As such, we chose SAR features as target features to enhance HiViT.
+SAR-HOG changes the gradient calculations for HOG features and uses gradient by ratio to solve for the speckle noise, thereby outperforming pixel values and HOG.
+Inspired by PixMIM, we prefer to directly use the target shape (i.e., gradient features) as a target feature  [^3] SAR-HOG uses the same multi-scale settings to illustrate that simple gradient features can effectively represent target shape as an SSL guide signal..
+In addition, multi-scale methods can improve the feature representations of various small targets common in remote sensing.
+Discussions of kernel settings used for computing gradients are illustrated in Fig.
+3.
+Scale can also affect feature quality, as a smaller scale is finer for small target edge extraction, while a larger scale is more suitable for large targets and noise suppression.
+Therefore, combining features of different scales (see Fig.
+3) offers improvements over a single scale for various image target sizes.
+IV-D Analysis
+As stated above, the primary contributions of SARatrX can be summarized as follows: the HiViT architecture avoids the loss of small target information.
+SSL-ImageNet & SAR use ImageNet pre-training weights to provide a good initialization for diversity perceptual capabilities.
 MGF ensures high-quality target features and suppresses speckle noise under SSL with SAR images.
-Benefiting the above insights, our SARATR-X can learn high-quality target features from noisy SAR remote sensing images in Table IV.
-Next, we analyze the diversity and scalability of SARATR-X.
-Figure 5: Averaged attention distance in different attention heads (x-axis is the attention head w.r.t layer number) on SSL models.
-Attention distance represents the range of the receptive field.
-We can focus on model architectures (Fig.
+By taking advantage of these insights, SARatrX can learn high-quality target features from noisy SAR remote sensing images, as seen in Table II.
+In the next section, we analyze the diversity and scalability of SARatrX.
+Figure 3: Discussions of single and multi-scale kernel settings for MGF.
+Here, the scale 1/2/3 assumes $r$ equal to 9/13/17, as the multi-scale contacts all scales.
+This multi-scale approach is more suitable than a single-scale technique for various targets in remote sensing images.
+Figure 4: Averaged attention distances for various attention heads (the x-axis is the attention head w.r.t layer number, and point colours represent different layers for better visualization) in the SSL models.
+Attention distance represents the range of a receptive field.
+We focused specifically on model architectures (Fig.
 (a) v.s.
@@
 Fig.
-(e)) to ensure diverse attention ranges in SAR target recognition, such as HiViT architecture, ImageNet weights, and SAR target features.
-Figure 6: Scaling ability of SARATR-X in dataset size, parameters, and training epochs with linear probing performance.
-Our method can benefit from these three aspects.
-However, it is important to note that excessive training epochs can lead to overfitting due to the dataset size and speckle noise.
-Visualization.
-Research [85] has shown that supervised pre-training and contrastive learning only model global information in the high layer, and MIM can model both local and global information.
-However, we observe that this phenomenon is not only related to the methods but also to the data properties.
+(e)) to ensure diverse attention ranges for SAR target recognition, including the HiViT architecture, ImageNet weights, and SAR target features.
+Figure 5: The scalability of SARatrX for dataset size, parameters, and training epochs with linear probing performance.
+While our method benefited from these three attributes, it is important to note that excessive training epochs often led to overfitting, due to the dataset size and speckle noise.
+Visualization - Prior research [95] has shown that supervised pre-training and contrastive learning only model global information in higher layers, while MIM can model both local and global information.
+However, we observed that this effect was not only related to the chosen method but also to the data properties.
 Fig.
-5 (a) shows that the ViT with MAE focuses on global information due to large SAR image scenes, which differs from MIM’s modeling properties.
-Therefore, HiViT has various attention ranges with its high spatial resolution and hierarchical structure in Fig.
+4 (a) demonstrates that ViT with MAE focuses on global information due to large SAR image scenes, which differs from MIM modeling properties.
+As such, HiViT includes various attention distances from 40 to 140 in different layers with its high spatial resolution and hierarchical structure, as seen in Fig.
 4 (b).
-In addition, using ImageNet’s weight as initialization can enhance this problem in Fig.
+In addition, using ImageNet weights for initialization also solved this problem, as seen in Fig.
 4 (c).
-As for the target feature, Fig.
-5 (d) displays the HOG feature enhances the noise interference, which harms feature diversity.
-MGF effectively extracts the shape information of the target, making the model more focused on diverse edge information in the lower layer.
-However, this approach removes texture and preserves edge information, which motivates the higher layers not to need to focus on texture details and diminishes the attention range in Fig.
+Similarly, Fig.
+4 (d) demonstrates that HOG features enhanced the noise interference, which limited feature diversity.
+MGF effectively extracts shape information from targets, focusing the model on diverse edge information in the lower layer.
+However, this approach removes textures and preserves edge information, which motivates higher layers to rely less on texture details and diminishes the attention range shown in Fig.
 4 (e).
-Therefore, we combine two-step training with MGF in Fig.
+Thus, we combined SSL-ImageNet &\&& SAR with MGF for two-step pre-training, as seen in Fig.
 4 (f).
-Scaling experiment.
-Although MIM is a good learner that scales with data and model resources [86], a question arises as to whether our method can ensure scalability for MIM when dealing with noisy data such as SAR.
+Scaling experiment - Although MIM learns effectively and scales with data and model resources [96], a question arises as to whether our method can ensure scalability for MIM when dealing with noisy data, such as SAR.
 Fig.
-6 presents the scaling experiment from three perspectives: dataset size, parameters, and training epochs.
-Despite our pre-training set comprising 186,660 images, which is smaller than ImageNet-1K, we observe a significant improvement in downstream tasks’ performance with increasing data and parameters.
-This phenomenon indicates that the foundational model can fully unleash its potential in SAR images by extracting high-quality features as guiding signals.
-However, just like the study [86], when the pre-training size contains 100,000 images, the model tends to overfit during extended training epochs.
-Moreover, SAR image noise and low resolution can further aggravate the overfitting.
-However, SARATR-X outperforms our previous study, SAR-JEPA, which has overfitted at 400 epochs with 94,776 SAR images.
-There is a need to continue investigating how to ensure high-quality feature representation when extending the SAR foundation models.
-5 Leveraging SARATR-X for Recognition
-We have discussed different aspects of SARATR-X, but there are many specific datasets and specialized models for SAR ATR.
-We thus compare our SARATR-X with other state-of-the-art methods, such as supervised learning ($\rm{CS}^{n}$et [9] and PD [95]), semi-supervised learning (EUAPS [100]), and self-supervised learning (MSFA [43] and BIDFC [92]).
-We focus on SAR recognition tasks, including image classification and detection.
-More detailed settings [^3] We removed downstream tasks’ test sets from the pre-training set samples.
-are in appendix B.
-TABLE VI: SAR classification results on MSTAR’s SOCs and EOCs settings.
-SOCs are in a similar distribution and have ten types of vehicle targets.
-EOCs include variations for imaging conditions that belong to robustness classification and out-of-distribution problems.
-EOCs use depression angle, configuration, and version variations to test robustness.
-We find that self-supervised (BIDFC and Ours) and semi-supervised (EUAPS) methods significantly perform under the few-shot task with additional unlabeled data, demonstrating the data’s importance.
-Besides, our method is robust to variations in imaging parameters.
-The metric is $\uparrow$, and the detail setting follows the $N$-shot [94].
-Bold indicates the best result, underline is the next best result.
-We provide detailed results in Appendix B.
+5 presents the results of a scaling experiment from three perspectives: dataset size, parameters, and training epochs.
+Despite our pre-training set comprising 186,660 images, which is smaller than ImageNet-1K, we observed significant rising curve in downstream task performance with increasing data and parameter quantities in Figs.
+5 (a) and (b).
+This result indicated that the foundational model could fully achieve its potential in SAR images by extracting high-quality features as guiding signals.
+However, as in [96], the model tended to overfit during extended training epochs when the pre-training set contained at least 100,000 images.
+In addition, SAR image noise and low resolution further aggravated the overfitting.
+Regardless, SARatrX outperformed our previous study (SAR-JEPA), which overfit at 400 epochs with 94,776 SAR images.
+Thus, there is a need to continue investigating new ways to ensure high-quality feature representations when extending SAR foundation models.
+V Leveraging SARatrX for Recognition
+We have discussed different aspects of SARatrX, but there are several datasets and specialized models available for SAR ATR.
+Therefore, we compared our proposed SARatrX algorithm with other state-of-the-art techniques, such as supervised learning ($\rm{CS}^{n}$et [97] and PD [98]), semi-supervised learning (EUAPS [99]), and self-supervised learning (MSFA [21] and BIDFC [19]).
+We focused on SAR recognition tasks, including image classification and detection.
+More detailed settings [^4] We removed the test sets of downstream tasks from the pre-training set samples.
+are provided in Appendix B.
+[…]
 SOCs-Standard operating conditions (10-way)
@@
 PD [98] 2024 46.7 58.9 70.2
-SARATR-X 2024 85.2 (+4.5) 91.4 (+6.1) 95.9 (+5.6)
+SARatrX 2024 85.2 (+4.5) 91.4 (+6.1) 95.9 (+5.6)
 EOCs-Depression angle variations (4-way)
@@
 CRID [102] 2023 62.1 62.3 74.5
-SARATR-X 2024 93.4 (+31.3) 97.3 (+33.2) 98.9 (+24.8)
+SARatrX 2024 93.4 (+31.3) 97.3 (+33.2) 98.9 (+24.8)
 EOCs-Target configuration variations (4-way)
@@
 CRID [102] 2023 62.8 65.7 74.1
-SARATR-X 2024 65.0 (+2.2) 74.0 (+8.3) 78.3 (+4.2)
+SARatrX 2024 65.0 (+2.2) 74.0 (+8.3) 78.3 (+4.2)
 EOCs-Target version variations (4-way)
@@
 CRID [102] 2023 53.5 56.2 67.2
-SARATR-X 2024 65.3 (+11.8) 76.5 (+20.3) 82.8 (+15.6)
-TABLE VII: SAR detection results on SARDet-100K, OGSOD, SSDD, and SAR-Aircraft.
-Our proposed SARATR-X has competitive performance on various detection datasets.
-Bold indicates the best result, underline is the next best result.
-More detailed results are in Appendix B.
+SARatrX 2024 65.3 (+11.8) 76.5 (+20.3) 82.8 (+15.6)
+[…]
 SARDet-100K (Object detection)
@@
 MSFA [21] 2024 56.4 88.2 61.5
-SARATR-X 2024 57.3 (+0.9) 88.7 (+0.5) 62.8 (+1.3)
+SARatrX 2024 57.3 (+0.9) 88.7 (+0.5) 62.8 (+1.3)
 OGSOD (Object detection)
@@
 YOLOv7 [109] 2022 45.1 79.2 -
-SARATR-X 2024 52.0 (+6.9) 85.9 (+6.7) 51.3
+SARatrX 2024 52.0 (+6.9) 85.9 (+6.7) 51.3
 SSDD (Ship detection)
@@
 $\rm{CS}^{n}$et [97] 2023 64.9 97.1 -
-SARATR-X 2024 67.5 (+2.6) 97.3 (+0.2) 83.5 (+7.3)
+SARatrX 2024 67.5 (+2.6) 97.3 (+0.2) 83.5 (+7.3)
 SAR-Aircraft (Aircraft detection)
@@
 SA-Net [24] 2023 - 77.7 62.8
-SARATR-X 2024 58.7 86.1 (+5.7) 64.7 (+3.3)
-Classification task.
-We study in Table VI the performance of SARATR-X on the MSTAR [1] dataset with Standard Operating Conditions (SOCs) and Extended Operating Conditions (EOCs).
-We first notice that SSL (BIDFC and ours) and semi-supervised (EUAPS) significantly outperform other methods under small samples with additional unlabeled samples.
-Our results surpass the previous best by large margins, demonstrating the value of foundation models in an era of rapidly growing SAR data.
-In particular, our SARATR-X shows robustness to the EOC setting of imaging condition variations.
-This result indicates that the foundation model learns stable features and relationships from diverse imaging conditions in a large number of target samples.
-Detection task.
-As illustrated in Table VII, we report box AP for SAR target detection with a horizontal bounding box on multi-category (SARDet-100K and OGSOD), ship (SSDD), and aircraft detection (SAR-Aircraft).
-SARATR-X performs better than our previous MSFA by 0.8 points on SARDet-100K.
-MSFA has more complex training processes and target features, which use multi-stage training between RGB and SAR images with three different target features and have a detection pre-training step.
-SARATR-X is more simple but more effective for SAR images.
-More significantly, SARATR-X outperforms or has comparable performance on various datasets compared to many specifically designed detection methods in Table VII.
-Of course, our study is only a preliminary exploration of SSL for SAR.
-More effective target features can be realized in a data-knowledge dual-driven manner by further mining the knowledge of SAR imaging mechanisms and properties.
-Moreover, with the larger dataset and parameters, the path of foundation models will hopefully lead to general SAR target recognition, but this goal requires researchers together.
-6 Conclusion and Future Perspectives
-In this paper, we proposed the SARATR-X for SAR ATR and systematically investigated a foundation model framework.
-First, a pre-training dataset was built from 14 open-source datasets, including various targets, scenes, and sensors.
-Then, the foundation model’s pre-training backbone, SSL methods, and downsteam tasks were discussed in detail.
-Importantly, SARATR-X demonstrated superior performance on different target recognition datasets, demonstrating the foundation model’s potential in this field.
-We believe that research on SAR foundation models, such as SARATR-X, has the potential for generalized feature representations in SAR images and accelerates towards all-day, all-weather target recognition in Earth observation.
-However, the research of the foundation model requires large data, which is a real problem for SAR images.
-SAR images are expensive and require specific imaging equipment and algorithms, and privacy and security prevent the data from opening.
+SARatrX 2024 58.7 86.1 (+5.7) 64.7 (+3.3)
+Classification task - Table IV details the performance of SARatrX for the MSTAR [18] dataset with standard operating conditions (SOCs) and extended operating conditions (EOCs).
+Notice that SSL (BIDFC and ours) and semi-supervised models (EUAPS) significantly outperformed other methods for small samples with additional unlabeled data.
+Our results also surpassed the previous best by large margins.
+For example, SOC 1-shot accuracy increased by 4.5% and EOCs 1-shot accuracy increased by 15.1% on average.
+demonstrating the value of FMs in an era of rapidly growing SAR data.
+In particular, SARatrX exhibited a robustness to the EOC setting of variable imaging conditions.
+This result indicated that the foundation model could learn stable features and relationships from diverse imaging conditions in a large number of samples.
+Detection task - As illustrated in Table V, we reported a box mAP for SAR target detection with a horizontal bounding box for multi-category (SARDet-100K and OGSOD), ship (SSDD), and aircraft detection (SAR-Aircraft).
+SARatrX outperformed our previous MSFA by 0.8 points on SARDet-100K.
+MSFA also includes complex training processes and target features, which employ multi-stage training between RGB and SAR images with three different target features (HOG [83], Haar-like [84], WSTG [115]), including a detection pre-training step.
+SARatrX is thus simpler yet more effective for SAR images.
+Notably, SARatrX outperformed or offered comparable performance for multiple datasets, compared to several specifically designed detection methods shown in Table V.
+Of course, our study is only a preliminary exploration of SSL for SAR image interpretation and more effective target features could be achieved in a data-knowledge dual-driven manner by further mining information on SAR imaging mechanisms and properties.
+Furthermore, given a larger dataset and computing power, the path of FMs will hopefully lead to generalized SAR interpretation include target recognition, scene classsification, semantic segmentation, and change detectionn, but this will require additional research.
+VI Conclusion and Future Perspectives
+This study proposed SARatrX for SAR ATR and systematically investigated a foundation model framework.
+First, a pre-training dataset SARDet-180K was constructed from 14 open-source datasets, including various targets, scenes, and sensors.
+The foundation model’s pre-training backbone, SSL methods, and downstream tasks were then discussed in detail.
+Importantly, SARatrX demonstrated superior performance on different target recognition datasets, demonstrating the potential of FMs in this field.
+We believe that further research on SAR foundation models, including SARatrX, has the potential to generalize feature representations of SAR images and accelerate all-day, all-weather target recognition in Earth observations.
+However, foundation model research requires large data, which is problematic for SAR images, as they are expensive and require specific imaging equipment and algorithms.
+Privacy and security may also prevent the data from becoming open source.
 Therefore, we are particularly grateful to the publishers of open-source SAR target datasets.
-By making SARATR-X publicly available, we aim to accelerate the progress of the foundation model in SAR target recognition by enabling researchers to use our dataset and code to design better methods or explore downstream applications.
-Although this work systematically investigated a foundation model framework, several limitations and challenges require exploration in future work.
-The SAR images are derived from open-source SAR datasets, and the targets are mainly vehicles, ships, aircraft, oil tanks, etc.
-Collecting target samples from increasingly unlabelled SAR imagery could further expand the amount of data and the range of downstream applications.
-In addition, investigating the expert knowledge with text for multimodal interactions to describe the relationship between targets and scenes can further enhance the representation capability of the foundation model.
-Due to the visual variability of SAR images, textual descriptions are an effective tool for improving interpretability and comprehensibility.
-In conclusion, we have verified the SARATR-X’s ability to adapt to diverse SAR target datasets, showing high performance and generalizability in classification and detection.
-By taking full advantage of the rapid growth of SAR images, the SSL-based foundation model opens the door to a generalized feature representation and SAR target recognition.
-We believe it’s time to embrace fundamental models for SAR image interpretation.
-Appendix A Implementation Details of Section 4 Experiments of SARATR-X
+By making SARatrX publicly available, we aim to accelerate the progress of FMs in SAR target recognition by enabling researchers to use our dataset and code to design better methods or explore downstream applications.
+Although this work systematically investigated a foundation model framework, several limitations and challenges will require exploration in a future study.
+The SAR images were derived from open-source SAR datasets, and the targets primarily included vehicles, ships, aircraft, oil tanks, etc.
+Thus, collecting target samples from increasingly unlabeled imagery could further expand the amount of data to reach a million level and the range of downstream applications.
+In terms of model architecture, learning the dynamic contextual information on space-time is also an important issue.
+In addition to target shape features, various scenarios and target signatures for SAR self-supervised learning need to be explored.
+Finally, investigating expert knowledge with text used for multimodal interactions, and describing the relationship between targets and scenes, could further enhance the representation capabilities of the foundation model for SAR ATR.
+In conclusion, we have demonstrated the ability of SARatrX to adapt to diverse SAR target datasets, achieving high performance and generalizability in classification and detection tasks.
+By taking full advantage of the rapid growth of SAR images, this SSL-based foundation model opens the door to a generalized SAR target recognition.
+Appendix A Implementation Details for Section 4 and SARatrX Experiments
 Here are the details of the dataset and training settings.
-A.1 Pre-traing dataset setting
-Figure 7: Visualization of the fourteen datasets included in our pre-training. In this way, a pre-trained dataset for SAR ATR can be built with multiple targets, scenes, and sensors.
+A-A Pre-traing dataset setting
+The timeline of measured and simulated datasets used for SAR target recognition is illusate in Fig.
+6, and we chose the typical open-source dataset as pre-training.
 As shown in Fig.
@@
 Here are brief descriptions of each dataset’s targets, scenes, and sensors.
+Figure 6: The timeline of measured and simulated datasets used for SAR target recognition (SAR magnitude images were processed in pseudo-color for better visualization).
+SAR target datasets have increased rapidly since 2020, greatly enriching the diversity of available targets, scenes, and sensors.
+The different colours represent the target categories included in the dataset.
+Target categories primarily include fine-grained vehicles, ships, and aircraft.
+A mixed sample implies the target exhibits more than one category, such as the combination of various vehicles, ships, and aircraft in IRIS-SAR.
+These scenes may also contain cities, harbors, airports, etc.
+The sensor devices offer different resolutions and bands for satellite or airborne platforms.
+However, the number of samples (No.
+samples) in most datasets is only a few thousand across several categories, due to high costs and the difficulty of annotation.
+This inspired us to propose SARatrX, with an integrated pre-training dataset SARDet-180K.
+Figure 7: Visualization of the fourteen datasets included in our pre-training. In this way, a pre-trained dataset for SAR ATR can be built with multiple targets, scenes, and sensors.
 AIR-SARShip [62] is a ship detection dataset based on the Chinese C-band Gaofen-3 satellite.
@@
 The dataset covers various ship sizes in different oceanic conditions with diverse clutter and noise interference.
-A.2 Classification dataset for performance Test
-TABLE VIII: Description of our SAR classification dataset, named SAR-VSA, which contains 25 fine-grained targets. # Train: Number of train samples. # Test: Number of test sample images.
+A-B Classification dataset for performance Test
+We select three target classification datasets, including 25 fine-grained targets from vehicles, ships, aircraft, and others, to evaluate the comprehensive performance of SSL and the foundation model for SAR target recognition.
+The new SAR classification dataset named SAR-Target in Table VI.
+MSTAR [18] is the most commonly used SAR vehicle datasetw.
+It has many experimental setting variants, while we refer to the  [25] to adopt the most commonly used ten-class classification settings, such as infantry vehicle, patrol car, personnel carrier, main battle tank, and truck.
+FUSAR-Ship [29] contains 15 primary ship categories and many non-ship targets based on the Gaofen-3 satellite in scenes such as sea, land, coast, river, and island.
+Based on the experimental setting of [116], we have ten ocean target types, such as four fine-grained ships, bridges, and ocean scene slices.
+SAR-ACD [29] contains five types of aircraft based on the Gaofen-3 satellite in three civil airports.
+Since the released dataset does not separate the training and test data, we randomly select partial samples as the training set and others as the test set.
+Fine-grained recognition of aircraft targets is a more challenging task due to the smooth surface of the aircraft, resulting in insignificant SAR image features.
+A-C Hyperparameter settings
+TABLE VI: Description of our SAR classification dataset, named SAR-VSA, which contains 25 fine-grained targets. # Train: Number of train samples. # Test: Number of test sample images.
 Fine-grained category # Train # Test
@@
 aircraft (Boeing 787) 113 391
-We select three target classification datasets, including 25 fine-grained targets from vehicles, ships, aircraft, and others, to evaluate the comprehensive performance of SSL and the foundation model for SAR target recognition.
-The new SAR classification dataset named SAR-Target in Table VIII.
-MSTAR [1] is the most commonly used SAR vehicle datasetw.
-It has many experimental setting variants, while we refer to the  [40] to adopt the most commonly used ten-class classification settings, such as infantry vehicle, patrol car, personnel carrier, main battle tank, and truck.
-FUSAR-Ship [64] contains 15 primary ship categories and many non-ship targets based on the Gaofen-3 satellite in scenes such as sea, land, coast, river, and island.
-Based on the experimental setting of [76], we have ten ocean target types, such as four fine-grained ships, bridges, and ocean scene slices.
-SAR-ACD [64] contains five types of aircraft based on the Gaofen-3 satellite in three civil airports.
-Since the released dataset does not separate the training and test data, we randomly select partial samples as the training set and others as the test set.
-Fine-grained recognition of aircraft targets is a more challenging task due to the smooth surface of the aircraft, resulting in insignificant SAR image features.
-A.3 Hyperparameter settings
-Here are the detailed settings of our pre-training and downstream tasks as shown in Table IX and X.
-Pre-training.
-Our default pre-training setting is in Table IX, and other hyperparameters of each method use the default settings from their papers and codes.
-Our pre-training is applied on 8 NVIDIA RTX3090 GPUs with 200 epochs and 800 batch sizes.
-Compared to the training settings of MAE, we add ColorJitter ($\rm{contrast}=0.5$) to increase data richness.
-Moreover, we modify the batch size and epoch according to 8 GPUs.
-It is worth noting that although MAE uses the normalized pixel value to enhance the feature representation in the visible spectral images, we find that the normalized pixel value cannot be used due to the SAR image noise and prevents the training loss from decreasing properly.
-TABLE IX: Pre-training setting.
+TABLE VII: Pre-training setting.
 Config Value
@@
 augmentation ResizedCrop, HFlip, ColorJitter
-Classification setting.
-All models use the same training settings in downstream classification tasks.
-Table X gives the default setting.
-Our few-shot learning setting is based on the Dassl toolbox [101, 102] and is averaged over 10 random experiments.
-Since we focus on the small sample case of the downstream classification task, we use the linear probing method in MAE to finetune models and avoid overfitting.
-TABLE X: Classification settings (fine-tuning and linear probing).
+TABLE VIII: Classification settings (fine-tuning and linear probing).
 Config Value
@@
 warmup learning rate 1e-5
-Figure 8: Partial fine-tuning results of HiViT-B and SSL-ImageNet & SAR.
-Tuning 0 blocks is linear probing.
-With the increasing number of fine-tuned transformer blocks, our method are consistently better than MAE and experience overfitting later.
-Partial fine-tuning.
-Fig.
-8 shows why we chose linear probing for few-shot evaluation, as HiViT overfits when many blocks are fine-tuned.
-Experimental results show that our SAR-KGPA consistently obtains better representations than MAE, and overfitting occurs later.
-Appendix B Implementation Details of Section 5 Leveraging SARATR-X for Recognition
-B.1 Dataset description
-We choose MSTAR, the most commonly used dataset in SAR target classification.
-SOCs are in similar image conditions, and the training set’s depression angle under SOC is 17°, while the test set is 15°.
-Ten categories of targets include BMP2, BRDM2, BTR60, BTR70, T62, T72, 2S1, D7, ZIL131, and ZSU234 as shown in Table VIII.
-The EOCs setting is the imaging condition variations to test robustness.
-Existing methods have been saturated with different experimental settings of MSTAR [40], but few samples and depression angle variations remain challenging.
-We use SARDet-100K and OGSOD datasets, which have many test samples and categories, to evaluate the detection performance fully.
-The OGSOD comparison results are derived from the original article’s [71] single-modal approach using only SAR images.
-SSDD and SAR-Aircraft are ship and aircraft categories.
-B.2 Hyperparameter settings
-Based on our scaling experiment, we use HiViT-B with 600 epochs pre-trained on SSL-ImageNet & SAR as the foundation model for classification and detection tasks.
-Classification setting is follow Table X. The only difference is that we use partial fine-tuning for better performance, and the last 6 blocks are added to the fine-tuned.
-Detection setting is follows the default setting in HiViT, and we adjust the learning rate to 5e-4.
-We used the same settings for each dataset fine-tuning, see our GitHub configuration based on the mmdetection [11] framework for details.
-B.3 Detailed results
-We provide detail classification and detection result on Table XI and XII.
-Although the proposed method outperforms existing methods in mAP, there is still scope for improvement in some refined metrics.
-TABLE XI: Detailed classification results on MSTAR’s SOCs and EOCs. Bold indicates the best result, underline is the next best result.
+TABLE IX: Detailed classification results on MSTAR’s SOCs and EOCs. “-” refers to unknown. Bold indicates the best result, underline is the next best result.
 SOCs: Standard operating conditions (10 way)
@@
 PD [98] 2024 46.7 58.9 62.0 67.5 70.2 83.7 - -
-SARATR-X 2024 85.2 91.4 93.9 95.1 95.9 97.7 98.1 98.5
+SARatrX 2024 85.2 91.4 93.9 95.1 95.9 97.7 98.1 98.5
 EOCs: Depression angle variations (4 way)
@@
 CRID [102] 2023 62.1 62.3 - - 74.5 85.9 - 87.0
-SARATR-X 2024 93.4 97.3 98.5 98.0 98.9 99.5 99.6 99.4
+SARatrX 2024 93.4 97.3 98.5 98.0 98.9 99.5 99.6 99.4
 EOCs: Target configuration variations (4-way)
@@
 CRID [102] 2023 62.8 65.7 - - 74.1 78.7 - 84.1
-SARATR-X 2024 65.0 74.0 74.0 84.2 78.3 87.0 88.9 93.5
+SARatrX 2024 65.0 74.0 74.0 84.2 78.3 87.0 88.9 93.5
 EOCs: Target version configuration variations (4-way)
@@
 CRID [102] 2023 53.5 56.2 - - 67.2 79.9 - 87.8
-SARATR-X 2024 65.3 76.5 76.3 83.4 82.8 85.0 92.1 97.0
-TABLE XII: Detailed detection results.
+SARatrX 2024 65.3 76.5 76.3 83.4 82.8 85.0 92.1 97.0
+TABLE X: Detailed detection results.
 The metrics are mAP, $\rm{mAP_{50}}$ (@@@@50), $\rm{mAP_{75}}$ (@@@@75), $\rm{mAP_{small}}$ (@@@@s), $\rm{mAP_{medium}}$ (@@@@m), and $\rm{mAP_{large}}$ (@@@@l).
+“-” refers to unknown.
 Bold indicates the best result, underline is the next best result.
@@
 MSFA [21] 2024 56.4 88.2 61.5 50.5 66.5 62.5
-SARATR-X 2024 57.2 (+0.8) 88.5 (+0.3) 62.6 (+1.1) 51.2 (+0.7) 69.6 (+3.1) 64.7 (+2.2)
+SARatrX 2024 57.2 (+0.8) 88.5 (+0.3) 62.6 (+1.1) 51.2 (+0.7) 69.6 (+3.1) 64.7 (+2.2)
 OGSOD* (Object detection)
@@
 YOLOv7 [109] 2022 45.1 79.2 - - - -
-SARATR-X 2024 51.4 (+6.3) 85.2 (+6.0) 51.7 46.8 75.3 77.1
+SARatrX 2024 51.4 (+6.3) 85.2 (+6.0) 51.7 46.8 75.3 77.1
 SSDD (Ship detection)
@@
 $\rm{CS}^{n}$et [97] 2023 64.9 97.1 - - - -
-SARATR-X 2024 67.4 (+2.5) 97.3 (+0.2) 83.1 (+6.9) 68.1 (+13.0) 66.1 (-2.1) 60.5 (-10.1)
+SARatrX 2024 67.4 (+2.5) 97.3 (+0.2) 83.1 (+6.9) 68.1 (+13.0) 66.1 (-2.1) 60.5 (-10.1)
 SAR-Aircraft (Aircraft detection)
@@
 SA-Net [24] 2023 - 77.7 62.8 - - -
-SARATR-X 2024 58.7 86.1 (+5.7) 64.7 (+3.3) 20.0 57.2 49.7
+SARatrX 2024 58.7 86.1 (+5.7) 64.7 (+3.3) 20.0 57.2 49.7
+Here are the detailed settings of our pre-training and downstream tasks as shown in Table VII and VIII.
+Pre-training - Our default pre-training setting is in Table VII, and other hyperparameters of each method use the default settings from their papers and codes.
+Our pre-training is applied on 8 NVIDIA RTX3090 GPUs with 200 epochs and 800 batch sizes.
+Compared to the training settings of MAE, we add ColorJitter ($\rm{contrast}=0.5$) to increase data richness.
+Moreover, we modify the batch size and epoch according to 8 GPUs.
+It is worth noting that although MAE uses the normalized pixel value to enhance the feature representation in the visible spectral images, we find that the normalized pixel value cannot be used due to the SAR image noise and prevents the training loss from decreasing properly.
+Figure 8: Partial fine-tuning results of HiViT-B and SSL-ImageNet & SAR.
+Tuning 0 blocks is linear probing.
+With the increasing number of fine-tuned transformer blocks, our method are consistently better than MAE and experience overfitting later.
+Classification setting - All models use the same training settings in downstream classification tasks.
+Table VIII gives the default setting.
+Our few-shot learning setting is based on the Dassl toolbox [121, 122] and is averaged over 10 random experiments.
+Since we focus on the small sample case of the downstream classification task, we use the linear probing method in MAE to finetune models and avoid overfitting.
+Partial fine-tuning - Fig.
+8 shows why we chose linear probing for few-shot evaluation, as HiViT overfits when many blocks are fine-tuned.
+Experimental results show that our SAR-KGPA consistently obtains better representations than MAE, and overfitting occurs later.
+Appendix B Implementation Details for Section 5 and the Leveraging of SARatrX for Recognition Tasks
+B-A Dataset description
+We choose MSTAR, the most commonly used dataset in SAR target classification.
+SOCs are in similar image conditions, and the training set’s depression angle under SOC is 17°, while the test set is 15°.
+Ten categories of targets include BMP2, BRDM2, BTR60, BTR70, T62, T72, 2S1, D7, ZIL131, and ZSU234 as shown in Table VI.
+The EOCs setting is the imaging condition variations to test robustness.
+Existing methods have been saturated with different experimental settings of MSTAR [25], but few samples and depression angle variations remain challenging.
+We use SARDet-100K and OGSOD datasets, which have many test samples and categories, to evaluate the detection performance fully.
+The OGSOD comparison results are derived from the original article’s [22] single-modal approach using only SAR images.
+SSDD and SAR-Aircraft are ship and aircraft categories.
+B-B Hyperparameter settings
+Based on our scaling experiment, we use HiViT-B with 600 epochs pre-trained on SSL-ImageNet & SAR as the foundation model for classification and detection tasks.
+Classification setting is follow Table VIII. The only difference is that we use partial fine-tuning for better performance, and the last 6 blocks are added to the fine-tuned.
+Detection setting is follows the default setting in HiViT, and we adjust the learning rate to 5e-4.
+We used the same settings for each dataset fine-tuning, see our GitHub configuration based on the mmdetection [123] framework for details.
+B-C Detailed results
+We provide detail classification and detection result on Table IX and X. Although the proposed method outperforms existing methods in mAP, there is still scope for improvement in some refined metrics.

### v3

@@
-SARatrX: Towards Building A Foundation Model for SAR Target Recognition
+SARATR-X: Towards Building A Foundation Model for SAR Target Recognition
 Abstract
-Despite the remarkable progress in synthetic aperture radar automatic target recognition (SAR ATR), recent efforts have concentrated on the detection or classification of a specific and coarse category, e.g., vehicles, ships, airplanes, or buildings.
-One of the fundamental limitations of the top-performing SAR ATR methods is that the learning paradigm is supervised, task-specific, limited-category, closed-world learning, which depends on massive amounts of accurately annotated samples that are expensively labeled by expert SAR analysts and has limited generalization capability and scalability.
-In this work, we make the first attempt towards building a foundation model for SAR ATR, termed SARatrX.
-SARatrX learns generalizable representations via self-supervised learning (SSL) and provides a basis for label-efficient model adaptation to generic SAR target detection and classification tasks.
-Specifically, SARatrX is trained on 0.18 M unlabelled SAR target samples, which are curated by combining contemporary benchmarks and constitute the largest publicly available dataset till now.
-Considering the characteristics of SAR images, a backbone tailored for SAR ATR is carefully designed, and a two-step SSL method endowed with multi-scale gradient features was applied to ensure the feature diversity and model scalability of SARatrX.
-The capabilities of SARatrX are evaluated on classification under few-shot and robustness settings and detection across various categories and scenes, and impressive performance is achieved, often competitive with or even superior to prior fully supervised, semi-supervised, or self-supervised algorithms.
-Our SARatrX and the curated dataset are released at https://github.com/waterdisappear/SARatrX to foster research into foundation models for SAR ATR and SAR image interpretation.
+Despite the remarkable progress in synthetic aperture radar automatic target recognition (SAR ATR), recent efforts have concentrated on detecting and classifying a specific category, e.g., vehicles, ships, airplanes, or buildings.
+One of the fundamental limitations of the top-performing SAR ATR methods is that the learning paradigm is supervised, task-specific, limited-category, closed-world learning, which depends on massive amounts of accurately annotated samples that are expensively labeled by expert SAR analysts and have limited generalization capability and scalability.
+In this work, we make the first attempt towards building a foundation model for SAR ATR, termed SARATR-X.
+SARATR-X learns generalizable representations via self-supervised learning (SSL) and provides a cornerstone for label-efficient model adaptation to generic SAR target detection and classification tasks.
+Specifically, SARATR-X is trained on 0.18 M unlabelled SAR target samples, which are curated by combining contemporary benchmarks and constitute the largest publicly available dataset till now.
+Considering the characteristics of SAR images, a backbone tailored for SAR ATR is carefully designed, and a two-step SSL method endowed with multi-scale gradient features was applied to ensure the feature diversity and model scalability of SARATR-X.
+The capabilities of SARATR-X are evaluated on classification under few-shot and robustness settings and detection across various categories and scenes, and impressive performance is achieved, often competitive with or even superior to prior fully supervised, semi-supervised, or self-supervised algorithms.
+Our SARATR-X and the curated dataset are released at https://github.com/waterdisappear/SARATR-X to foster research into foundation models for SAR image interpretation.
 Index Terms:
-Synthetic Aperture Radar (SAR), Object Recognition, Object Detection, Foundation Model, Self-supervised learning, Deep Learning, Masked Image Modeling
-https://github.com/waterdisappear/SARatrX
+Synthetic Aperture Radar, Target Recognition, Object Detection, Foundation Model, Self-supervised learning, Deep Learning, Masked Image Modeling
 I Introduction
-Synthetic aperture radar (SAR) [1, 2, 3, 4], developed from electromagnetic scattering in microwave frequency bands, plays a crucial role in active Earth observations, functioning effectively across diverse weather conditions and lighting environments.
-With the rapid development of SAR imaging techniques, high-resolution SAR images can be accessed more easily than before, enabling even more research opportunities for SAR image intelligent interpretation.
+Synthetic aperture radar (SAR) [1, 2, 3, 4], developed from electromagnetic scattering in microwave bands, plays a crucial role in active Earth observations, functioning effectively across diverse weather conditions and lighting environments.
+With the rapid development of SAR imaging techniques, high-resolution SAR images can be accessed more easily than before, enabling even more research opportunities for intelligent interpretation of SAR images.
 SAR automatic target recognition (ATR), aiming at automatically localizing and classifying objects of interest (e.g., vehicles, ships, airplanes, or buildings) in SAR images, is a longstanding, important yet challenging problem in SAR image intelligent interpretation [5, 6, 7, 8].
-SAR ATR plays an essential role in civil and national defense applications such as modern airport management, disaster management, urban planning and infrastructure monitoring, military reconnaissance, and maritime surveillance.
+SAR ATR [^1] In this paper, we focus on SAR ATR’s classification and detection tasks.
+The classification aims to recognize the class and type of a single object, and the detection identifies the class and location of a single or multiple objects.
+plays an essential role in civil and national defense applications such as modern airport management, disaster management, urban planning and infrastructure monitoring, military reconnaissance, and maritime surveillance.
 Therefore, it has become an active research area for several decades [9, 10, 11, 12, 13, 14].
-In the past decade, deep learning has brought tremendous success for SAR ATR [15, 5, 16, 17, 8].
-Despite the significant progress, there are at least the following fundamental challenges that will need to be addressed to further advance the field of SAR ATR.
-Figure 1: The motivation and performance of SARatrX.
-(a) SAR ATR includes various imaging conditions, targets, scenes, and sensors.
-However, due to high costs, the datasets are often collected in specific settings for certain tasks.
-For example, MSTAR [18] is a ten-type vehicle target classification dataset in the X-band and grass scenarios, while SAR-Aircraft is a seven-type aircraft target detection dataset collected from three airports and a C-band satellite.
-Several specialized algorithms have been proposed for application to these datasets.
-However, the differing target characteristics, scene background information, and sensor parameters have complicated the generalization of existing algorithms.
-As such, this paper aims to develop a SAR ATR foundation model, a generalized method for conducting various tasks.
-(b) SARatrX performed well across 5 datasets with 8 settings.
-It was superior to existing SSL methods (BIDFC [19]) for target classification in the fine-grained vehicle MSTAR dataset [18] with a few-shot setting.
-In addition, it performed well under extended operating conditions (EOCs) [20] (i.e., imaging conditions with variable depression angle (EOCs-Depression), target configuration (EOCs-Config), and version (EOCs-Version)).
-SARatrX also demonstrated competitive object detection performance with existing supervised methods applied to various categories (SARDet-100K [21] and OGSOD [22]), as well as specific categories for ships (SSDD [23]) and aircraft (SAR-AIRcraft [24]).
-Our study shows the potential of a foundation model for generalized SAR ATR.
+In the past decade, deep learning has brought tremendous success for SAR ATR [15, 16, 17].
+Despite the significant progress, the following fundamental challenges will need to be addressed to advance the field of SAR ATR.
 Firstly, task-specific property.
-One of the fundamental limitations of the current SAR ATR methods [25, 26, 27, 28, 29, 30] is that one model is trained and evaluated on one specific task.
+One of the fundamental limitations of the current ATR methods [18, 19, 20, 21, 22] is that one model is trained and evaluated on one specific task.
 The detection and classification of a specific coarse category (e.g., vehicles, ships, airplanes, or buildings in Fig.
@@
 Secondly, heavy reliance on supervised learning.
-Recent progress in SAR ATR, while substantial, has been limited to supervised learning, which heavily depends on massive amounts of accurately annotated target samples that are expensively labeled by expert SAR analysts and have limited generalization capability and scalability.
+Recent progress in SAR ATR [23, 24, 25], while substantial, has been limited to supervised learning, which heavily depends on massive amounts of accurately annotated target samples that are expensively labeled by expert SAR analysts and have limited generalization capability and scalability.
 However, the scarcity of expert SAR analysts cannot meet such an exhaustive requirement, leaving vast amounts of SAR images unlabelled and unexploited.
@@
 This raises significant challenges when one intends to transfer prior knowledge from the natural image domain.
-Different strong prior knowledge of SAR imagery, including speckle noise, discrete target appearances, and the lack of geometry, texture, and contour cues, needs special consideration when designing backbone architectures.
+Different strong prior knowledge of SAR imagery, including speckle noise, discrete target appearances, and the lack of geometry, texture, and contour cues, needs special consideration when designing backbone architectures and learning strategies.
 Most of the current mainstream backbones and methods designed on natural images are not suitable for the aforementioned information.
@@
 Foundation models [30], pretrained on extensive data in a task-agnostic manner (generally via self-supervised learning), can be flexibly adapted to a wide range of downstream tasks.
-Self-supervised learning (SSL) [36, 37, 7, 38] can be used to mitigate label inefficiencies by exploring supervision in the data directly, thereby reducing the reliance on expensive expert labeling while efficiently scaling the data and models.
+Self-supervised learning (SSL) [31, 32, 33, 34, 35] can be used to mitigate label inefficiencies by exploring supervision in the data directly, thereby reducing the reliance on expensive expert labeling while efficiently scaling the data and models.
 FMs shine in a broad range of areas, including natural language processing, computer vision, speech recognition, and medical image analysis.
-FMs have also been explored in remote sensing image understanding, but it is mostly limited to the evaluation of optical data.
+As summarized in Table I, FMs have also been explored in remote sensing image understanding, but they are mostly limited to the evaluation of optical data.
 To our knowledge, the huge potential of FMs for SAR image interpretation remains completely locked.
+Figure 1: Various specialized SAR ATR datasets and tasks.
+SAR ATR includes various imaging conditions (i.e.
+operating condition), such as targets, scenes, and sensors.
+However, the datasets are often collected in specific settings for certain tasks due to high costs.
+For example, MSTAR [36] is a ten-type vehicle target classification dataset in the X-band and grass scenarios, and SAR-Aircraft is a seven-type aircraft detection dataset collected from three airports and a C-band satellite.
+Specialized algorithms have been proposed for these datasets.
+However, the differing target characteristics, scene information, and sensor parameters have complicated the generalization of existing algorithms.
+As such, this paper aims to develop a SAR ATR foundation model, a generalized method for conducting various tasks.
 In our preliminary work, a novel SSL method for SAR imagery named SAR Joint-Embedding Predictive Architecture (SAR-JEPA) [37] was proposed and demonstrated promising results.
-Given the aforementioned discussion, to fully unlock the potential of LMs for SAR image interpretation, we present the first attempt toward building a foundation model for SAR ATR and propose SARatrX that learns generalizable representations via SSL and provides a basis for label-efficient model adaptation to generic SAR target detection and classification tasks.
+Given the aforementioned discussion, to fully unlock the potential of FMs for SAR image interpretation, we present the first attempt toward building a foundation model for SAR ATR and propose SARATR-X that learns generalizable representations via SSL and provides a basis for label-efficient model adaptation to generic SAR target detection and classification tasks.
 While conceptually simple, existing methods for building FMs in other domains cannot be just directly applied to SAR ATR.
-We are still facing significant challenges, which we will discuss below and address with SARatrX.
+We are still facing significant challenges, which are discussed and addressed below.
 Pre-training datasets must include diverse target categories and imaging conditions to accommodate various downstream tasks.
-However, SAR ATR lacks a large-scale dataset, such as ImageNet [40], and the most common MSTAR dataset only includes fine-grained vehicle categories that are not suitable for larger-scale pre-training.
-As such, several SSL methods combine different datasets.
-Specifically, SARDet-100K [21] incorporates 9 SAR target detection sets.
-With the number of open-source datasets used for SAR ATR increasing, most of the open-source datasets were integrated into this study as part of the pre-training.
-A total of 14 classification and detection datasets with different target categories and imaging conditions were included in this study as a new pre-training dataset, SARDet-180K, to explore the potential of FMs, as seen in Table I.
-Model backbones aim to achieve better spatial representations in remote sensing images, especially for small target signatures in large imagery.
-Transformers and convolutional neural networks are among the most common architectures used for these tasks.
-Transformers offer better spatial resolution without downsampling, while convolutional neural networks (CNNs) provide higher efficiency through convolution kernels.
-The performance of different architectures was tested in MSFA [21] with SAR detection tasks, the results of which have shown the Swin transformer backbone to be the most effective for remote sensing images.
-As such, HiViT [41] was selected for use in this study, as it has the advantages of a Swin transformer and can drop patches in Masked Image Modeling (MIM).
+However, SAR ATR lacks a large-scale dataset, such as ImageNet [38], and the most used MSTAR dataset only includes fine-grained vehicle categories that are not suitable for larger-scale pre-training (see Table II).
+As such, SARDet-100K [39] incorporates 9 SAR target detection datasets.
+With the number of open-source SAR ATR datasets increasing, most of the datasets were integrated into this study as part of the pre-training.
+A total of 14 classification and detection datasets with different target categories and imaging conditions were included as a new pre-training dataset, SARDet-180K, to explore the potential of FMs, as seen in Table III.
+Model backbones aim to achieve better spatial resolution representations in remote sensing images, especially for small targets in large imagery.
+Transformers and convolutional neural networks are among the most common architectures for these tasks.
+As shown in Table I, the Transformer offers better spatial resolution without downsampling, which is most commonly used in recent studies.
+As such, HiViT [40] was selected as it has the advantages of a Swin transformer high-resolution input and can drop patches in masked image modeling (MIM).
 Self-supervised learning is complicated by SAR image quality, which is negatively affected by speckle noise in coherent imaging.
-The coherent imaging resulting visual features are also not as distinct or rich as in natural RGB images.
-Contrastive learning [19, 42] uses data augmentation and preprocessing to reduce noise, while MIM [43, 39, 21] applies various target feature for guided signals to suppress noise.
-As such, the primary task of SAR SSL is the enhance the quality of feature learning and guide signals.
+The coherent imaging resulting visual features are also not as distinct or rich as in nature RGB images.
+Contrastive learning [41, 42] uses data augmentation and preprocessing to reduce noise, while MIM [43, 37, 39] applies various target feature for guided signals to suppress noise (see Table II).
+As such, the primary task of SAR SSL is to enhance the quality of feature learning and guide signals.
 For example, PGIL [6] leveraged a sub-frequency feature of complex SAR images to learn physics information, while our SAR-JEPA [37] applied multi-scale gradient ratios to solve for multiplicative speckle noise and capture target shapes.
 Furthermore, multi-stage training [39] from ImageNet to SAR diminished the effects of noise on model diversity, as seen in Fig.
-4.
-Thus, we applied two-step pretraining from ImageNet to SAR to increase model diversity during pretraining with SAR images.
-Besides, multiscale gradient features were used as high-quality guide signals for MIM with SAR images.
+5.
+Thus, we applied two-step pre-training from ImageNet to SAR to increase model diversity during pre-training with SAR images.
+Besides, multi-scale gradient features were used as high-quality guide signals for MIM with SAR images.
 Evaluation tasks need to comprehensively evaluate the performance of a foundation model for different tasks and settings.
-The availability of 3 open-source target datasets was utilized by first constructing a fine-grained classification dataset, SAR-VSA, with 25 categories to evaluate the effectiveness of the proposed improvements.
-A comprehensive comparison was then performed between the proposed SARatrX and existing methods for public classification and detection tasks.
-SARatrX achieved superior performance in 5 datasets across 8 task settings, as shown in Fig.
-1, which is competitive with prior methods on various SAR ATR tasks (few-shot vehicle classification, robust classification, detection with specific categories or various categories).
-We hope that this work could advance the development of the intersection of general SAR target recognition and FMs.
+Three open-source target datasets were utilized by first constructing a fine-grained classification dataset, SAR-VSA, with 25 categories to evaluate the effectiveness of the proposed improvements.
+A comprehensive comparison was then performed between the proposed SARATR-X and existing methods for public classification and detection tasks.
+SARATR-X achieved superior performance in 5 datasets across 8 task settings, as shown in Fig.
+2, which is competitive with prior methods on various SAR ATR tasks (classification with few-shot and robustness setting and detection with specific categories or various categories).
+We hope that this work could advance the development of the intersection of general SAR target recognition and foundation models.
 The primary contributions of this study can be summarized as follows:
+Figure 2: Results on classification and detection tasks.
+SARATR-X performed well across 5 datasets with 8 settings.
+It was superior to existing SSL methods (BIDFC [41]) for target classification in the fine-grained vehicle MSTAR dataset [36] with a few-shot setting.
+In addition, it performed well under extended operating conditions (EOCs) [44] (i.e., imaging conditions with variable depression angle (EOCs-Depression), target configuration (EOCs-Config), and version (EOCs-Version)).
+SARATR-X also demonstrated competitive object detection performance with existing supervised methods applied to various categories (SARDet-100K [39] and OGSOD [45]), as well as specific categories for ships (SSDD [46]) and aircraft (SAR-AIRcraft [47]).
+Our study shows the potential of a foundation model for SAR ATR.
   1. $\bullet$
-    We present the first foundation model called SARatrX, which learns generalizable representations via SSL from large-scale unlabelled data and provides a basis for generic SAR target detection and classification tasks.
+    We present the first foundation model called SARATR-X, which learns generalizable representations via SSL from large-scale unlabelled data and provides a cornerstone for generic SAR target detection and classification tasks.
   2. $\bullet$
     We systematically investigate a foundation model framework for SAR ATR.
-We build a currently the largest publicly available pre-training dataset SARdet-180k and fully discussed the model architecture and proposed SSL method with many comparisons.
+We build the largest publicly available pre-training dataset, SARdet-180k, and fully discuss the model architecture and proposed SSL method with many comparisons.
   3. $\bullet$
-    The ability of SARatrX is comprehensively evaluated with various SAR ATR tasks, such as few-shot classification, robust classification, ship detection, aircraft detection, and detection with various categories tasks.
+    SARATR-X is evaluated comprehensively with various SAR ATR tasks, such as few-shot classification, robust classification, ship detection, aircraft detection, and detection with various categories.
 The remainder of this paper is organized as follows.
@@
 Sec.
-III discusses the proposed foundation model (SARatrX).
+III discusses the proposed foundation model (SARATR-X).
 Secs.
@@
 II Related Work
-Visual foundation models are actively being used for remote sensing applications and several different algorithms have recently been proposed for various modalities and tasks.
-However, this study focuses on a foundation model used for SAR ATR (i.e., SAR image-based target classification and object detection).
-In the following sections, we introduce recent developments in remote sensing foundation models.
+TABLE I: A list of visual FMs used for remote sensing.
+There are many FMs are available for various modalities and tasks, but existing research (FG-MAE, SkySense, and OFA-Net) has focused on scene-level tasks with the SAR modality.
+SARATR-X represents an extension of our previous work (SAR-JEPA and MSFA [37, 39]) and advances the state of FMs for SAR ATR.
+SSL methods can be divided into two categories: contrastive (i.e., contrastive learning used to obtain invariant features) and generative (i.e., masked image modeling used to generate and predict features).
+SD: stable diffusion.
+Swin: Swin transformer.
+Foundation Model Year Modality Dataset Backbone SSL Tasks
+SatMAE [48] 2022 Multi-spectral, RGB fMoW RGB/Sentinel ViT Generative Scene classification and semantic segmentation
+RVSA [49] 2022 RGB MillionAID ViT, ViTAE Generative Scene classification, object detection, and semantic segmentation
+RingMo [50] 2022 RGB Self-built ViT, Swin Generative Scene classification, object detection, semantic segmentation, and change detectionn
+RingMo-Sense [51] 2023 RGB Self-built Video Swin Generative object detection and tracking, images segmentation, and 3 prediction tasks
+CMID [52] 2023 RGB MillionAID ResNet, Swin Generative & Contrastive Scene classification, object detection, and semantic segmentation
+GFM [53] 2023 RGB GeoPile Swin Generative Scene classification, semantic segmentation, change detection, and super-resolution
+DiffusionSat [54] 2023 RGB fMoW, Satlas, SpaceNet SD Generative Image generation, super-resolution, temporal generation, and in-painting
+Scale-MAE [55] 2023 RGB fMoW RBG ViT Generative Scene classification, object detection, and semantic segmentation
+FG-MAE [43] 2023 Multi-spectral, SAR SSL4EO-S12 ViT Generative Scene classification and semantic segmentation
+SMLFR [56] 2024 RGB GeoSense ConvNeXt Generative Object detection, semantic segmentation, and change detection
+SkySense [57] 2024 3 modalities HSROIs, TMsI, TSARI Swin, ViT Contrastive Scene classification, object detection, semantic segmentation, and change detection
+OFA-Net [58] 2024 4 modalities 5 multi-modalities datasets ViT Generative Scene Classification and semantic segmentation
+SARATR-X 2024 SAR ImageNet & 14 SAR datasets HiViT Generative Target classification and object detection
+TABLE II: Related SSL work in SAR ATR.
+They has involved the use of various datasets, architectures, and SSL algorithms.
+Inspired by previous work, this study presents the first SAR ATR foundation model, SARATR-X.
+Method Year Dataset Backbone SSL Tasks Description
+RotANet [59] 2021 MSTAR CNN Generative Target classsification (MSTAR) Predicting rotational patterns of targets as a loss regularization term for classification task.
+UACL [60] 2021 MSTAR, FUSAR-Ship CNN Contrastive Target classsification (MSTAR) Contrastive defense method to enhance model robustness to different adversarial attack.
+PGIL [6] 2022 Sea-ice, Urban ResNet-18 Contrastive Scene classsification (sea-ice, urban) Contrastive learning between sub-frequency features of complex images and deep features of amplitude images.
+BIDFC [41] 2022 MSTAR, OpenSARShip ResNet-18 Contrastive Target classsification (MSTAR, OpenSARShip) Weakly contrastive learning for pre-training in fine-grained datasets with similar sample.
+TSCL [42] 2023 MSTAR ResNet-50 Contrastive Target classsification (MSTAR, OpenSARShip) Pre-processing of SAR images for noise removal to improve image quality for Contrastive learning.
+FG-MAE [43] 2023 SSL4EO-S12 ViT-S Generative Scene classification (EuroSAT, BigEarthNet-MM), semantic segmentation (DFC2020) Hand-craft feature HOG as the target signal for MIM in SAR images.
+SAR-JEPA [37] 2024 MSAR, SAR-Ship, SARSim, SAMPLE ViT-B Generative Target classification (MSTAR, FUSARShip, SAR-ACD) Local reconstruction and multi-scale gradient feature for MIM with small object in noise SAR images.
+MSFA [39] 2024 ImageNet, DOTA, SARDet-100k ConvNeXt, Swin-B, Van Generative Object detection (SARDet-100K, SSDD, HRSID) Multi-stage pre-training strategy for MIM from RGB images to SAR images.
+SARATR-X 2024 ImageNet, 14 SAR target datasets HiViT-B Generative Target classification and object detection (5 datasets and 8 settings) A first foundation model for target recognition in SAR Images.
+Visual foundation models are actively being used for remote sensing applications, and different algorithms have recently been proposed for various modalities and tasks.
+This study focuses on a foundation model used for SAR ATR (i.e., SAR image-based target classification and object detection).
+In the following sections, we introduce recent developments in remote sensing and SAR foundation models.
 II-A Foundation models for remote sensing
-Remote sensing foundation models [44] have received widespread attention in recent years and have achieved effective learning across various modalities and tasks.
+Remote sensing foundation models [61, 62] have received widespread attention in recent years and have achieved effective learning across various modalities and tasks.
 Many of these studies have used existing large-scale pre-training datasets or have collected large quantities of samples from different sources.
@@
 However, due to the difficulty of annotating SAR images, the collection of public datasets used by EarthGPT only contains 10,554 SAR ship images, much less than the 84,838 infrared or 907,945 optical images.
-As a result, relatively little research is available on SAR FMs, due to the scarcity of high-resolution SAR target datasets (see Table I).
-This study explores visual FMs based on SAR images, using target recognition to promote research in this field.
+As a result, this study explores visual FMs based on unlabeled SAR images to improve the SAR ATR model’s scalability with large-scale samples.
 II-B Related SSL in SAR
-SSL for SAR ATR has been investigated in multiple studies.
-Early SSL was often used as a regularization loss for classification tasks, as discussed below.
+SSL for SAR ATR has been investigated in multiple studies, as detailed in Table II.
+Early SSL was often used as a auxiliary task for classification tasks, as discussed below.
 RotANet [59] predicts the rotational patterns of MSTAR vehicle targets by capturing azimuthal features for classification tasks.
-UACL [61] combines data augmentation and adversarial samples for contrastive learning, to improve model robustness to various adversarial attacks.
+UACL [60] combines data augmentation and adversarial samples for contrastive learning to improve model robustness to various adversarial attacks.
 PGIL [6] employs contrastive learning between complex SAR image sub-frequency features and deep amplitude image features, incorporating physical knowledge into classification tasks.
@@
 As described above, increasing SAR datasets and SSL studies have inspired us to develop a foundation model for SAR ATR.
-This approach centers around pre-training datasets, model backbones, SSL methods, and evaluation tasks, to provide a systematic benchmark for SAR ATR foundation models.
-TABLE I: A description of the SARatrX data, which included 14 open-source SAR datasets used for pre-training.
+We focus on pre-training datasets, model backbones, SSL methods, and evaluation tasks to provide a systematic benchmark for SAR ATR foundation models.
+First, we establish a diverse dataset for pre-training and integrate 14 SAR target datasets to ensure that most SAR target samples are included and that samples are balanced across categories.
+Besides, based on target size and image resolution, we slice bigger image sizes than 1000 pixels to increase the number of samples and extract small target slices.
+Second, we compare different model architectures to select a suitable model architecture for remote sensing recognition.
+Third, we propose a two-step pre-training strategy to achieve diversity in attention distance for different target sizes and effective training and scaling in the presence of SAR image noise interference.
+Natural image pre-training weights are used as SAR image pre-training initialization to enhance attentional diversity and multi-scale gradient features (MGFs) to suppress speckle noise.
+Finally, we perform a linear probing evaluation of proposed SSL strategies on the self-constructed SAR classification dataset and compare SARATR-X with existing methods on the public SAR target recognition dataset.
+III-A Pre-training Dataset
+Previous research primarily employed MSTAR [36] as a pre-training dataset.
+While MSTAR provides high-quality vehicle targets, it only contains a few thousand commonly used samples.
+The images also suffer from background bias caused by a single imaging scene [67].
+In contrast, the ImageNet-1K pre-training set contains 1.4 million images with different categories and scenes.
+Since diverse target, scene, and sensor conditions constitute a large data sampling space in real-world scenarios, constructing a large pre-training dataset for the foundation model is central.
+The increasing availability of SAR target datasets is a primary motivation for achieving this goal.
+Although SAR images are expensive and no single dataset contains all popular target categories or imaging conditions, collecting target samples from various open-source datasets can still provide a pre-training set with distinct categories, scenes, and sensors.
+Besides, considering that MSTAR’s background bias is due to the strong correlation between target classes and acquisition locations, we need to integrate different target datasets to increase the diversity of scenes and sensor conditions while learning target contextual information.
+As such, we constructed a new pre-training dataset, SARDet-180K, consisting of 186,600 SAR target samples from 14 open-source [^2] In this paper, we did not consider target datasets that are not open source or datasets for SAR terrain classification.
+Furthermore, we did not choose all open-source ship dataset as pre-training to avoid too many marine scenarios causing class imbalance [68], and other open-source ship datasets about 28,749 images were not added, including FUSAR-Ship [69], LS-SSDD [70], DSSDD [71], SRSDD [72], and RSDD [73].
+Besides, the class imbalance can be alleviated using resampling and hard-sample mining during the pre-training and fine-tuning in practice.
+SARDet-180K includes 46,096 vehicles, 73,973 ships, 35,861 aircraft, and 46,830 other samples.
+SAR target datasets’ all images, as described in Table III.
+This set aims, to the extent possible, to include common target categories (terrestrial and maritime targets such as vehicles, ships, aircraft, oil tanks, bridges, etc.), scenes (typical scenes such as cities, harbors, airports, oceans, etc.), and sensors (satellite, airborne, and simulation platforms of varying resolutions and bands).
+TABLE III: A description of the SARATR-X data, which included 14 open-source SAR datasets used for pre-training.
 Large SAR imagery in the detection datasets contained more targets and scene types than the annotations.
-Some large images were cropped to increase the number of available training samples.
+Some large images in Sandia MiniSAR and SAR-AIRcraft were cropped to increase the number of available training samples based on their target size and image resolution.
+More detailed descriptions are in Appendix A.
+Cls.: Classification.
+Det.: Detection.
 # Img.: Number of images.
@@
 Pol.: Polarization.
-Dataset Year # Imgs. Img. Size # Targets # Scenes Res. (m) Band Pol. Description
-AIR-SARShip [62] 2019 801 $512\sim 1000$ $\geq 1$ $\geq 3$ $1\sim 3$ C Single Ship detection dataset in complex scenes
-HRSID [63] 2020 5,604 800 $\geq 1$ $\geq 2$ $0.5\sim 3$ C/X Quad Ship detection and instance segmentation dataset
-Sandia MiniSAR [64] 2006 3,927 224 $\geq 1$ $\geq 7$ 0.1 Ku Single Terrestrial targets in urban areas, airports, deserts, and others
-MSAR [65, 66] 2022 28,499 $256\sim 2048$ $\geq 4$ $\geq 6$ 1 C Quad Terrestrial and maritime targets detection dataset
-MSTAR [18] 1995 14,577 $128\sim 193$ 10 1 0.3 X Single Fine-grained vehicle classification dataset
-OGSOD [22] 2023 18,331 256 $\geq 3$ $\geq 2$ 3 C Double Detection dataset for bridges, oil tanks, and harbours
-OpenSARShip [67] 2017 26,679 $9\sim 445$ 14 10 $2.3\sim 17.4$ C Double Fine-grained maritime target slices
-SADD [30] 2022 883 224 $\geq 1$ $\geq 2$ $0.5\sim 3$ X Single Aircraft detection dataset
-SAMPLE [68] 2019 5,380 128 10 2 0.3 X Single Simulation and measured vehicle dataset
-SAR-AIRcraft [24] 2023 18,818 512 $\geq 7$ $\geq 3$ 1 C Single Aircraft detection dataset
-SARSim [69, 70] 2017 21,168 139 14 3 0.3 X Single Simulation vehicle dataset
-SAR-Ship [71] 2019 39,729 256 $\geq 1$ $\geq 4$ $3\sim 25$ C Quad Ship detection dataset in complex scenes
-SIVED [72] 2023 1,044 512 $\geq 1$ $\geq 4$ $0.1\sim 0.3$ X/Ku/Ka Single Synthetic vehicle detection dataset
-SSDD [23] 2021 1,160 $214\sim 668$ $\geq 1$ $\geq 2$ $1\sim 15$ C/X Quad Ship detection dataset
-III-A Establishing a Diverse Dataset for Pretraining
-Previous research primarily employed MSTAR [18] as a pretraining dataset.
-While MSTAR provides high-quality vehicle targets, it only contains a few thousand commonly used samples.
-The images also suffer from background bias caused by a single imaging scene [73].
-In contrast, the ImageNet-1K pre-training set contains 1.4 million images with different categories and scenes.
-Since diverse target, scene, and sensor conditions constitute a large data sampling space in real-world scenarios, constructing a large pre-training dataset for the foundation model is central.
-The increasing availability of SAR target datasets as seen in Table.
-I, is a primary motivation for achieving this goal.
-Although SAR images are expensive and no single dataset contains all popular target categories or imaging conditions, collecting target samples from various open-source datasets can still provide a pre-training set with distinct categories, scenes, and sensors.
-As such, we constructed a new pre-training dataset, SARDet-180K, consisting of 186,600 SAR target samples from 14 open-source [^1] In this paper, we did not consider target datasets that are not open source or datasets for SAR terrain classification.
-Furthermore, we chose the typical open-source ship dataset as pre-training to avoid too many marine scenarios, and other open-source ship datasets about 28,749 images were not added, including FUSAR-Ship [74], LS-SSDD [75], DSSDD [76], HRSID [63], SRSDD [77], and RSDD [78].
-SAR target datasets’ all images, as described in Table I.
-This set aims, to the extent possible, to include common target categories (terrestrial and maritime targets such as vehicles, ships, aircraft, oil tanks, bridges, etc.), scenes (typical scenes such as cities, harbors, airports, oceans, etc.), and sensors (satellite, airborne, and simulation platforms of varying resolutions and bands).
+Dataset Year Task # Imgs. Img. Size # Targets # Scenes Res. (m) Band Pol. Target description
+MSTAR [36] 1995 Cls. 14,577 $128\sim 193$ 10 1 0.3 X Single Fine-grained vehicle dataset
+Sandia MiniSAR [74] 2006 Det. 3,927 224 $\geq 1$ $\geq 7$ 0.1 Ku Single Terrestrial targets in urban, deserts, and others
+SARSim [75, 76] 2017 Cls. 21,168 139 14 3 0.3 X Single Simulation vehicle dataset
+SAMPLE [77] 2019 Cls. 5,380 128 10 2 0.3 X Single Simulation and measured vehicle dataset
+SIVED [78] 2023 Det. 1,044 512 $\geq 1$ $\geq 4$ $0.1\sim 0.3$ X/Ku/Ka Single Synthetic vehicle dataset
+OpenSARShip [79] 2017 Cls. 26,679 $9\sim 445$ 14 10 $2.3\sim 17.4$ C Double Fine-grained ship slices
+SAR-Ship [80] 2019 Det. 39,729 256 $\geq 1$ $\geq 4$ $3\sim 25$ C Quad Ship dataset in complex scenes
+AIR-SARShip [81] 2019 Det. 801 $512\sim 1000$ $\geq 1$ $\geq 3$ $1\sim 3$ C Single Ship dataset in complex scenes
+HRSID [82] 2020 Det. 5,604 800 $\geq 1$ $\geq 2$ $0.5\sim 3$ C/X Quad Instance-level ship dataset
+SSDD [46] 2021 Det. 1,160 $214\sim 668$ $\geq 1$ $\geq 2$ $1\sim 15$ C/X Quad Ship dataset
+SADD [23] 2022 Det. 883 224 $\geq 1$ $\geq 2$ $0.5\sim 3$ X Single Aircraft dataset
+SAR-AIRcraft [47] 2023 Det. 18,818 512 $\geq 7$ $\geq 3$ 1 C Single Aircraft ataset
+MSAR [83, 84] 2022 Det. 28,499 $256\sim 2048$ $\geq 4$ $\geq 6$ 1 C Quad Terrestrial and maritime targets
+OGSOD [45] 2023 Det. 18,331 256 $\geq 3$ $\geq 2$ 3 C Double Targets include bridges, oil tanks, and harbours
 III-B Model Architecture
-Two model backbones were considered for SAR target recognition.
+Two model backbone types were considered for SAR target recognition.
 The first utilized a vision transformer (ViT) [85], commonly used in SSL, offering good scalability of model parameters.
 The second architecture employed ConvNeXt-V2 [86], which offered the same scalability as ViT but maintained the efficiency of a convolutional neural network.
-In addition to the scalability of model parameters, remote sensing tasks must also consider image properties.
+In addition to the scalability of model parameters, remote sensing tasks also need to consider image properties.
 For example, SAR targets typically exhibit a small foreground and a dynamic context range.
-Our previous study, MSFA [21], demonstrated that Swin transformers could outperform ViTs with a hierarchical structure, yet are unsuitable when using drop patches in MIM to preserve computing resources.
+Swin transformers could outperform ViTs with a hierarchical structure yet are unsuitable when using drop patches in MIM to preserve computing resources.
 Therefore, we finally considered a variant of ViT, a hierarchical vision transformer (HiViT) [40], which improved the input spatial resolution and retained ViT properties for MIM.
-Figure 2: The pre-training process included two steps.
+Figure 3: Two-step pre-training process.
 The first involved performing MIM on ImageNet data to obtain better initialization weights for model diversity, as shown in Fig.
 5 (c).
-The open-source pre-training weights used in step 1 could easily be obtained, thereby reducing the pre-training time.
-The second involved performing MIM on SAR images with high-quality guide signals.
-Multi-scale gradient features were used to suppress speckle noise and extract target edges.
-III-C Proposed Pretraining
-MIM was used as a pretext task, and masked autoencoders (MAE) [81] were employed to drop patches and preserve computational resources.
+The second involved performing MIM on SAR images with high-quality guide signals that are multi-scale gradient features suppressing speckle noise and extracting target edges.
+III-C Proposed Pre-training Method
+MIM was used as a pretext task for SSL pre-training, and masked autoencoders (MAE) [87] were employed to drop patches and preserve computational resources.
 MIM can help foundation models achieve SAR image interpretation by recognizing contextual relationships in objects (a key point when applying MIM).
-SAR utilizes a type of coherent imaging, which involves speckle noise that can interfere with a pretext task.
-As such, SARatrX employs two pretraining steps to construct a foundation model, as shown in Fig.
-2.
-The first step provides ImageNet weight to avoid the interference of SAR speckle noise in the early stages of the second pretraining step.
-Moreover, we use multiscale gradient features to suppress speckle noise throughout the second SAR pretraining step.
-The first step involves performing MIM with ImageNet data to obtain better initialization weights because visible light images contain more and bettter signtures compared to SAR images.
-We simplified the multistage pretraining of MSFA, which performs SSL on ImageNet with a backbone, detection task pretraining on DOTA with a whole framework, and detection task finetuning on SAR images.
-SARatrX uses the pre-training weights from ImageNet as initialization weights for the SAR pre-training step.
+However, SAR utilizes a type of coherent imaging, which involves speckle noise that can interfere with a pretext task.
+As such, SARATR-X employs two pre-training steps to construct a foundation model, as shown in Fig.
+3.
+The first step provides ImageNet weight to increase attention diversity and avoid the interference of SAR speckle noise in the early stages of the second step.
+We use multi-scale gradient features to suppress speckle noise throughout the second SAR pre-training step.
+The first step involves performing MIM with ImageNet data to obtain better initialization weights because visible light images contain more and better signatures than SAR images.
+We simplified the multistage pre-training of MSFA, which performs SSL on ImageNet with a backbone, detection task pre-training on DOTA with a whole framework, and detection task finetuning on SAR images.
+SARATR-X uses the pre-training weights from ImageNet as initialization weights for the SAR pre-training step.
 This approach enhances the diversity of attention during SAR pre-training, as shown in Fig.
-4.
+5.
 In contrast, random initialization leads to the convergence of attention toward the same pattern as in SAR pre-training with MAE.
-The ImageNet pre-training backbone weights were obtained from an open source, thus reducing pre-training time significantly.
-This process of using pre-trained ImageNet weights is referred to as SSL-ImageNet & SAR.
+Besides, the natural image weights also compensate well for the lack of diversity in the bottom layer of HiViT in SAR image pre-training and decreased diversity in the top layer due to MGFs.
+The ImageNet pre-training backbone weights were obtained from an open source to reduce pre-training time.
+This process of using pre-trained ImageNet weights is called SSL-ImageNet & SAR.
 The second step involves performing MIM with SAR images.
@@
 Different feature combinations can be used to achieve the best results [39], but the simplest gradient features follow our previous SAR-JEPA approach to avoid excessive runtimes during complex feature selection.
-As such, multi-scale gradient features (MGFs) [39] were used to suppress speckle noise and extract target shapes.
-Multi-scale gradient feature - The classical differential gradient is not a constant false alarm rate operator due to multiplicative speckle noise in SAR images.
-This implies that speckle can cause gradient calculations to produce false points in strong target regions.
-Previous studies [87, 88] have demonstrated that a computing ratio is suitable for multiplicative noise.
+As such, MGFs [37] were used to suppress speckle noise and extract target shapes.
+Multi-scale gradient feature - Multiplicative speckle noise causes distortions in the strong scattering of the region, which interferes with the SSL for the target features.
+Therefore, we would like to suppress the noise employing the target edge information for learning.
+However, Differential gradient have false points in strong target regions due to multiplicative speckle noise in SAR images [93, 94].
+This result is because multiplicative noise leads to scattering points with significant strong and weak amplitude variations in the strong scattering region, especially for strongly scattering metallic targets.
+Therefore, using a gradient ratio within a region can improve stability.
+It can reduce the interference in the strong and weak points inner target regions due to speckle noise.
 In this study, MGF employed a gradient by ratio [91, 92] to obtain the relevant gradient features $G_{\text{m}}$:
@@
 $M_{1}(i)$ and $M_{2}(i)$ denote area averages on opposite sides of a current pixel along the direction $i$, where $i=1$ represents the horizontal direction and $i=3$ indicates the vertical direction.
-These area averages can be calculated from an input image and four fixed convolution kernels (see Fig.
-2).
+The area averages can be calculated from an input image and four fixed convolution kernels (see Fig.
+3).
 Eqs.
@@
 Due to the dynamic range required for various targets in remote sensing [95], MGF is constructed with convolutional kernels of different sizes.
-We set the kernel scale $r$ equal to 9, 13, and 17 to obtain $G_{\text{m1}}$, $G_{\text{m2}}$, and $G_{\text{m3}}$, and the whole convolutional kernel size is odd square $12r+$ 1 in order to calculate the average ratio in four different directions.
+We set the kernel scale $r$ equal to 9, 13, and 17 to obtain $G_{\text{m1}}$, $G_{\text{m2}}$, and $G_{\text{m3}}$, and the whole convolutional kernel size is odd square $12r+$ 1 to calculate the average ratio in four different directions.
 III-D Evaluation with Recognition Tasks
-Fine-grained classification datasets comprised of vehicles, ships, and aircraft were merged to form a new SAR classification dataset called SAR-VSA (Vehicles, Ships, and Aircraft).
-This set was used to compare SSL model performance with the few-shot settings described in Sec.
+Fine-grained classification datasets comprised of vehicles, ships, and aircraft were merged to form a new SAR classification dataset called SAR-VSA (Vehicles, Ships, and Aircraft) in Table III.
+We aim to increase the number of classes and assess the proposed improvements in feature quality.
+SAR-VSA compares SSL model performance with few-shot settings in Sec.
 IV.
-We then report the SARatrX results for existing classification and detection settings, datasets, and algorithms in Sec.
-V.
-IV SARatrX Experiments
-We first performed SSL on the pre-training dataset without label information and then fine-tuned the pre-trained model on a classification dataset using few-shot classification tasks and linear probing settings to analyze improvements made to SARatrX.
+We then report the SARATR-X results for existing classification and detection settings, datasets, and algorithms in Sec.
+V to show the powerful potential of a foundation model.
+IV SARATR-X Experiments
+We first performed SSL on the pre-training dataset without label information and then fine-tuned the pre-trained model on a classification dataset using few-shot classification tasks and linear probing settings to analyze improvements made to SARATR-X.
 We also discuss the scalability of the proposed technique.
 Pre-training was first performed on eight NVIDIA RTX3090 GPUs.
-The SAR pre-training SARDet-180K dataset consisted of 14 SAR datasets (see Table I).
+The SAR pre-training SARDet-180K dataset consisted of 14 SAR datasets (see Table III).
 Specifically, a few-shot SAR classification dataset, SAR-VSA, included 25 fine-grained targets from three SAR datasets (MSTAR [18], FUSAR-Ship [22], and SAR-ACD [22]).
@@
 Detailed settings can be found in Appendix A.
-TABLE II: The results for a classification dataset containing 25 categories.
+TABLE IV: Results for a classification dataset containing 25 categories.
 Linear probing was performed during few-shot SAR target classification.
-We recommend using pre-trained ImageNet weights during initialization before pre-training the SAR datasets.
+According to the results, we recommend using pre-trained ImageNet weights during initialization before pre-training the SAR datasets.
 We also recommend the HiViT model backbone for recognizing small targets in SAR images.
@@
 HiViT 66M SSL-ImageNet &\&& SAR ImageNet &\&& SAR Ours 76.5 80.8 85.1
-TABLE III: Comparisons of various target features for MIM.
+TABLE V: Comparisons of various target features for MIM.
 Many target features were found to be unsuitable for the multiplicative speckle noise in SAR images.
-This result inspired us to pursue different gradient features for SAR SSL.
+These results inspired us to pursue different gradient features for SAR SSL.
 The pre-training settings in SSL-SAR and the model backbones formed the base version, while average velocity served as the classification metric.
@@
 IV-A Comparison of Model Backbones
-Table II compares different model backbones used for SAR ATR, including ConvNeXt-V2 [80], ViT [79], and HiViT [41].
-ConvNeXt-V2 and ViT represent the two primary architectures: a CNN and a transformer.
-Specifically, HiViT combines ViT with a Swin transformer [92].
+Table IV compares different model backbones used for SAR ATR, including ConvNeXt-V2 [86], ViT [85], and HiViT [40].
 The results indicated that ViT outperformed ConvNeXt-V2.
@@
 On the other hand, multiple downsampling steps in ConvNeXt-V2 resulted in the loss of small targets, while ViT maintained the same spatial resolution in different layers.
-HiViT, a visual transformer with ViT’s flexibility and hierarchical representation, outperformed ViT and ConvNeXt-V2 in our self-supervised experiments.
-In particular, HiViT employed small (4 × 4) input patches, capturing small target features well.
+HiViT outperformed ViT and employed small (4 × 4) input patches, capturing small target features well.
 Fig.
-4 demonstrates that HiViT also offered a superior variable attention distance compared to ViT, due to the small target information common in remote sensing.
-Therefore, we use HiViT as backbone of our SARatrX.
+5 demonstrates that HiViT also offered a superior variable attention distance compared to ViT due to the small target information common in remote sensing.
+Therefore, we use HiViT as the backbone of our SARATR-X.
 IV-B Strategy of two-step pre-training
 Here, we discuss the two-step pre-training strategy included to make full use of available model weights and SAR datasets.
-Table II includes four pre-training settings: SL-ImageNet, SSL-ImageNet, SSL-SAR, and SSL-ImageNet & SAR.
+Table IV includes four pre-training settings: SL-ImageNet, SSL-ImageNet, SSL-SAR, and SSL-ImageNet & SAR.
 SL-ImageNet was pre-trained on ImageNet using supervised learning [^3] We used open-source weights from GitHub to conduct the experiments.
@@
 However, ImageNet pre-training weights did provide a good initialization for lower features, such as shape and texture in SSL, with visible spectral remote sensing [98] and medical images [99].
-Our experiments also confirmed that using SSL-ImageNet as initialization weights improved the pre-training performance of SAR images (see Table II) and attention diversity (see Fig.
-4).
-As such, SARatrX employed the SSL-ImageNet & SAR settings to complement the richness of pre-training.
+Our experiments also confirmed that using SSL-ImageNet as initialization weights improved the pre-training performance of SAR images (see Table IV) and attention diversity (see Fig.
+5).
+As such, SARATR-X employed the SSL-ImageNet & SAR settings to complement the richness of pre-training.
 IV-C Design of target signals for SAR images
-After investigating the backbone and strategy, we focused on target features for SSL methods used with SAR images.
+After investigating the model backbone and learning strategy, we focused on target features for SSL methods used with SAR images.
 One key point for MIM is designing high-quality guide signals due to the unique multiplicative speckle noise in SAR images.
 This means we need to suppress noise and enhance target features.
-As seen in Table III, we considered five target features (pixel values) [81], including a low pass filter [90], HOG features [43], deep features [91], SAR-HOG [85], and gradient by ratio [86, 85]).
-All SSL methods used the SSL-SAR setting and base version.
+As seen in Table V, we considered five target features (pixel values) [87], including a low pass filter [96], HOG features [43], deep features [97], SAR-HOG [91], and gradient by ratio [92, 91]).
+All SSL methods use the SSL-SAR setting and base version.
 We first considered whether existing methods based on ViT were suitable for SAR image classification.
@@
 In addition, I-JEPA [97] proposed deep networks for use as target feature encoders to capture deep semantic features.
-However, this can lead to training overfitting noise and a failure to learn effective feature representations.
+However, this can lead to training overfitting noise and a failure to effective features.
+Figure 4: Discussions of single and multi-scale kernel settings for MGF.
+Here, the scale 1/2/3 assumes $r$ equal to 9/13/17, as the multi-scale contacts all scales.
+This multi-scale approach is more suitable than a single-scale technique for various targets in remote sensing images.
 As such, we chose SAR features as target features to enhance HiViT.
@@
 Discussions of kernel settings used for computing gradients are illustrated in Fig.
-3.
+4.
 Scale can also affect feature quality, as a smaller scale is finer for small target edge extraction, while a larger scale is more suitable for large targets and noise suppression.
 Therefore, combining features of different scales (see Fig.
-3) offers improvements over a single scale for various image target sizes.
+4) offers improvements for various image target sizes.
 IV-D Analysis
-As stated above, the primary contributions of SARatrX can be summarized as follows: the HiViT architecture avoids the loss of small target information.
+As stated above, the primary contributions of SARATR-X can be summarized as follows: the HiViT architecture avoids the loss of small target information.
 SSL-ImageNet & SAR use ImageNet pre-training weights to provide a good initialization for diversity perceptual capabilities.
 MGF ensures high-quality target features and suppresses speckle noise under SSL with SAR images.
-By taking advantage of these insights, SARatrX can learn high-quality target features from noisy SAR remote sensing images, as seen in Table II.
-In the next section, we analyze the diversity and scalability of SARatrX.
-Figure 3: Discussions of single and multi-scale kernel settings for MGF.
-Here, the scale 1/2/3 assumes $r$ equal to 9/13/17, as the multi-scale contacts all scales.
-This multi-scale approach is more suitable than a single-scale technique for various targets in remote sensing images.
-Figure 4: Averaged attention distances for various attention heads (the x-axis is the attention head w.r.t layer number, and point colours represent different layers for better visualization) in the SSL models.
+By taking advantage of these insights, SARATR-X can learn high-quality target features from noisy SAR remote sensing images, as seen in Table IV.
+In the next section, we analyze the diversity and scalability of SARATR-X.
+Figure 5: Averaged attention distances for various attention heads (the x-axis is the attention head w.r.t layer number, and point colors represent different layers for better visualization) in the SSL models.
 Attention distance represents the range of a receptive field.
@@
 (e)) to ensure diverse attention ranges for SAR target recognition, including the HiViT architecture, ImageNet weights, and SAR target features.
-Figure 5: The scalability of SARatrX for dataset size, parameters, and training epochs with linear probing performance.
-While our method benefited from these three attributes, it is important to note that excessive training epochs often led to overfitting, due to the dataset size and speckle noise.
+Figure 6: Scalability of SARATR-X for dataset size, parameters, and training epochs with linear probing performance.
+While our method benefited from these three attributes, it is important to note that excessive training epochs often led to overfitting due to the dataset size.
 Visualization - Prior research [100] has shown that supervised pre-training and contrastive learning only model global information in higher layers, while MIM can model both local and global information.
@@
 5 (a) demonstrates that ViT with MAE focuses on global information due to large SAR image scenes, which differs from MIM modeling properties.
-As such, HiViT includes various attention distances from 40 to 140 in different layers with its high spatial resolution and hierarchical structure, as seen in Fig.
+As such, HiViT includes various attention distances from 40 to 140 in different layers with its high spatial resolution and hierarchical structure but has decreased diversity in the bottom layer, as seen in Fig.
 5 (b).
-In addition, using ImageNet weights for initialization also solved this problem, as seen in Fig.
+In addition, using ImageNet weights for initialization solved this problem, as seen in Fig.
 5 (c).
@@
 5 (e).
-Thus, we combined SSL-ImageNet &\&& SAR with MGF for two-step pre-training, as seen in Fig.
+Thus, we combined SSL-ImageNet &\&& SAR with MGF for two-step pre-training in Fig.
 5 (f).
-Scaling experiment - Although MIM learns effectively and scales with data and model resources [96], a question arises as to whether our method can ensure scalability for MIM when dealing with noisy data, such as SAR.
-Fig.
-5 presents the results of a scaling experiment from three perspectives: dataset size, parameters, and training epochs.
-Despite our pre-training set comprising 186,660 images, which is smaller than ImageNet-1K, we observed significant rising curve in downstream task performance with increasing data and parameter quantities in Figs.
-5 (a) and (b).
-This result indicated that the foundational model could fully achieve its potential in SAR images by extracting high-quality features as guiding signals.
-However, as in [96], the model tended to overfit during extended training epochs when the pre-training set contained at least 100,000 images.
-In addition, SAR image noise and low resolution further aggravated the overfitting.
-Regardless, SARatrX outperformed our previous study (SAR-JEPA), which overfit at 400 epochs with 94,776 SAR images.
-Thus, there is a need to continue investigating new ways to ensure high-quality feature representations when extending SAR foundation models.
-V Leveraging SARatrX for Recognition
-We have discussed different aspects of SARatrX, but there are several datasets and specialized models available for SAR ATR.
-Therefore, we compared our proposed SARatrX algorithm with other state-of-the-art techniques, such as supervised learning ($\rm{CS}^{n}$et [97] and PD [98]), semi-supervised learning (EUAPS [99]), and self-supervised learning (MSFA [21] and BIDFC [19]).
-We focused on SAR recognition tasks, including image classification and detection.
-More detailed settings [^4] We removed the test sets of downstream tasks from the pre-training set samples.
-are provided in Appendix B.
+However, the benefits of different modalities, especially a sufficient number of visible images, for SAR data need to be explored more.
+In the future, we plan to collect the satellite and airborne multimodal paired data and use various features, including filter pixels, handmade features, and learned features, to ensure the migration of the underlying feature extraction and the higher-level semantic information.
+Continuous learning [53, 101] is also needed to prevent catastrophic forgetting.
 […]
@@
 PD [106] 2024 46.7 58.9 70.2
-SARatrX 2024 85.2 (+4.5) 91.4 (+6.1) 95.9 (+5.6)
+SARATR-X 2024 85.2 (+4.5) 91.4 (+6.1) 95.9 (+5.6)
 EOCs-Depression angle variations (4-way)
@@
 CRID [104] 2023 62.1 62.3 74.5
-SARatrX 2024 93.4 (+31.3) 97.3 (+33.2) 98.9 (+24.8)
+SARATR-X 2024 93.4 (+31.3) 97.3 (+33.2) 98.9 (+24.8)
 EOCs-Target configuration variations (4-way)
@@
 CRID [104] 2023 62.8 65.7 74.1
-SARatrX 2024 65.0 (+2.2) 74.0 (+8.3) 78.3 (+4.2)
+SARATR-X 2024 65.0 (+2.2) 74.0 (+8.3) 78.3 (+4.2)
 EOCs-Target version variations (4-way)
@@
 CRID [104] 2023 53.5 56.2 67.2
-SARatrX 2024 65.3 (+11.8) 76.5 (+20.3) 82.8 (+15.6)
+SARATR-X 2024 65.3 (+11.8) 76.5 (+20.3) 82.8 (+15.6)
 […]
@@
 MSFA [39] 2024 56.4 88.2 61.5
-SARatrX 2024 57.3 (+0.9) 88.7 (+0.5) 62.8 (+1.3)
+SARATR-X 2024 57.3 (+0.9) 88.7 (+0.5) 62.8 (+1.3)
 OGSOD (Object detection)
@@
 YOLOv7 [114] 2022 45.1 79.2 -
-SARatrX 2024 52.0 (+6.9) 85.9 (+6.7) 51.3
+SARATR-X 2024 52.0 (+6.9) 85.9 (+6.7) 51.3
 SSDD (Ship detection)
@@
 $\rm{CS}^{n}$et [118] 2023 64.9 97.1 -
-SARatrX 2024 67.5 (+2.6) 97.3 (+0.2) 83.5 (+7.3)
+SARATR-X 2024 67.5 (+2.6) 97.3 (+0.2) 83.5 (+7.3)
 SAR-Aircraft (Aircraft detection)
@@
 SA-Net [47] 2023 - 77.7 62.8
-SARatrX 2024 58.7 86.1 (+5.7) 64.7 (+3.3)
-Classification task - Table IV details the performance of SARatrX for the MSTAR [18] dataset with standard operating conditions (SOCs) and extended operating conditions (EOCs).
+SARATR-X 2024 58.7 86.1 (+5.7) 64.7 (+3.3)
+Scaling experiment - Although MIM learns effectively and scales with data and model resources [121], a question arises as to whether our method can ensure scalability for MIM when dealing with noisy data, such as SAR.
+Fig.
+6 presents the results of a scaling experiment from three perspectives: dataset size, parameters, and training epochs.
+Despite our pre-training set comprising 186,660 images, which is smaller than ImageNet-1K, we observed a significant rising curve in downstream task performance with increasing data and parameter quantities in Figs.
+6 (a) and (b).
+This result indicated that the foundational model could fully achieve its potential in SAR images by extracting high-quality features as guiding signals.
+However, as in [121], the model tended to overfit during extended training epochs when the pre-training set contained about 100,000 images in ImageNet.
+In addition, SAR image noise and low resolution further aggravated the overfitting.
+Regardless, SARATR-X outperformed our previous study (SAR-JEPA), which overfitted at 400 epochs with 94,776 SAR images.
+Thus, there is a need to continue investigating new ways to ensure high-quality feature representations when extending SAR foundation models.
+V Leveraging SARATR-X for Recognition
+Figure 7: Visualization of detection on SARDet-100K.
+False alarms and missed detections are common in SAR images, especially when similar targets overlap and complex scenes are present.
+While our method effectively improves the detection effect by learning the contextual information in the image, target detection in complex scenes and low-quality images is still very difficult.
+We have discussed different aspects of SARATR-X, but there are many datasets and specialized models available for SAR ATR.
+Therefore, we compared our proposed SARATR-X algorithm with other state-of-the-art techniques, such as supervised learning ($\rm{CS}^{n}$et [118] and PD [106]), semi-supervised learning (EUAPS [105]), and self-supervised learning (MSFA [39] and BIDFC [41]).
+We focused on SAR recognition tasks, including image classification and object detection in Fig.
+2.
+More detailed settings [^5] We removed the test sets of downstream tasks from the pre-training set samples, and other method results are from their original article tables.
+are provided in Appendix B.
+Classification task - Table VI details the performance of SARATR-X for the MSTAR [36] dataset with standard operating conditions (SOCs) and extended operating conditions (EOCs).
 Notice that SSL (BIDFC and ours) and semi-supervised models (EUAPS) significantly outperformed other methods for small samples with additional unlabeled data.
@@
 Demonstrating the value of FMs in an era of rapidly growing SAR data.
-In particular, SARatrX exhibited a robustness to the EOC setting of variable imaging conditions.
+In particular, SARATR-X exhibited robustness to the EOC setting of variable imaging conditions.
 This result indicated that the foundation model could learn stable features and relationships from diverse imaging conditions in a large number of samples.
-Detection task - As illustrated in Table V, we reported a box mAP for SAR target detection with a horizontal bounding box for multi-category (SARDet-100K and OGSOD), ship (SSDD), and aircraft detection (SAR-Aircraft).
-SARatrX outperformed our previous MSFA by 0.8 points on SARDet-100K.
+Detection task - As illustrated in Table VII, we reported a box mAP for SAR target detection with a horizontal bounding box for multi-category (SARDet-100K and OGSOD), ship (SSDD), and aircraft detection (SAR-Aircraft).
+SARATR-X outperformed our previous MSFA by 0.8 points on SARDet-100K.
 MSFA also includes complex training processes and target features, which employ multi-stage training between RGB and SAR images with three different target features (HOG [89], Haar-like [90], WSTG [122]), including a detection pre-training step.
-SARatrX is thus simpler yet more effective for SAR images.
-Notably, SARatrX outperformed or offered comparable performance for multiple datasets, compared to several specifically designed detection methods shown in Table V.
-Of course, our study is only a preliminary exploration of SSL for SAR image interpretation and more effective target features could be achieved in a data-knowledge dual-driven manner by further mining information on SAR imaging mechanisms and properties.
-Furthermore, given a larger dataset and computing power, the path of FMs will hopefully lead to generalized SAR interpretation include target recognition, scene classsification, semantic segmentation, and change detectionn, but this will require additional research.
+SARATR-X is thus simpler yet more effective for SAR images.
+The detection visualization is shown in Fig.
+7, and we have fewer missed detections and false alarm results.
+Notably, SARATR-X outperformed or offered comparable performance for multiple datasets, compared to several specifically designed detection methods shown in Table VII.
+Of course, our study is only a preliminary exploration of SSL for SAR image interpretation.
+More effective target features could be achieved in a data-knowledge dual-driven manner by further mining information on SAR imaging mechanisms and properties.
+Furthermore, given a larger dataset and computing power, the path of FMs will hopefully lead to generalized SAR interpretation, including target recognition, scene classification, semantic segmentation, and change detection, but this will require additional research.
 VI Conclusion and Future Perspectives
-This study proposed SARatrX for SAR ATR and systematically investigated a foundation model framework.
+This study proposed a foundation model SARATR-X for SAR ATR.
 First, a pre-training dataset SARDet-180K was constructed from 14 open-source datasets, including various targets, scenes, and sensors.
 The foundation model’s pre-training backbone, SSL methods, and downstream tasks were then discussed in detail.
-Importantly, SARatrX demonstrated superior performance on different target recognition datasets, demonstrating the potential of FMs in this field.
-We believe that further research on SAR foundation models, including SARatrX, has the potential to generalize feature representations of SAR images and accelerate all-day, all-weather target recognition in Earth observations.
-However, foundation model research requires large data, which is problematic for SAR images, as they are expensive and require specific imaging equipment and algorithms.
-Privacy and security may also prevent the data from becoming open source.
+Importantly, SARATR-X demonstrated superior performance on different target recognition datasets, demonstrating the potential of FMs in this field.
+We believe that further research on SAR foundation models, including SARATR-X, has the potential to generalize feature representations of SAR images and benefit all-day, all-weather target recognition in Earth observations.
+However, FMs research requires large data and SAR images are expensive and require specific imaging equipment and algorithms.
+Privacy and security also prevent the data from becoming open source.
 Therefore, we are particularly grateful to the publishers of open-source SAR target datasets.
-By making SARatrX publicly available, we aim to accelerate the progress of FMs in SAR target recognition by enabling researchers to use our dataset and code to design better methods or explore downstream applications.
-Although this work systematically investigated a foundation model framework, several limitations and challenges will require exploration in a future study.
+By making SARATR-X publicly available, we aim to accelerate the FMs in SAR target recognition by enabling researchers to use our code and weight to design better methods and explore downstream applications.
+Figure 8: Visualization of the fourteen datasets included in our pre-training.
+In this way, with the contribution of open-source researchers and the community, a pre-trained dataset for SAR ATR can be built with multiple targets, scenes, and sensors.
+Although this work performs systematical investigations, several limitations and challenges will require exploration in a future study.
 The SAR images were derived from open-source SAR datasets, and the targets primarily included vehicles, ships, aircraft, oil tanks, etc.
 Thus, collecting target samples from increasingly unlabeled imagery could further expand the amount of data to reach a million level and the range of downstream applications.
+The pre-training set image size [123] and automatic target slicing will become important issues that need to be investigated due to the larger image size and contextual range of remote sensing images compared to natural images.
 In terms of model architecture, learning the dynamic contextual information on space-time is also an important issue.
 In addition to target shape features, various scenarios and target signatures for SAR self-supervised learning need to be explored.
-Finally, investigating expert knowledge with text used for multimodal interactions, and describing the relationship between targets and scenes, could further enhance the representation capabilities of the foundation model for SAR ATR.
-In conclusion, we have demonstrated the ability of SARatrX to adapt to diverse SAR target datasets, achieving high performance and generalizability in classification and detection tasks.
-By taking full advantage of the rapid growth of SAR images, this SSL-based foundation model opens the door to a generalized SAR target recognition.
-Appendix A Implementation Details for Section 4 and SARatrX Experiments
+Finally, investigating expert knowledge with text used for multimodal interactions and describing the relationship between targets and scenes could enhance the representation capabilities of FMs for SAR ATR.
+In conclusion, we have demonstrated the ability of SARATR-X to adapt to diverse SAR target datasets, achieving high performance and generalizability in classification and detection tasks.
+By taking full advantage of the rapid growth of SAR images, this SSL-based foundation model opens the door to generalized SAR target recognition.
+Appendix A Implementation Details for Section 4 and SARATR-X Experiments
 Here are the details of the dataset and training settings.
 A-A Pre-traing dataset setting
-The timeline of measured and simulated datasets used for SAR target recognition is illusate in Fig.
-6, and we chose the typical open-source dataset as pre-training.
+we chose the most open-source dataset as pre-training.
 As shown in Fig.
-7, we collect data from open-source datasets based on our previous research [39, 21].
+8, we collect data from open-source datasets based on our previous research [37, 39].
 Now, our pre-training dataset contains 14 open-source SAR target datasets.
 Here are brief descriptions of each dataset’s targets, scenes, and sensors.
-Figure 6: The timeline of measured and simulated datasets used for SAR target recognition (SAR magnitude images were processed in pseudo-color for better visualization).
+Figure 9: Timeline of measured and simulated datasets used for SAR target recognition (SAR magnitude images were processed in pseudo-color for better visualization).
 SAR target datasets have increased rapidly since 2020, greatly enriching the diversity of available targets, scenes, and sensors.
@@
 samples) in most datasets is only a few thousand across several categories, due to high costs and the difficulty of annotation.
-This inspired us to propose SARatrX, with an integrated pre-training dataset SARDet-180K.
-Figure 7: Visualization of the fourteen datasets included in our pre-training. In this way, a pre-trained dataset for SAR ATR can be built with multiple targets, scenes, and sensors.
+This inspired us to propose SARATR-X, with an integrated pre-training dataset SARDet-180K.
+MSTAR [36] is the most commonly used target classification dataset released by the Defense Advanced Research Projects Agency, USA.
+Its sensor is an X-band radar with HH polarization mode and 0.3 m resolution.
+It contains ten categories of military vehicles with various imaging angles, target variants, and other conditions but with simple grass scenes and not completely open source.
+Sandia MiniSAR [74] is a 0.1 m resolution dataset based on a Ku-band airborne platform released by Sandia National Laboratories.
+The dataset contains scenes and targets such as aircraft on tarmacs, buildings in urban areas, and vehicles in desert areas but lacks official annotations.
+SARSim [75, 76] is a fine-grained vehicle dataset created by Terma A/S, Denmark.
+The simulation system used for this dataset can generate X-band SAR images with resolutions ranging from 0.1m to 0.3m from CAD models.
+SARSim provides 21,168 vehicle samples in 7 categories (truck, car, motorbike, bus, tank, bulldozer, and pickup) and 3 scenes (grass, roads, and a mean of the two) with 7 imaging depression angles.
+SIVED [78] is a vehicle detection dataset with rotatable bounding box.
+It consists of vehicle slices from the MSTAR dataset [36] and vehicles in urban areas from the Sandia MiniSAR and FARAD datasets [74], and scenes include car parks, buildings, trees, roads, and others.
+SAMPLE [77] is a synthetic and measured paired fine-grained vehicle dataset released by the Air Force Research Laboratory, USA.
+This dataset is simulated in X-band and 0.3 m resolution.
+The public version provides 5,380 images of ten categories of vehicle targets at partial imaging angles.
+OpenSARShip [79] is a ship slices dataset based on the European C-band Sentinel-1 satellite.
+Its resolution is 2.3 m to 17.4 m with VV and VH polarization.
+The dataset contains many ship slices from 10 busy ports.
+It has a diverse range of ship types but a significant category imbalance.
+SAR-Ship [80] is a ship target detection dataset in complex scenes based on Chinese Gaofen-3 and European Sentinel-1 satellites.
+The public version of this dataset contains 39,729 images from two satellites in different imaging modes and resolutions.
+The dataset provides ship targets of various sizes in complex ocean scenes such as nearshore, distant seas, harbors, and islands.
 AIR-SARShip [81] is a ship detection dataset based on the Chinese C-band Gaofen-3 satellite.
@@
 HRSID consists of 5,604 cropped SAR images with 0.5 to 3 m resolutions.
-The scene is a busy area of maritime transport, such as harbors and estuarine cities, and the annotation targets are ships of different sizes.
-Sandia MiniSAR [64] is a 0.1 m resolution dataset based on a Ku-band airborne platform released by Sandia National Laboratories.
-The dataset contains scenes and targets such as aircraft on tarmacs, buildings in urban areas, and vehicles in desert areas but lacks official annotations.
+The scene is a busy area of maritime transport, such as harbors and estuarine cities, and the annotation targets are civil ships of different sizes.
+SSDD [46] is a commonly used SAR ship detection dataset.
+It is constructed based on Canadian RadarSat-2, German TerraSAR-X, and European Sentinel-1 satellites and contains different scenarios for the inshore and offshore of China and India.
+The dataset covers various ship sizes in different oceanic conditions with diverse clutter and noise interference.
+SADD [23] is an aircraft detection dataset collected from the German X-band TerraSAR-X satellite.
+Its resolution is 0.5 m to 3 m with HH polarization.
+The dataset contains densely parked aircraft of different sizes on airport tarmacs, runways, and airport cemeteries.
+It has a large number of small-sized planes as well as the airport perimeter area.
+SAR-AIRcraft [47] is a aircraft detection dataset based on the Chinese C-band Gaofen-3 satellite with 1 m resolution and single polarization.
+The dataset collects seven types of aircraft of different sizes from three civil airports.
+It can support fine-grained aircraft detection and classification studies.
 MSAR [83, 84] is a multi-class target detection dataset based on the Chinese C-band HISEA-1 satellite in large-scale scenes.
@@
 The labeled target categories include aircraft, oil tanks, bridges, and ships.
-MSTAR [18] is the most commonly used target classification dataset released by the Defense Advanced Research Projects Agency, USA.
-Its sensor is an X-band radar with HH polarization mode and 0.3 m resolution.
-It contains ten categories of military vehicles with various imaging angles, target variants, and other conditions but with single grass scenes.
 OGSOD [45] is a city object detection dataset collected from the Chinese C-band Gaofen-3 satellite with VV and VH polarization modes, and its resolution is 3 m.
@@
 It is annotated with static objects, including bridges, harbors, and oil tanks in urban areas.
-OpenSARShip [67] is a ship slices dataset based on the European C-band Sentinel-1 satellite.
-Its resolution is 2.3 m to 17.4 m with VV and VH polarization.
-The dataset contains many ship slices from 10 busy ports.
-It has a diverse range of ship types but a significant category imbalance.
-SADD [30] is an aircraft detection dataset collected from the German X-band TerraSAR-X satellite.
-Its resolution is 0.5 m to 3 m with HH polarization.
-The dataset contains densely parked aircraft of different sizes on airport tarmacs and runways.
-It has a large number of small-sized planes as well as the airport perimeter area.
-SAMPLE [68] is a synthetic and measured paired fine-grained vehicle dataset released by the Air Force Research Laboratory, USA.
-This dataset is simulated in X-band and 0.3 m resolution.
-The public version provides 5,380 images of ten categories of vehicle targets at partial imaging angles.
-SAR-AIRcraft [24] is a aircraft detection dataset based on the Chinese C-band Gaofen-3 satellite with 1 m resolution and single polarization.
-The dataset collects seven types of aircraft of different sizes from three civil airports.
-It can support fine-grained aircraft detection and classification studies.
-SARSim [69, 70] is a fine-grained vehicle dataset created by Terma A/S, Denmark.
-The simulation system used for this dataset can generate X-band SAR images with resolutions ranging from 0.1m to 0.3m from CAD models.
-SARSim provides 21,168 vehicle samples in 7 categories (truck, car, motorbike, bus, tank, bulldozer, and pickup) and 3 scenes (grass, roads, and a mean of the two) with 7 imaging depression angles.
-SAR-Ship [71] is a ship target detection dataset in complex scenes based on Chinese Gaofen-3 and European Sentinel-1 satellites.
-The public version of this dataset contains 39,729 images from two satellites in different imaging modes and resolutions.
-The dataset provides ship targets of various sizes in complex ocean scenes such as nearshore, distant seas, harbors, and islands.
-SIVED [72] is a vehicle detection dataset with rotatable bounding box.
-It consists of vehicle slices from the MSTAR dataset [18] and vehicles in urban areas from the Sandia MiniSAR and FARAD datasets [64], and scenes include car parks, buildings, trees, roads, and others.
-SSDD [23] is a commonly used SAR ship detection dataset.
-It is constructed based on Canadian RadarSat-2, German TerraSAR-X, and European Sentinel-1 satellites and contains different scenarios for the inshore and offshore of China and India.
-The dataset covers various ship sizes in different oceanic conditions with diverse clutter and noise interference.
 A-B Classification dataset for performance Test
-We select three target classification datasets, including 25 fine-grained targets from vehicles, ships, aircraft, and others, to evaluate the comprehensive performance of SSL and the foundation model for SAR target recognition.
-The new SAR classification dataset named SAR-Target in Table VI.
-MSTAR [18] is the most commonly used SAR vehicle datasetw.
-It has many experimental setting variants, while we refer to the  [25] to adopt the most commonly used ten-class classification settings, such as infantry vehicle, patrol car, personnel carrier, main battle tank, and truck.
-FUSAR-Ship [29] contains 15 primary ship categories and many non-ship targets based on the Gaofen-3 satellite in scenes such as sea, land, coast, river, and island.
-Based on the experimental setting of [116], we have ten ocean target types, such as four fine-grained ships, bridges, and ocean scene slices.
-SAR-ACD [29] contains five types of aircraft based on the Gaofen-3 satellite in three civil airports.
-Since the released dataset does not separate the training and test data, we randomly select partial samples as the training set and others as the test set.
-Fine-grained recognition of aircraft targets is a more challenging task due to the smooth surface of the aircraft, resulting in insignificant SAR image features.
-A-C Hyperparameter settings
-TABLE VI: Description of our SAR classification dataset, named SAR-VSA, which contains 25 fine-grained targets. # Train: Number of train samples. # Test: Number of test sample images.
+TABLE VIII: Description of our SAR classification dataset, named SAR-VSA, which contains 25 target classes. # Train: Number of train samples. # Test: Number of test samples.
 Fine-grained category # Train # Test
@@
 aircraft (Boeing 787) 113 391
-TABLE VII: Pre-training setting.
+TABLE IX: Pre-training setting.
 Config Value
@@
 augmentation ResizedCrop, HFlip, ColorJitter
-TABLE VIII: Classification settings (fine-tuning and linear probing).
+TABLE X: Classification settings (fine-tuning and linear probing).
 Config Value
@@
 warmup learning rate 1e-5
-TABLE IX: Detailed classification results on MSTAR’s SOCs and EOCs. “-” refers to unknown. Bold indicates the best result, underline is the next best result.
+Figure 10: Partial fine-tuning results of HiViT-B and SSL-ImageNet & SAR.
+Tuning 0 blocks is linear probing.
+With the increasing number of fine-tuned transformer blocks, our method is consistently better than MAE and experiences overfitting later under the 5-shot setting.
+TABLE XI: Detailed classification results on MSTAR’s SOCs and EOCs. “-” refers to unknown. Bold indicates the best result, underline is the next best result.
 SOCs: Standard operating conditions (10 way)
@@
 PD [106] 2024 46.7 58.9 62.0 67.5 70.2 83.7 - -
-SARatrX 2024 85.2 91.4 93.9 95.1 95.9 97.7 98.1 98.5
+SARATR-X 2024 85.2 91.4 93.9 95.1 95.9 97.7 98.1 98.5
 EOCs: Depression angle variations (4 way)
@@
 CRID [104] 2023 62.1 62.3 - - 74.5 85.9 - 87.0
-SARatrX 2024 93.4 97.3 98.5 98.0 98.9 99.5 99.6 99.4
+SARATR-X 2024 93.4 97.3 98.5 98.0 98.9 99.5 99.6 99.4
 EOCs: Target configuration variations (4-way)
@@
 CRID [104] 2023 62.8 65.7 - - 74.1 78.7 - 84.1
-SARatrX 2024 65.0 74.0 74.0 84.2 78.3 87.0 88.9 93.5
+SARATR-X 2024 65.0 74.0 74.0 84.2 78.3 87.0 88.9 93.5
 EOCs: Target version configuration variations (4-way)
@@
 CRID [104] 2023 53.5 56.2 - - 67.2 79.9 - 87.8
-SARatrX 2024 65.3 76.5 76.3 83.4 82.8 85.0 92.1 97.0
-TABLE X: Detailed detection results.
+SARATR-X 2024 65.3 76.5 76.3 83.4 82.8 85.0 92.1 97.0
+We select three target classification datasets, including 25 fine-grained targets from vehicles, ships, aircraft, and others, to evaluate the comprehensive performance of SSL and the foundation model for SAR target recognition.
+The new SAR classification dataset is named SAR-Target in Table VIII.
+MSTAR [36] is the most commonly used SAR vehicle datasetw.
+It has many experimental setting variants, while we refer to the  [18] to adopt the most commonly used ten-class classification settings, such as infantry vehicle, patrol car, personnel carrier, main battle tank, and truck.
+FUSAR-Ship [69] contains 15 primary ship categories and many non-ship targets based on the Gaofen-3 satellite in scenes such as sea, land, coast, river, and island.
+Based on the experimental setting of [128], we have ten ocean target types, such as four fine-grained ships, bridges, ocean scenes, and false alarm slices.
+SAR-ACD [22] contains five types of aircraft based on the Gaofen-3 satellite in three civil airports.
+Since the released dataset does not separate the training and test data, we randomly select partial samples as the training set and others as the test set.
+Fine-grained recognition of aircraft targets is a more challenging task due to the smooth surface of the aircraft, resulting in insignificant SAR image features.
+A-C Hyperparameter settings
+TABLE XII: Detailed detection results.
 The metrics are mAP, $\rm{mAP_{50}}$ (@@@@50), $\rm{mAP_{75}}$ (@@@@75), $\rm{mAP_{small}}$ (@@@@s), $\rm{mAP_{medium}}$ (@@@@m), and $\rm{mAP_{large}}$ (@@@@l).
@@
 MSFA [39] 2024 56.4 88.2 61.5 50.5 66.5 62.5
-SARatrX 2024 57.2 (+0.8) 88.5 (+0.3) 62.6 (+1.1) 51.2 (+0.7) 69.6 (+3.1) 64.7 (+2.2)
+SARATR-X 2024 57.2 (+0.8) 88.5 (+0.3) 62.6 (+1.1) 51.2 (+0.7) 69.6 (+3.1) 64.7 (+2.2)
 OGSOD* (Object detection)
@@
 YOLOv7 [114] 2022 45.1 79.2 - - - -
-SARatrX 2024 51.4 (+6.3) 85.2 (+6.0) 51.7 46.8 75.3 77.1
+SARATR-X 2024 51.4 (+6.3) 85.2 (+6.0) 51.7 46.8 75.3 77.1
 SSDD (Ship detection)
@@
 $\rm{CS}^{n}$et [118] 2023 64.9 97.1 - - - -
-SARatrX 2024 67.4 (+2.5) 97.3 (+0.2) 83.1 (+6.9) 68.1 (+13.0) 66.1 (-2.1) 60.5 (-10.1)
+SARATR-X 2024 67.4 (+2.5) 97.3 (+0.2) 83.1 (+6.9) 68.1 (+13.0) 66.1 (-2.1) 60.5 (-10.1)
 SAR-Aircraft (Aircraft detection)
@@
 SA-Net [47] 2023 - 77.7 62.8 - - -
-SARatrX 2024 58.7 86.1 (+5.7) 64.7 (+3.3) 20.0 57.2 49.7
-Here are the detailed settings of our pre-training and downstream tasks as shown in Table VII and VIII.
-Pre-training - Our default pre-training setting is in Table VII, and other hyperparameters of each method use the default settings from their papers and codes.
+SARATR-X 2024 58.7 86.1 (+5.7) 64.7 (+3.3) 20.0 57.2 49.7
+Here are the detailed settings of our pre-training and downstream tasks as shown in Table IX and X.
+Pre-training - Our default pre-training setting is in Table IX, and other hyperparameters of each method use the default settings from their papers and codes.
 Our pre-training is applied on 8 NVIDIA RTX3090 GPUs with 200 epochs and 800 batch sizes.
@@
 It is worth noting that although MAE uses the normalized pixel value to enhance the feature representation in the visible spectral images, we find that the normalized pixel value cannot be used due to the SAR image noise and prevents the training loss from decreasing properly.
-Figure 8: Partial fine-tuning results of HiViT-B and SSL-ImageNet & SAR.
-Tuning 0 blocks is linear probing.
-With the increasing number of fine-tuned transformer blocks, our method are consistently better than MAE and experience overfitting later.
 Classification setting - All models use the same training settings in downstream classification tasks.
-Table VIII gives the default setting.
+Table X gives the default setting.
 Our few-shot learning setting is based on the Dassl toolbox [129, 130] and is averaged over 10 random experiments.
@@
 10 shows why we chose linear probing for few-shot evaluation, as HiViT overfits when many blocks are fine-tuned.
-Experimental results show that our SAR-KGPA consistently obtains better representations than MAE, and overfitting occurs later.
-Appendix B Implementation Details for Section 5 and the Leveraging of SARatrX for Recognition Tasks
+Experimental results show that our method consistently obtains better representations than MAE, and overfitting occurs later.
+Appendix B Implementation Details for Section 5 and the Leveraging of SARATR-X for Recognition Tasks
 B-A Dataset description
@@
 SOCs are in similar image conditions, and the training set’s depression angle under SOC is 17°, while the test set is 15°.
-Ten categories of targets include BMP2, BRDM2, BTR60, BTR70, T62, T72, 2S1, D7, ZIL131, and ZSU234 as shown in Table VI.
-The EOCs setting is the imaging condition variations to test robustness.
+Ten categories of targets include BMP2, BRDM2, BTR60, BTR70, T62, T72, 2S1, D7, ZIL131, and ZSU234 as shown in Table VIII.
+EOCs settings are the imaging condition variations to test robustness.
 Existing methods have been saturated with different experimental settings of MSTAR [18], but few samples and depression angle variations remain challenging.
@@
 Based on our scaling experiment, we use HiViT-B with 600 epochs pre-trained on SSL-ImageNet & SAR as the foundation model for classification and detection tasks.
-Classification setting is follow Table VIII. The only difference is that we use partial fine-tuning for better performance, and the last 6 blocks are added to the fine-tuned.
-Detection setting is follows the default setting in HiViT, and we adjust the learning rate to 5e-4.
-We used the same settings for each dataset fine-tuning, see our GitHub configuration based on the mmdetection [123] framework for details.
+Classification setting is follow Table X. The only difference is that we use partial fine-tuning for better performance, and the last 6 blocks are used to the fine-tuned.
+Detection setting follows the default setting in HiViT, and we adjust the learning rate to 5e-4.
+We used the same settings for each dataset fine-tuning, and our GitHub configuration is based on the mmdetection [131] framework for details.
 B-C Detailed results
-We provide detail classification and detection result on Table IX and X. Although the proposed method outperforms existing methods in mAP, there is still scope for improvement in some refined metrics.
+We provide detailed classification and detection results in Table XI and XII.
+Although the proposed method outperforms existing methods in mAP, there is still scope for improvement in some refined metrics.